A Complete Bibliography of the Journal of the Royal Statistical Society, Series A family: 1980–1989

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

09 November 2023 Version 1.02

Title word cross-reference

2 × 2 [611, 210, 453]. L_{1} [627]. R^{2} [401, 615, 509].

1 [13, 536, 343, 320, 709]. 10th [779]. 149th [380]. 151st [618]. 152nd [684]. 153rd [758]. 17th [11]. 18th [11]. 1900 [756]. 1933 [11]. 1964 [597]. 1971 [68]. 1976 [15]. 1979 [174, 231]. 1982 [478]. 1983 [779]. 1985 [917]. 1987 [942].

2 [24, 99, 12, 534]. 20th [743]. 21A [478]. 21B [478]. 2nd [153, 101, 25, 26].

3 [251, 144].

4 [482]. 42nd [773].

5th [174].

6 [622].

79 [146]. 7th [174].

A-Circular [481]. A. [364]. Abraham [634, 459]. Absolute [701]. Abstract [327]. Acceptance [355]. Accident [680, 952, 456]. Account [655]. Accounting [897, 677, 237]. Accounts [842, 777]. Accuracy [974, 66]. Acid [934]. Action [844, 825]. Activity [769, 123]. Actual [765]. Actuarial [792]. Actuaries [231]. Addendum [277]. Address [726]. Adjusted [648]. Adjusting [851].

Adjustment [63, 207, 528, 209, 237]. Administrative [894]. Advanced [801, 178, 480, 52, 716, 471]. Advances [358, 792, 876, 877]. Affected [528]. Against [11, 550]. Age [639, 316]. Age-Period-Cohort [639]. Aggregation [508]. Agresti [573]. Agriculture [446, 775, 488]. AIDS [764, 767, 743, 768, 763, 772, 773, 765]. Air [933]. Aitchison [744]. Akahira [342]. Akaike [884]. Akin [98]. Alan [558, 801, 573]. Albert [899]. Algebra [872]. Algorithms [795, 629]. Alice [418]. Allais [287]. Allen [545, 745]. Allied [664]. Allocation [658, 213, 120]. Allocative [704]. Aloke [663]. Alternative [921, 210, 579]. Altman [473]. Altruism [345]. Alvey [956]. Ambartzumian [402]. America [231]. American [262]. Analyses [657, 554]. Analysing [258, 282, 81, 125, 17]. Analysis [351, 216, 392, 549, 280, 326, 707, 714, 832, 126, 250, 85, 868, 13, 859, 174, 934, 889, 377, 70, 143, 18, 639, 555, 430, 890, 583, 605, 21, 944, 698, 907, 821, 479, 477, 749, 990, 99, 612, 92, 718, 885, 873, 203, 744, 36, 688, 61, 107, 622, 570, 506, 246, 409, 513, 616, 595, 581, 712, 343, 891, 649, 200, 214, 252, 292, 301, 480, 484, 553, 690, 734, 145, 865, 433, 354, 302, 624, 397, 26, 534, 556, 259, 746, 170, 625, 400, 17, 793, 416, 860, 319, 176, 17, 290, 414, 824, 829]. Analysis [671, 880, 795, 954, 901, 520, 463, 908, 603, 102, 706, 389, 221, 867, 132, 579, 673, 968, 574, 109, 223, 755, 842, 486, 297, 150, 72, 748, 573]. Analytic [960]. Andersen [141]. Anderson [168, 873, 900, 622, 421, 746, 15, 125, 422, 343, 142, 553, 574, 534]. Andrew [346]. Annette [469]. Annual [97, 195, 286, 383, 533, 621, 683, 730, 856, 15, 19, 94, 192, 283, 380, 530, 618, 684, 758]. ANOVA [753]. Anscombe [512]. Answers [296]. Anthony [585, 422]. APL [512]. Applicable [872]. Application [372, 216, 678, 335, 949, 239, 398, 456, 475]. Applications [959, 402, 915, 370, 495, 799, 689, 871, 966, 413, 600, 297, 542, 717, 466, 670, 410, 479, 477, 943, 141, 666, 108, 964, 295, 800, 984, 556, 515]. Applied [600, 46, 263, 461, 876, 464, 852, 136, 753, 326, 672, 290, 868, 824, 232, 990, 902, 629, 490, 815, 625]. Appreciation [264]. Approach [138, 440, 732, 69, 25, 26, 216, 493, 557, 690, 417]. Approaches [549, 918, 579, 257, 398, 540, 921, 411]. Appropriations [120]. Approximating [2, 185]. Approximation [412, 267]. Area [935, 796]. Arithmetical [947]. Armitage [814]. Arnold [30, 319, 939, 625]. Arrow [750, 536, 320]. Arthur [560, 298]. Ascher [575]. Asmussen [815]. Aspects [806, 770, 863, 118, 462, 647]. Assembly [773]. Assessing [979, 809]. Assessment [505, 740, 241, 864, 977, 418, 904]. Assessments [278]. Association [262, 231]. Asymptotic [342, 130, 342, 881, 169, 602]. Asymptotics [946]. Atkinson [901, 706]. Attack [932]. Attribute [521]. Auditors [560]. Austin [584]. Author [771]. Authorities [657]. Automatic [569]. Automation [919]. Autoregressive [491, 315]. Aven [859]. Average [189, 613]. Aviation [67]. Avoidable [348].

B [967, 81, 658, 199, 782, 832, 969, 825, 224, 752, 671, 910, 719, 881, 46, 521, 554, 792, 887, 152, 146, 876, 24, 491, 463, 477, 949, 141, 51, 623, 882, 864, 877, 69, 132, 299, 823, 353, 252, 253, 664, 751, 862, 329, 939, 953, 349, 471, 578, 519, 883, 746, 289, 755, 131, 78, 180, 468]. B.-C [969]. Babbage [585]. Bach [217]. Back [33, 57, 87, 115, 134, 158, 186, 205, 235, 255, 272, 310, 332, 362, 373, 395, 424, 450, 502, 524, 547, 564, 590, 607, 635, 643, 651, 674, 692, 699, 722, 736, 759, 804, 845, 892, 929, 970, 993]. Background [11]. Bailey [460]. Bain [18]. Balakrishnan [843]. Ball [685]. Balzer [384]. Bancroft [321]. Bandit [658]. Barnett [576, 322, 323]. Barrie [300, 721]. Bartels [143]. Bartholomew [344, 288]. Barucha [99]. Barucha-Reid [99]. Based [886, 399, 732, 398]. Basford [915]. Basic [886, 600, 46, 862, 25]. Basics [753]. Basis [245]. Basset [816]. Batchelor [245]. Batschelet [324]. Bawa [79]. Baxter [67]. Bayes [90, 242]. Bayesian [554, 926, 885, 411, 948, 117, 437, 242, 121, 508, 405, 779, 944, 302, 624]. Beale [695, 940]. Beard [499, 604]. Beaumont [461, 778]. Beckmann [599]. Beech [113]. Been [528]. before [756]. Begg [384]. Behaviour [152, 527, 198]. Behavioural [49, 770]. Being [910]. Belief [152, 7]. Belsley [166]. Belson [403]. Belt [653]. Bendat [707]. Benefit [378]. Bennett [70, 708, 338, 71]. Bensaber [783]. Berbee [74]. Berenson [902]. Berger [474]. Bergstrom [404]. Berkson [393]. Berkum [888]. Bernard [50]. Bernardo [779, 405]. Berry [658, 814]. Best [261]. Bethea [623]. Bets [613]. Betting [5]. between [239]. Beyond [753]. Bharucha [731, 13]. Bharucha-Reid [731, 13]. Bias [849]. Bibby [385]. Bibliography [642, 50]. Bickel [406]. Bilinear [718]. Billard [659]. Billingsley [73, 747]. Bina [110]. Binomial [611]. Bioassay [151]. Biological [294, 168, 222]. Biologists [249]. Biology [149, 324, 488]. Biomathematics [460]. Biomedical [146, 890]. Biometry [446]. Biostatistical [789]. Biostatistics [296, 224, 409, 296]. Birthday [406]. Bischler [167]. Bispectral [718]. Bivariate [241, 456]. Black [42]. Bland [817]. Blast [931]. Bliss [31]. Block [797, 219]. Blom [630]. BMDP [146]. BMDP-79 [146]. Boas [589]. Bogue [167]. Böker [818]. Boldy [263]. Book [294, 886, 130, 955, 967, 297, 348, 351, 793, 537, 600, 816, 669, 667, 878, 197, 960, 713, 797, 866, 828, 589, 262, 472, 668, 834, 836, 84, 216, 558, 81, 465, 756, 481, 839, 75, 392, 416, 658, 903, 950, 47, 884, 182, 199, 947, 965, 989, 790, 73, 869, 835, 511, 782, 23, 860, 29, 54, 41, 249, 518, 344, 914, 540, 542, 633, 516, 911, 912, 418, 817, 167, 560, 604, 720, 319, 750, 390, 15, 125, 326, 660, 672, 707, 714, 832, 28, 176, 177, 225, 300, 838, 17, 53, 126, 150]. Book [153, 250, 290, 85, 789, 868, 22, 13, 414, 711, 859, 48, 969, 824, 825, 149, 224, 752, 829, 717, 843, 226, 671, 910, 466, 630, 719, 946, 174, 468, 880, 342, 881, 959, 46, 263, 478, 355, 521, 661, 670, 961, 879, 461, 492, 538, 957, 515, 536, 801, 889, 410, 786, 554, 791, 920, 904, 792, 70, 143, 168, 512, 795, 522, 18, 870, 555, 887, 288, 152, 708, 82, 181, 146, 890, 402, 583, 77, 196, 267, 19, 605, 954, 901, 514, 129, 577, 580, 520, 151, 232, 404, 67, 229, 963, 330, 584, 991]. Book [915, 21, 951, 324, 788, 927, 154, 498, 111, 731, 876, 541, 350, 405, 517, 779, 926, 944, 757, 248, 220, 16, 802, 925, 818, 907, 942, 49, 50, 80, 148, 171, 215, 781, 218, 415, 408, 821, 826, 918, 391, 544, 24, 347, 491, 687, 747, 45, 101, 420, 463, 479, 493, 477, 489, 127, 217, 385, 585, 631, 110, 247, 749, 928, 949, 956, 222, 990, 99, 943, 27, 141, 359, 51, 356, 495, 718, 885, 732, 666, 873, 902, 924, 945, 203, 623, 744, 820, 844, 906, 908, 735, 882, 386, 601, 799, 576]. Book [603, 733, 74, 106, 108, 228, 327, 921, 626, 464, 900, 628, 322, 665, 104, 663, 575, 691, 917, 102, 251, 475, 688, 819, 629, 964, 706, 389, 12, 40, 52, 107, 221, 494, 622, 867, 79, 864, 783, 837, 877, 875, 69, 384, 659, 863, 132, 293, 710, 840, 539, 940, 689, 403, 166, 175, 299, 470, 11, 952, 823, 909, 246, 473, 144, 298, 346, 388, 487, 409, 513, 794, 784, 325, 353, 871, 557, 899, 923, 888, 827, 169, 745, 822, 874, 715, 490, 780, 485, 581, 83, 291, 712, 803, 992, 686, 343, 523]. Book [579, 716, 219, 754, 178, 597, 891, 958, 460, 587, 71, 72, 86, 142, 200, 214, 227, 252, 253, 265, 269, 292, 295, 301, 352, 459, 476, 480, 482, 484, 535, 553, 559, 586, 598, 664, 673, 690, 751, 862, 872, 966, 321, 323, 245, 421, 734, 412, 145, 497, 800, 25, 407, 543, 778, 919, 905, 231, 831, 865, 922, 320, 329, 76, 939, 941, 582, 953, 349, 296, 230, 916, 968, 984, 987, 14, 128, 471, 574, 578, 796, 354, 798, 913, 98, 287, 302, 345, 357, 411, 462, 474, 496, 624, 685, 519, 20, 103, 588, 358, 962]. Book [109, 985, 815, 26, 172, 406, 709, 534, 556, 467, 841, 328, 883, 266, 602, 510, 627, 173, 413, 483, 488, 746, 202, 223, 387, 830, 170, 986, 632, 268, 289, 469, 748, 753, 988, 147, 755, 948, 842, 861, 131, 662, 785, 721, 573, 486, 814, 833, 42, 44, 68, 100, 105, 179, 198, 201, 417, 625, 78, 180, 43, 264, 419, 787, 599, 570, 199]. Boorman [345]. Borooah [462]. Boullion [623]. Bovas [459]. Bowman [992]. Box [780, 417, 281, 401, 615]. Box-Jenkins [417, 281]. Bozdogan [860]. Bradford [584]. Braid [217]. Braithwaite [152]. Breslow [246]. Brian [494]. Briefing [773]. Brillinger [535, 597, 214]. Britain [266, 42, 727, 594, 932, 735, 985, 100]. British [895, 370, 653, 742, 917, 985, 68]. Broemeling [861, 624]. Brook [625]. Brown [224, 79, 887, 146]. Brownian [412]. Bruno [656]. Build [231]. Building [633, 132, 780, 709]. Bullen [903]. Bunday [862]. Bunke [709, 709]. Bunn [660]. Bureau [919]. Buringer [196]. Burnham [168, 421]. Burr [29]. Burton [712]. Business [21, 489, 990, 439, 269]. Buying [527, 984]. Byron [586].

C [793, 465, 416, 914, 542, 720, 126, 414, 969, 226, 946, 244, 786, 904, 143, 901, 638, 963, 24, 45, 101, 247, 928, 51, 885, 74, 626, 464, 900, 628, 691, 475, 819, 964, 706, 487, 874, 581, 86, 200, 265, 690, 966, 321, 421, 865, 329, 939, 984, 796, 913, 830, 170, 988, 948, 785, 68, 105, 625, 599]. C.-S [465]. Calculus [959, 410]. Calibration [39]. Cambridge [15]. Campbell [833]. Campion [552]. Canada [840]. Cancer [348, 161, 975, 514, 976, 246, 513, 647, 887].

Cancers [701]. Capital [582, 110]. Capture [421]. Capture-Recapture [421]. Carbon [39]. Carbon-14 [39]. Cardiovascular [258, 282]. Care [312]. Carl [711, 354]. Carlo [330]. Carlos [500]. Carrier [112]. Carter [513, 296, 296]. Carvalho [72]. Case [653, 240, 246]. Case-Control [246]. Case-Study [240]. Casebook [224]. Casino [5]. Casualties [653]. Casualty [931]. Categorical [297, 17, 824, 573, 400]. Categories [581]. Catherine [351, 550]. Causa [550]. Causal [583]. Causality [41]. Cause [506]. Causes [348, 575]. CB [331]. CBE [813]. Celia [67]. Censored [754]. Census [610, 569, 921, 435]. Censuses [359]. Centile [848]. Central [386, 144, 919]. Centuries [11]. Century [743]. Chain [139]. Chains [587, 412, 356, 412]. Chakravarti [169]. Challenges [776]. Chambers [463, 598, 270]. Change [140, 85, 861]. Changes [21, 340]. Changing [11, 985]. Characterization [822]. Charles [572, 803, 585]. Chater [346]. Chatfield [464, 170]. Chemical [577]. Chemicstry [889]. Chemotherapy [513]. Cheng [465]. Chernoff [781]. Chester [31]. Chi [946, 62]. Chi-Square [946]. Chi-Squared [62]. Chiang [465]. Childhood [978, 976, 896, 977]. Children [809]. Cho [791]. Cho-Yook [791]. Choice [826, 616, 595, 968, 79]. Chow [626]. Christensen [661, 710, 941]. Chronic [794]. Chung [942, 465]. cinlar [942]. Circular [481, 324]. Cirrhosis [639]. City [100]. Civil [67, 931]. Claims [720]. Clark [197, 868, 987]. Clarke [963]. Class [742]. Classes [886]. Classification [481, 702, 293]. Classification-Eye [481]. Classified [50, 581, 642]. ‘Classifying [76]. Cleland [904]. Cleveland [463, 662]. Cliff [466, 863]. Clifford [581]. Clinical [376, 392, 390, 889, 749, 378, 899, 398, 833, 793, 540, 229, 557]. Clinimetrics [870]. Clogg [581]. Closed [421]. Cluster [657, 889, 605]. Clustering [960, 915]. Clusters [240]. Cochran [588, 467, 304]. Coefficient [241, 491]. Coffman [859]. Cohen [325]. Cohort [639]. Colin [987]. Collected [465, 597]. Collection [22, 350, 211, 823]. Collinearity [166]. Colm [225]. Combination [336]. Combinatorial [402]. Comment [339]. Commerce [439]. Commission [248]. Committee [118]. Commodities [9]. Communities [821]. Company [259]. Comparative [127, 210, 660, 323]. Compared [639]. Comparison [657, 888, 874, 103, 136, 210, 315]. Comparisons [945]. Compilation [19]. Complex [92, 136]. Component [315]. Components [704, 964]. Compositional [744]. Comprehensive [527, 292]. Compstat [598]. Computational [598, 690, 220]. Computations [950]. Computer [399, 859, 146, 301, 932, 585, 659, 745]. Computer-Based [399]. Computing [429, 129, 512]. Concept [766]. Concepts [54, 342, 420, 886]. Concerning [740]. Concomitant [528]. Conduct [793]. Conference [15]. Confidence [521]. Configural [960]. Conjectured [64]. Conover [171]. Cons [894]. Consequences [528]. Consideration [338]. Considerations [616, 595]. Constance [419]. Constancy [175]. Constructing [9]. Construction [266]. Consultant [825]. Consumer [917, 198]. Contemporary [918]. Contingency [453]. Continuous [678]. Contribution [610, 467]. Contributions [838, 630, 77, 350, 871, 496]. Control [246, 988, 29, 355, 391, 544, 353, 253]. Controversies [54]. Convergence [386, 822]. Cook [407]. Cooke [294]. Cooper [289]. Cormack [145]. Cornell [514]. Cornfield [55]. Coronary [743]. Corporate [698]. Correction [303, 606].

Corrections [615, 616]. Correlation [241]. Correspondence [555]. Corrigenda [185, 184]. Corsi [626]. Cost [705]. Council [97, 195, 286, 383, 533, 621, 683, 730, 856, 980, 507, 550]. Countable [412]. Counterexamples [925]. Counting [108]. Countries [504]. Country [504]. Course [522, 820, 464, 173, 483, 902, 52]. Covello [864]. Coverage [727]. Cox [247, 819]. Coxon [468, 468]. Cramér [365, 650]. Creaming [117]. Crime [369, 696]. Criterion [5, 884, 401, 615, 455]. Critics [932]. Crop [678, 963, 810, 59]. Cross [39, 581, 583]. Cross-Classified [581]. Cross-Tabulation [583]. Cross-Validation [39]. Crow [943]. Csorgo [408]. Curnow [488]. Current [209]. Curve [122, 260]. Curves [848]. Cycle [21, 269]. Cycles [895]. Cyclic [827]. Cyert [944].

D [294, 886, 667, 797, 668, 658, 903, 47, 54, 249, 344, 167, 750, 15, 125, 660, 832, 176, 789, 824, 825, 466, 263, 521, 536, 889, 786, 168, 288, 152, 498, 779, 215, 408, 491, 493, 477, 217, 928, 902, 882, 733, 665, 917, 102, 629, 494, 622, 384, 863, 293, 166, 745, 992, 343, 597, 142, 484, 535, 553, 862, 245, 421, 734, 412, 800, 307, 320, 230, 987, 798, 624, 358, 172, 534, 883, 861, 131, 785, 721, 833, 264]. D’Agostino [782]. Daily [397]. Dalhousie [174]. Damned [385]. Daniel [229, 101, 409]. Danish [767]. Darwin [708]. Data [190, 297, 351, 313, 216, 849, 549, 767, 860, 516, 280, 593, 707, 17, 150, 725, 22, 848, 414, 824, 671, 880, 934, 786, 377, 795, 583, 520, 851, 101, 463, 479, 62, 92, 873, 603, 102, 680, 389, 61, 166, 211, 513, 616, 595, 581, 649, 257, 214, 471, 354, 109, 397, 647, 259, 136, 223, 209, 755, 315, 400, 81, 832, 143, 890, 924, 744, 576, 322, 388, 899, 923, 712, 746, 748, 662, 573, 661, 941]. Data-Oriented [216]. Dating [312]. Davenport [962]. David [326, 225, 414, 72, 214, 100, 945]. Davies [468, 577]. Davis [17]. Dawson [408]. Day [246]. Dean [346]. Death [506]. Deaton [198]. Debt [504]. Decade [366]. Decision [955, 781, 864, 132, 474, 398, 109, 26, 121, 43, 826, 178, 908, 784]. Decision-Making [26]. Decision-Support [398]. Decisions [668, 516, 390]. Decoding [725]. Decomposing [704]. Dedicated [350]. Defence [931]. DeGroot [866, 944]. Deistler [986]. Delivering [705]. Demand [646, 21]. Demographic [22, 104]. Demonstrations [886]. Demystifying [48]. Dennis [407]. Density [38, 168, 820, 14, 627, 755]. Deoxyribonucleic [934]. Dependence [719]. DePriest [757]. Derivatives [228]. Descriptive [555]. Design [793, 138, 797, 392, 8, 648, 85, 442, 749, 403, 480, 230, 833, 625, 126, 688, 865]. Design-Adjusted [648]. Designs [888, 219, 91, 797, 663, 299, 827]. Detection [922]. Determination [797]. Developing [504, 27]. Development [810, 100, 104]. Developments [443, 431, 442, 446, 440, 430, 438, 444, 436, 378, 439, 429, 432, 445, 433, 434, 437, 223, 150, 251, 12, 482]. Devroye [627, 820]. Dey [663]. Dhrymes [347]. Diagnostic [901, 706, 60]. Diagnostics [166]. Dice [516]. Dictionary [225]. Diet [239]. Dieulefait [500]. Different [505, 809]. Differential [959]. Diffusion [412]. Digby [821]. Directors [231]. Directory [928, 691]. Dirichlet [527]. Disaggregated [698]. Discoveries [117]. Discrete [600, 141]. Discriminant [698, 259]. Discrimination [136, 293]. Discussion [771, 282]. Disease [979, 794]. Display [543]. Displays [725]. Disposable [897]. Dissection [605]. Distance [140]. Distributed [347]. Distribution [767, 801, 705, 241, 2, 185, 454, 614, 850, 456, 329, 105, 961, 50]. Distribution-Free [961, 50]. Distributions [661, 2, 185, 943, 823, 316, 182, 734, 76, 939, 941]. Divisible [76]. Division [1, 184]. Dixon [146]. Dobson [469]. Dodge [748].

Doksum [406]. Doll [348]. Domain [687, 535]. Domesday [570]. Domestic [163]. Domison [100]. Don [911]. Donald [911, 711, 513, 490, 266, 113]. Double [8]. Douglas [757, 545, 80, 473]. Downton [561]. Dr [813]. Draper [290, 780]. Droesbeke [867]. Drost [946]. Drozzinov [967]. DSc [813]. Dual [297, 368, 211]. Dudley [388, 581]. Duncan [23, 581]. Dunn [868, 349, 471]. Duran [623]. Durant [394]. Duration [454, 314, 614]. During [312]. Dyke [963]. Dynamic [70, 391, 544, 594, 716, 180]. Dynamics [294, 764, 763].

E. [155]. Early [582]. Earnings [697]. Easton [869]. Ecological [226, 821]. Ecology [149, 145, 875]. Econometric [21, 626, 200, 751]. Econometrics [536, 24, 444, 302, 750, 959, 515, 965, 861]. Economic [377, 944, 990, 203, 207, 72, 916, 987, 237, 599]. Economics [19, 384, 388, 320, 42, 41, 632, 198]. Economies [43]. Economist [911]. Economists [716]. Economy [969, 462]. Ed [101]. Edgeworth [337]. Edgington [905, 147]. Editing [569]. Edition [153, 928, 25, 26]. Editors [937]. Education [366, 592]. Educational [81, 809, 742, 974]. Edward [757, 559, 543]. Edwards [199, 947]. EEC [551]. Effect [705, 208, 897]. Effectiveness [637]. Effects [653, 735, 932]. Efficiency [704, 678, 342, 965, 342, 498, 179]. Efron [224]. Eggleston [470]. Ehrenberg [984]. Eighth [283]. Elderly [240]. Elections [806]. Electoral [610]. Electrical [660]. Element [662]. Elementary [29, 781, 966]. Elements [586]. Elizabeth [803]. Elliot [691]. Elliott [928, 410, 346]. Elliptic [61]. Else [7]. Emergence [376]. Empirical [281, 907, 799, 780]. Employment [735, 36, 68, 247]. Encyclopedia [481, 481].

Engineering [78, 180, 948]. Engineers [623, 841]. England [639, 647, 435]. English [657]. Enhanced [610]. Enhancement [741]. Entering [742]. Entropy [47, 885, 948]. Environment [654]. Environmental [776, 979, 418]. Epidemic [768, 770, 772]. Epidemics [743, 863]. Epidemiology [769, 418, 978, 51, 794, 788]. Equal [248, 735]. Equations [101, 76]. Equity [179]. Eric [499, 406]. Erickson [948, 885]. Erling [141]. Error [481]. Errors [376, 91]. Escher [217]. Essay [162]. Essays [152, 987, 302, 496, 599, 582, 664]. Essentials [953]. Esther [301]. Estimate [481]. Estimates [348]. Estimating [455]. Estimation [648, 38, 404, 347, 885, 207, 239, 317, 840, 211, 169, 673, 14, 627, 755, 518, 820, 230, 168, 964, 79]. Estimators [342, 368]. Eternal [217]. Ethan [98]. Ethier [822]. Ethnic [334, 809]. Eubank [906]. Eugene [350]. Eugenics [708]. Eugenio [500]. European [916]. Evaluating [626]. Evaluation [312, 122, 455, 726]. Evelyn [695]. Events [317, 211]. Everitt [471, 825, 949, 349, 578]. Evidence [278, 667, 957, 470]. Evolution [739]. Ewens [148]. Exact [611]. Examination [281, 505, 593]. Examinations [215]. Example [671]. Examples [201, 351, 605]. Exercises [492, 24, 201]. Expectations [384, 667]. Expected [287]. Expenditure [313, 740]. Experiences [16]. Experimental [488, 625, 797]. Experimentation [631, 963]. Experiments [138, 678, 442, 851, 748, 658, 126, 749, 688, 865, 387]. Experimets [888]. Expert [276]. Explanatory [143]. Exploration [117, 471]. Exploratory [414, 143]. Exploring [583, 786]. Extraction [63]. Extrapolation [768]. Extremes [537]. Eye [481]. Eynon [746].

F [878, 828, 836, 216, 416, 947, 914, 319, 326, 153, 752, 719, 946, 880, 879, 786, 890, 67, 779, 818, 491, 101, 479, 359, 51, 732, 873, 389, 867, 539, 346, 490, 803, 734, 582, 916, 128, 913, 462, 413, 632, 43]. F. [337]. F.R.S. [695]. Factorial [663, 888, 299]. Facts [984]. Failed [448]. Failure [698, 259]. Family [313]. Farebrother [950]. Farmer [660]. Federer [299]. Feingold [575]. Feinstein [870]. Female [735]. Fertility [904, 91]. Festschrift [406]. Field [678]. Fienberg [264]. Fifth [406]. Fiftieth [530]. Fifty [428]. File [726]. Fill [806]. Filters [209]. Finance [914]. Financial [118, 259, 64]. Findings [960]. Findley [326]. Finetti [656]. Fingleton [671]. Finite [844, 734, 328, 795, 154]. Finney [249]. First [717, 522, 111, 427, 902, 173]. Fishburn [628]. Fisher [708, 264, 40]. Fit [946, 509, 782]. Fitting [848, 397, 101]. Fix [404]. Fix-Point [404]. Flamant [229]. Fleiss [749, 291]. Fleming [914, 913]. Florens [579, 411]. Fluctuations [582]. Fluoridation [550]. Folio [823]. Food [850, 520]. Force [551]. Forecast [117, 455]. Forecasting [281, 21, 698, 16, 853, 259, 66, 417, 316, 660, 990, 459, 86]. Forecasting. [15]. Forecasts [974, 64, 336]. Formal [338]. Formalism [47]. Formation [110]. Formulation [347]. Fortran [605]. Forty [94, 192, 283]. Forty-Eighth [283]. Forty-Seventh [192]. Forty-Sixth [94]. Foucart [783]. Foundation [409]. Foundations [484, 181]. Fowlkes [823]. Fractal [519]. Fractional [663]. Frame [726]. Framework [897]. Francis [512, 225, 927, 127, 52]. Frank [478, 561, 243, 858, 156]. Fraser [813, 172]. FRC [813]. FRCP [813]. Frederick [414]. Free [961, 50]. Freedman [412, 412]. Freeman [824]. French [785, 784]. Frequency [535]. Freund [102]. Friend [215]. Fristedt [658]. Front [3, 34, 58, 89, 116, 135, 159, 188, 206, 236, 256, 274, 311, 333, 363, 375, 396, 425, 451, 503, 526, 548, 565, 591, 609, 636, 644, 652, 676, 693, 700, 723, 738, 760, 805, 847, 893, 930, 972]. Frontiers [353]. FRS [813]. Frude [951]. Functional [167, 76]. Functioning [118]. Functions [955, 967, 732]. Fundamentals [711]. Further [164, 614, 338, 165]. Future [448, 921, 447, 117].

G [537, 797, 416, 884, 965, 720, 707, 300, 153, 859, 224, 671, 910, 630, 355, 461, 792, 514, 963, 915, 876, 16, 821, 491, 45, 489, 27, 885, 908, 799, 106, 228, 626, 688, 877, 539, 473, 353, 822, 891, 480, 559, 421, 778, 349, 471, 588, 467, 753, 948, 842, 721, 814, 201, 625]. G. [360]. Gabor [802]. Gabr [718]. Gain [5, 6]. Galen [513]. Galwey [956]. Game [839]. Games [5, 844, 542]. Gani [350, 664, 472]. Garneri [783]. Garwood [858]. Gary [751]. Gas [117]. Gastroenterology [398]. Geary [572]. Geer [907]. Gelfand [871]. Gelsema [686]. Gemmell [304]. General [94, 93, 192, 193, 283, 380, 318, 381, 530, 529, 618, 571, 617, 684, 758, 669, 153, 160, 773, 587]. Generalised [967, 154]. Generalization [239]. Generalizations [567]. Generalized [946, 239, 456, 485]. Generation [642]. Genetics [345, 148, 98]. Genn [83]. Genstat [956]. Gentle [129]. Geoffrey [497]. Geographic [258, 282]. Geographical [549]. Geometric [69]. Geometry [402, 800, 98, 519]. George [501, 305, 545, 780]. Geostatistics [197]. Gerhard [546]. Getoor [942]. Getz [149]. Gibbons [665]. Gilchrist [580]. Giri [126, 865]. Given [613]. Glazebrook [721]. Godel [217]. Goldberg [634, 666]. Golden [217]. Goldsmith [577]. Goldstein [85, 20]. Good [261, 100]. Goodman [17, 581]. Goodness [946, 509, 782]. Goodness-of-Fit [946, 782]. Goods [917]. Gopal [108]. Gordesch [598]. Gordon [113]. Gore [473]. Gottman [301, 292, 301]. Gottman-Williams [301]. Government [248, 238]. Grade [316]. Graham [872].

Grant [740]. Grant-Related [740].

Graphical [725, 901, 706, 463]. Graphics [757]. Graphing [662]. Graphs [628]. Gray [928, 501, 691]. Graybill [413]. GREA [740]. Greenwood [164, 2, 185, 165]. Greer [173]. Gregory [908]. Grenander [327]. Grether [72]. Griffiths [629]. Griliches [750, 536]. Groot [779]. Gross [711, 163]. Grouped [239]. Groups [809]. Groves [103]. Growth [487, 449, 247]. Guide [351, 468, 951, 301, 985, 746, 144]. Guided [631]. Gunnar [630]. Gunst [216]. Gupta [474, 860, 174]. Gyorfi [627].

H [537, 866, 839, 416, 860, 540, 633, 290, 85, 789, 824, 959, 961, 244, 889, 143, 152, 708, 196, 605, 404, 638, 779, 944, 80, 781, 477, 247, 928, 27, 945, 74, 575, 691, 384, 132, 513, 794, 353, 269, 690, 966, 734, 831, 357, 358, 985, 387, 861, 486, 68, 78, 180, 599]. H. [371]. Haberman [17, 150]. Habits [957]. Hackl [175]. Hagen [287]. Haggett [863]. Hair [873]. Hajek [328]. Hakim [351]. Halifax [174]. Hall [475, 386]. Halsey [985]. Hamburg [25]. Hampel [732]. Han [321]. Hand [825, 293]. Handbook [819, 661, 750, 250, 536, 535, 320, 329]. Hannan [687, 986, 664]. Harald [520, 365, 650]. Harmonic [486, 315]. Harn [76]. Harnett [176]. Harold [517, 302, 982]. Harris [711, 899]. Harry [415, 552]. Hartley [785, 233]. Harvey [20, 200, 265]. Hashmi [22]. Hauptmann [599]. having [581]. Hayek [582]. Hazard [680]. Hazards [567]. Hazel [83]. Head [136]. Health [312, 263, 791, 927, 654, 811, 409, 120, 788, 218]. Healy [229]. Hebden [515]. Heck [880]. Heckman [712]. Hedayat [299]. Heights [807]. Heiler [476]. Helga [709]. Hellevik [583]. Hellmut [520]. Helmet [754]. Henry [112, 394]. Herbert [871]. Heredity [708]. Herman [233]. Herrendorfer [797]. Hertz [477]. Herzberg [478]. Heterogeneity [769]. Heyde [475]. Heyer [387].

Heyworth [177]. Hicks [41]. Hierarchical [702]. Higher [342, 366, 592]. Hilary [562]. Hill [584, 629]. Hinckley [264]. Hiorns [294]. Historical [371]. History [756, 11, 568]. HIV [769, 764, 763, 770, 766]. HIV/AIDS [764]. Hoa [987]. Hoaglin [414, 786]. Hobcraft [104]. Hochberg [874]. Hodges [406]. Hoeffding [169]. Hofstadter [217]. Hogarth [826]. Hogg [84, 295]. Holden [667]. Holland [218]. Home [932]. Honarary [620]. Honor [871, 169, 599]. Honorary [96, 194, 285, 382, 532, 682, 729, 855]. Honour [630, 152, 664, 987, 302, 496, 406]. Horses [613]. Hospitals [120]. Howard [105]. Hsu [465]. Hubert [296, 46, 151]. Human [122, 852, 260, 853, 631]. Hundred [94, 192, 283, 530]. Hunter [600, 600, 35]. Huskova [130, 602]. Hutchinson [952]. Hydrology [877]. Hyman [585]. Hypotheses [457, 287, 790]. Hypothesis [153].

Ian [751]. Ideas [606, 541, 191, 225]. Identification [60]. Identifying [166]. II [481, 689, 180, 39, 137, 294, 326, 686, 602]. III [162, 474]. Ikeda [909]. Image [757]. Impact [588]. Implications [189, 420]. Import [917]. Imports [917]. Improvement [212]. Imputation [883]. Incidence [594]. Income [313, 987, 897, 105]. Incomes [346]. Incomplete [400, 219]. Increases [976]. Increments [74]. Incubation [769, 767, 766]. Index [124, 139]. India [110]. Indices [9]. Individual [654]. Induced [701, 977]. Industrial [443, 895, 314]. Industrialisation [245]. Industry [378, 577]. Inequalities [182, 654, 903, 869]. Infection [769, 763, 766]. Infectious [769]. Inference [90, 367, 915, 575, 40, 616, 595, 178, 321, 242, 709, 223, 828, 843, 881, 538, 517, 106, 327, 665, 323, 354, 172]. Inferences [379].

Infinitely [76]. Influence [732, 407]. Influential [166]. Influenza [863]. Informal [773, 338]. Information [88, 187, 273, 374, 525, 608, 675, 737, 846, 971, 399, 884, 725, 357, 543, 955]. Initial [593]. Injured [136]. Inland [67]. Innovation [247]. Input [457]. Input-Output [457]. Insights [370]. Installations [975, 976, 896, 977]. Instead [610]. Institute [262, 15, 646]. Institutions [118]. Insurance [231, 720]. Integral [402]. Integrated [52, 44]. Intellectual [836, 11]. Interaction [549]. Intercropping [191]. Interest [441, 435]. Interface [659, 745]. Intermediate [300]. International [717, 779, 69, 796]. Interpretation [899, 647]. Interpreting [849, 369, 696, 322]. Interval [628, 1, 184, 628]. Intervention [808]. Interviews [103]. Intriligator [536, 536, 320]. Intrilligator [750, 750]. Introduction [762, 817, 901, 45, 718, 666, 706, 784, 992, 292, 349, 230, 574, 578, 268, 469, 42, 417, 491, 956, 170, 177, 583, 517, 80, 949, 601, 940, 689, 388, 992, 354, 841, 289]. Introductory [819, 28, 176, 461]. Investigation [594]. Involving [240]. Ionizing [161]. Iron [878]. Ironmonger [987]. Irreducible [587]. Irvine [48]. Irwin [308]. Ishiguro [884]. Isobel [197]. Issues [637, 540, 83, 677, 22]. Itô [909]. Italian [838]. Ito [601]. Ittner [31]. IV [877]. Ivor [127].

J [600, 667, 713, 472, 839, 392, 416, 965, 989, 23, 249, 344, 912, 167, 560, 750, 390, 707, 832, 176, 177, 17, 150, 789, 868, 22, 825, 752, 466, 630, 959, 46, 536, 801, 410, 786, 904, 792, 70, 168, 18, 887, 288, 708, 82, 146, 267, 19, 129, 151, 229, 915, 154, 498, 876, 350, 405, 517, 779, 757, 802, 925, 148, 171, 408, 918, 347, 687, 463, 479, 813, 247, 732, 873, 744, 844, 799, 106, 665, 102, 964, 622, 79, 877, 863, 293, 689, 346, 325, 353, 557, 827, 485, 291, 579, 716, 891]. J [958, 460, 71, 352, 480, 484, 559, 586, 598, 664, 751, 145, 800, 922, 320, 329, 296, 987, 796, 411, 474, 588, 962, 406, 328, 830, 986, 469, 948, 785, 833, 198, 625, 78, 180, 43, 599]. J. [596, 306, 371]. Jack [331]. Jackson [388]. James [129, 712, 559]. Jana [77]. Janssen [479, 43]. Jauncey [550]. Jeffreys [982, 302]. Jenkins [417, 281, 360, 16]. Jerome [55]. Jewry [370]. Jews [370]. Jimmie [262]. Johannes [459]. John [41, 414, 48, 385, 104, 597, 292, 301, 354, 42, 827, 422, 219]. Johnson [481, 481, 2, 185, 787, 496]. Joint [646]. Jose [72]. Joseph [554, 229, 749, 308, 393]. Journal [761]. Judgement [826]. Judgements [212]. Juhl [390]. Julia [515]. Jureckova [77].

Kadane [554]. Kadekodi [965]. Kahn [220, 103]. Kalbfleisch [106]. Kanal [686]. Karl [11]. Karr [828]. Karvonen [788]. Kathleen [53]. Keith [801]. Kelly [128]. Kemp [516]. Kempson [963]. Kempton [821]. Kendal [178]. Kendall [558, 480, 800, 829, 221, 423, 801]. Kennedy [713, 129]. Kenneth [750, 536]. Kesten [415]. Ketellapper [143]. Kevin [721]. Key [468]. Keys [60]. Khatri [45]. Kingdom [878, 836, 914, 879, 67, 913, 703, 59]. Kingman [416]. Kirkwood [953]. Kitagawa [884]. Kiyosi [601]. Klas [598]. Kleffe [964]. Kleijnen [830]. Klein [79]. Kleiner [463]. Klotz [789]. Knowledge [741, 398, 266]. Knowledge-Based [398]. Kogan [859]. Kohlas [352]. Kolmogorov [364]. Kotz [481]. Krauss [69]. Krelle [599]. Kriegbaum [247]. Krishnaiah [687, 715, 988, 714, 250, 251, 12, 107, 482, 535]. Krotki [22]. Krzanowski [954]. Kshirsagar [483]. Kubik [955]. Kuh [166]. Kurtz [822]. Kuznetsov [910].

L [793, 669, 667, 960, 216, 465, 481, 182, 790, 869, 518, 542, 418, 176, 17, 224, 843, 226, 881, 46, 961, 70, 168, 555, 82, 19, 577, 927, 942, 781, 749, 928, 943, 873, 902, 623, 820, 906, 733, 691, 389, 864, 659, 940, 299, 513, 794, 291, 686, 579, 72, 598, 245, 421, 496, 624, 103, 861, 785, 43]. L. [244, 638]. Laake [168]. Lab [886]. Labor [712]. Laboratory [899]. Labouchere [65]. Labour [551]. Lags [347, 222]. Lallouch [229]. Lambert [484, 716]. Lamborn [82]. Lancaster [179]. Lane [956, 864]. Lanke [630]. Lansdowne [695]. Large [555, 237]. Larson [78, 180, 517]. Last [428]. Laszlo [627]. Latent [17, 578]. Latham [562]. Lattice [108]. Launer [389]. Lawyers [279]. Leadbetter [537]. Learning [40]. Least [950, 523]. Least-Squares [523]. Lebart [555, 598]. Lechin [390]. Lecture [219, 920]. Lectures [11, 598, 869]. Ledolter [459]. Lee [18, 957, 68]. Leech [51]. Legal [275]. Legislation [653, 735]. Lehmann [406, 790, 518]. Length [316]. Lenz [353]. Leo [581]. Leonard [262, 708]. Leontief [457]. Leroy [922]. Lesson [791]. Letters [937]. Leukaemia [978, 896, 977]. Level [52]. Levels [161, 218]. Levine [47, 902]. Levitt [345]. Lewis [576]. Ley [35]. Library [280]. Life [720, 18, 689, 506, 231, 685, 419]. Light [7]. Limit [386, 475]. Limits [521, 131]. Lindgren [537, 630]. Lindholm [875]. Lindley [779, 668]. Lindner [49]. Line [168]. Linear [669, 319, 672, 153, 717, 232, 490, 485, 523, 872, 357, 624, 172, 709, 483, 986, 209, 17, 950, 989, 912, 493, 923, 958]. Linhart [831]. List [458, 59]. Literary [612]. Literature [399]. Little [832]. Liver [639]. Load [660]. Local [705, 207]. Location [968, 679]. Location-Shifts [679]. Log [17]. Log-linear [17]. Logistic [122, 852, 260]. Lognormal [454, 943]. London [427]. Long [340]. Long-Run [340]. Longitudinal [85, 400, 712]. Look [111]. Lord [550]. Losing [613]. Louis [20, 540]. Low [161]. Loyalty [140]. Loyalty-Distance [140]. Loynes [900]. Lu [465]. Luc [627]. Lukacs [350]. Lwanga [791]. Lynx [315].

M [130, 537, 669, 866, 81, 756, 839, 903, 47, 884, 869, 360, 914, 817, 390, 28, 177, 126, 711, 149, 843, 226, 468, 342, 478, 795, 555, 146, 67, 229, 963, 788, 927, 731, 405, 779, 944, 220, 16, 408, 826, 24, 687, 45, 463, 479, 990, 718, 732, 666, 902, 623, 733, 900, 69, 940, 473, 298, 513, 888, 169, 745, 579, 219, 178, 72, 252, 253, 292, 301, 480, 598, 664, 734, 145, 865, 922, 320, 296, 796, 913, 287, 685, 103, 962, 556, 841, 602, 483, 986, 833, 105, 417, 43]. MA [813]. MacDonald [222]. MacDougall [911]. Machin [833]. MacKenzie [266]. Maclean [83]. MacNeil [751]. MacNeill [792, 876, 877]. Macroeconomics [384, 42]. Maddex [305]. Magidson [17]. Magnus [959, 989, 912, 958]. Major [743, 245]. Making [472, 167, 26, 43, 864, 668]. Makov [734]. Makridakis [86]. Malaria [460]. Management [927, 864]. Managers [181]. Mandarin [911]. Mandelbrot [519]. Mandl [130, 602]. Mandrekar [486]. Manly [752]. Manpower [316]. Manual [963, 44]. Manuscript [739]. Many [806]. Marc [72]. Marcotorchino [479]. Marginalia [137]. Mark [787]. Market [712, 968]. Markov [154, 356, 587, 822, 412]. Martens [520]. Martin [695, 70, 599, 803]. Martingale [475]. Mason [216]. Matching [8]. Materials [677]. Mathematica [478]. Mathematical [294, 713, 262, 633, 337, 402, 267, 802, 763, 909, 484, 320, 349, 962, 600, 600, 149, 288, 148, 632]. Mathematicians [415]. Mathematics [784, 716, 685]. Maths [199]. Matlock [928]. Matrices [851, 356, 872, 237, 555, 413]. Matrix [959, 228]. Matter [33, 57, 87, 115, 3, 34, 58, 89, 134, 158, 186, 205, 116, 135, 159, 188, 235, 255, 272, 310, 206, 236, 256, 274, 332, 362, 373, 395, 311, 333, 363, 375, 424, 450, 502, 524, 547, 396, 425, 451, 503, 526, 564, 590, 607, 635, 548, 565, 591, 609, 643, 651, 674, 692, 636, 644, 652, 676, 699, 722, 736, 759, 693, 700, 723, 738, 804, 845, 892, 760, 805, 847, 929, 970, 993, 893, 930, 972]. Matthews [487]. Mature [117]. Matusita [603, 223]. Maunder [914, 67, 913, 878, 836, 879]. Maurice [558, 829, 221, 423]. Mavis [83]. Maximum [47, 885, 948]. Maximum-Entropy [948]. May [478]. McCloughrey [582]. McColl [550]. McCullagh [485, 837]. McLachlan [915]. McQuitty [960]. Md [408]. Mead [488]. Means [82, 903]. Measure [509, 73, 747]. Measurement [756, 707, 85, 897, 991, 218]. Mecke [800]. Medical [849, 817, 920, 584, 432, 231, 953, 814]. Meeting [975, 779, 94, 93, 192, 193, 283, 380, 318, 381, 530, 529, 618, 571, 617, 684, 758]. Megalithic [61]. Meinhert [793]. Mellor [152]. Memoirs [911]. Memorial [262]. Memoriam [162]. Method [960, 9, 991, 945, 117, 616, 595, 20, 209, 330]. Methoden [818]. Methodes [783]. Methodological [83]. Methodology [838]. Methods [351, 372, 669, 752, 226, 335, 288, 890, 901, 514, 577, 851, 50, 463, 979, 949, 359, 623, 733, 102, 706, 837, 840, 689, 909, 246, 794, 291, 258, 282, 523, 295, 352, 459, 421, 471, 709, 488, 202, 632, 948, 721, 814, 300, 819, 715, 490, 685, 358, 268, 44]. Meyer [789]. Michael [782, 750, 536, 361]. Microcomputer [886, 928, 691]. Microcomputer-Based [886]. Microcomputing [789]. Mikheev [788]. Miller [224, 753]. MINITAB [746]. Minorities [334]. Minton [102]. Misclassification [957]. Misconceptions [575]. Missing [832, 748]. Mitrinović [903]. Mixture [734, 915]. Model [457, 567, 527, 698, 680, 852, 780, 397, 709, 504, 140, 765, 889, 357, 633, 132, 831]. Model-Building [780]. Modeling [860, 887]. Modelling [190, 551, 637, 764, 549, 90, 149, 16, 653, 493, 763, 575, 701, 853, 690, 580, 710, 751, 471, 469]. Models [600, 678, 344, 319, 660, 17, 969, 717, 466, 915, 21, 491, 141, 175, 923, 411, 685, 932, 709, 64, 455, 315, 78, 180, 669, 18, 222, 718, 626, 485, 265, 578, 624, 172, 483, 43]. Modern [25, 42, 389, 794, 295, 105]. Modified [852]. Modigliani [916]. Moghissi [864]. Mohanty [108]. Money [582]. Monotonicity [6]. Montagnon [181]. Monte [330]. Moolgavkar [794]. Moore [54, 489]. Moran [983]. Morgan [917, 586, 245]. Morineau [555]. Morris [25, 354]. Morrison [490]. Mortality [639, 370, 239, 506, 258, 282]. Moses [224, 781]. Mosler [599]. Moss [20]. Mosteller [414, 786]. Motion [412]. Mouchart [579]. Mrs [550]. Muellbauer [198]. Multi [506, 121]. Multi-Bayesian [121]. Multi-State [506]. Multidimensional [468, 257, 109]. Multiple [874, 883]. Multivariate [182, 319, 567, 954, 45, 322, 253, 433, 574, 860, 714, 752, 829, 226, 174, 880, 555, 821, 873, 107, 221, 710, 539, 899, 787]. Murphy [176]. Murtagh [880].

N [967, 589, 364, 481, 75, 390, 126, 290, 843, 492, 957, 70, 951, 408, 821, 918, 956, 27, 909, 246, 822, 686, 227, 865, 987, 796, 354, 488]. Naddeo [838]. Nagel [889]. Nation [897]. National [897, 103, 777, 43, 120, 987, 842]. Nations [773]. Natural [445, 708]. Nature [519]. Nazem [990]. Near [975, 979]. Neave [961]. Need [916]. Negative [587, 356, 872]. Neglected [935]. Neil [968]. Nelder [485]. Nerlove [72]. Nersessian [918]. Netherlands [919]. Networks [128]. Neudecker [959]. Newer [420]. Next [447, 366]. Neyman [306, 419]. Nicholls [491]. Nightingale [119, 137, 162]. Nijkamp [109]. Nishisato [297]. Nixon [596]. No [144]. Noise [180]. Non [720, 727, 50, 341, 587, 379, 411, 268, 669, 493, 356, 872]. Non-Bayesian [411].

Non-Coverage [727]. Non-Life [720]. Non-Linear [493]. Non-Negative [587, 356, 872]. Non-Parametric [50, 341, 411, 268, 669]. Non-random [379]. Nonparametric [80, 171, 906, 733, 196, 665, 715, 627]. Nonresponse [883]. Normal [241, 754, 329, 131]. Norman [222, 780, 460, 496, 112]. Note [507, 123, 808, 65, 371, 766, 317, 614, 649, 66]. Notes [920, 260, 219]. Nuclear [975, 976, 677, 932, 896, 977]. Number [317, 211, 642]. Numbers [727, 139, 371, 592]. Nummelin [587]. Nutrition [420].

O [839, 15, 125, 868, 859, 577, 844, 622, 487, 343, 142, 553, 987, 287, 474, 534, 78, 180]. OBE [499]. Obituary [813, 982, 858, 983]. Observations [239, 539]. Occupational [788]. October [174]. Odds [613]. Odeh [521, 962, 131, 713]. O’Donovan [417]. Office [144, 932]. Official [431, 377, 811, 144]. O’Hagan [991]. Oil [117]. Okun [298]. Olaf [709]. Old [655]. O’Muircheartaigh [225]. One [94, 192, 283, 530, 806]. Online [399]. OPCS [726]. Open [969]. Operational [443, 263, 785]. Operations [352, 599]. Operators [587]. Opinion [37, 7, 550, 66]. Opportunities [248]. Optimal [79, 75, 888, 230]. Optimality [498]. Optimisation [391, 544]. Optimization [795, 949, 940, 716]. Ord [622, 466, 801, 863, 480, 145]. Order [678, 342, 719]. Ordered [628, 581]. Orders [628]. Ordinal [573]. Oriented [216]. Orloci [226]. Orthogonal [663]. Osbourne [795]. Oscar [308]. Otis [581, 421]. Ottar [583]. Otto [233]. Outlier [922]. Outliers [576]. Outlines [791]. Output [457]. Overview [931]. Owen [521, 131].

P [130, 669, 589, 465, 903, 73, 633, 714, 250, 174, 468, 961, 461, 957, 143, 168, 181, 577, 154, 220, 821, 391, 24, 347, 687, 747, 420, 463, 489, 956, 732, 924, 386, 74, 921, 626, 628, 251, 475, 819, 629, 12, 107, 837, 863, 132, 175, 952, 353, 715, 780, 485, 579, 716, 219, 484, 535, 598, 321, 421, 778, 922, 128, 796, 411, 496, 588, 406, 602, 746, 830, 988, 814, 146]. P-Series [146]. Paired [945, 888]. Pakistan [22]. Pandit [556]. Pao [465]. Pao-Lu [465]. Paper [339]. Papers [465, 350, 664, 558]. Paradox [287]. Paradoxes [802]. Parameter [648, 230]. Parametric [50, 341, 411, 268, 669]. Parents [742]. Pareto [939]. Parsons [26]. Partially [628]. Party [646]. Paruchuri [482]. Pascal [789, 947]. Pass [215]. Passive [957]. Pat [983]. Patel [329]. Path [108]. Patients [136]. Pattern [671, 61, 960]. Pattern-Analytic [960]. Patterns [370, 686]. Pau [43]. Paul [102, 270, 345, 100]. Pay [735]. PC [951]. Pearce [963]. Pearson [11, 155, 11]. Peel [667]. Penetration [917]. Pentikainen [604]. Percentage [164, 165]. Perception [725]. Percolation [415]. Percy [501]. Performance [505]. Period [639, 766]. Periodicals [32, 56, 114, 133, 157, 183, 204, 234, 254, 271, 309]. Periods [769]. Perkins [987]. Personal [372, 443, 431, 776, 442, 446, 440, 335, 740, 430, 438, 444, 657, 436, 439, 429, 432, 445, 433, 434, 437, 103]. Pesonen [604]. Peter [544, 523, 673, 109, 268]. Peters [839]. Petersen [688]. Peto [348]. Petrie [920]. Pharmaceutical [378]. Phenomena [803]. Phillip [104]. Phillips [24, 67]. Philosophical [918, 40]. Philosophy [384]. Photosynthetic [875]. Phylogenies [367]. Physical [495]. Physics [909]. Phytotoxic [933]. Piersol [707]. Pioneer [585]. Pitt [225]. Planners [202]. Planning [965]. Plans [8, 521, 131]. Plant [123, 30]. Platek [796]. Plews [28]. Ploeg [462, 632]. Plots [823, 901, 706]. Plus [769]. Pocock [557]. Point [518, 671, 828, 404]. Points [164, 165]. Policies [916]. Policy [4, 100, 420, 346]. Political [462]. Polls [37, 66]. Pollutants [933]. Pollution [979]. Polynomials [731]. Population [764, 167, 705, 19, 148, 921, 894, 98, 435, 328, 840]. Populations [294, 641, 122, 852, 260, 853, 168, 754, 421]. Portfolio [79]. Ports [67]. Position [443, 431, 442, 446, 440, 430, 438, 444, 436, 439, 429, 432, 445, 433, 434, 437]. Poskett [484]. Possible [979]. Postcode [726]. Potential [443, 610, 431, 442, 446, 440, 430, 438, 444, 436, 439, 429, 432, 445, 433, 434, 437]. Prabhu [227]. Practical [886, 493, 900, 378, 197, 560, 16, 171]. Practice [558, 139, 622, 686, 343, 534, 404, 918, 473, 553]. Practitioners [830]. Prais [10, 339]. Prakasa [881]. Pratical [557]. Pratiques [783]. Predicting [504]. Prediction [253, 523, 486]. Prentice [418, 794]. Prequential [440]. Present [443, 431, 442, 446, 440, 430, 438, 444, 436, 439, 429, 432, 445, 433, 434, 437]. President [95, 284, 531, 619, 681, 728, 854]. Price [9, 124, 139, 613, 340]. Prices [64, 777, 298]. Priestley [252, 253, 664]. Primary [312]. Primer [53, 752]. Principle [160]. Principles [954]. Printing [879]. Prior [241, 357]. Probabilistic [13, 99, 909, 78, 180]. Probabilities [36, 679]. Probability [600, 774, 869, 630, 492, 275, 876, 350, 517, 802, 613, 666, 601, 436, 212, 966, 815, 841, 53, 461, 522, 220, 925, 470, 992, 866, 589, 416, 182, 73, 670, 991, 926, 747, 106, 778]. Probable [481]. Problem [849]. Problems [816, 658, 727, 50, 347, 40, 677, 504, 369, 885]. Procedure [117]. Procedures [707, 196, 874]. Proceedings [94, 93, 192, 193, 283, 380, 318, 381, 426, 530, 529, 618, 571, 617, 684, 758, 15, 174, 779, 659, 745, 717]. Process [918]. Processes [828, 466, 111, 876, 942, 799, 436, 867, 822, 484, 664, 955, 537, 344, 907, 227, 78]. Processing [757, 559, 919]. Product [163]. Production [704, 895, 577]. Professor [496, 339, 499, 338, 307, 30]. Programming [391, 544, 633, 154]. Programs [146, 605, 301]. Progress [703, 809, 987]. Projection [694]. Projections [592, 167]. Prokaryotes [875]. Proof [470]. Properties [537]. Property [836]. Proportion [317]. Proportional [567, 680]. Proportions [521, 291]. Pros [894]. Prospects [448, 447]. Proth [479]. Provide [312]. Provinces [117]. Psych [813]. Psychiatric [38]. Psychology [826]. Public [441, 4, 435]. Publication [849]. Publishing [879]. Pukkila [717]. Punktprozessen [818]. Puntanen [717]. Puri [669]. Purpose [21]. Pursuit [694].

Qualifications [742]. Qualitative [17, 150]. Qualitative/Categorical [17]. Quality [441, 29, 355, 811, 353, 988]. Quantile [823]. Quantile-Quantile [823]. Quantitative [348, 725, 671, 543]. Quantities [298]. Quantity [189]. Question [208]. Questions [403, 296]. Quetelet [119, 137, 162]. Queue [325]. Queueing [711, 862, 289]. Queues [815]. Quick [268]. Quinn [491]. Quotes [385, 385].

R [294, 348, 537, 713, 589, 84, 950, 47, 989, 912, 604, 714, 832, 177, 250, 290, 789, 224, 226, 174, 959, 521, 661, 961, 410, 70, 143, 168, 795, 870, 887, 152, 708, 890, 402, 267, 514, 404, 67, 229, 963, 330, 111, 926, 944, 942, 821, 826, 24, 687, 217, 27, 371, 885, 732, 873, 623, 906, 799, 900, 251, 688, 964, 389, 12, 40, 107, 79, 864, 783, 166, 487, 794, 923, 715, 780, 992, 597, 958, 71, 482, 535, 245, 421, 145, 407, 543, 941, 953, 796, 345, 588, 962, 26, 488, 202]. R [289, 753, 988, 948, 131, 785, 625, 264]. R. [264]. Race [696]. Radiation [161, 701, 977]. Radiation-Induced [701, 977]. Rahman [492]. Rainfall [397]. Raktoe [46, 299].

Ralph [782]. Randles [80]. Random [955, 537, 436, 1, 184, 778, 642, 78, 180, 379, 707, 111, 731, 491, 74]. Randomised [390]. Randomization [905, 147]. Rao [226, 408, 718, 964, 796, 588, 881, 687, 988]. Raouls [579]. Rare [641]. Rasch [797, 733]. Rates [551, 976, 931, 291, 386]. Ratio [5, 6, 259]. Rational [384]. Rationale [740]. Rationality [784]. Ratkowsky [493]. Rayner [238]. Re [36, 122, 338]. Re-Consideration [338]. Re-Employment [36]. Re-Evaluation [122]. Read [329]. Readable [199]. Readings [873]. Real [897]. Reason [910]. Recall [20]. Recapture [421]. Recognition [686]. Recommended [59]. Reconciliation [777]. Record [368]. Recovery [916]. Recursive [673]. Redfern [921]. Rees [104]. Reference [848, 577]. Reform [974]. Regional [550, 68, 104]. Registers [610, 894]. Registration [647]. Regression [290, 442, 213, 901, 493, 92, 341, 239, 706, 452, 175, 616, 595, 922, 358, 120, 625, 672, 868, 232, 906, 407, 216, 907, 102, 166, 513]. Regulation [523]. Reid [731, 13, 99, 419]. Rejoinder [213]. Related [537, 839, 312, 13, 740, 555, 408, 99, 474]. Relating [457, 369]. Relationship [239]. Relative [371, 340, 701]. Relevant [191, 647, 455]. Reliability [313, 18, 626, 575, 237, 988]. Relief [850]. Religion [913]. Religious [11]. Removal [421]. Renewal [74]. Repairable [575]. Repeat [984]. Repeat-Buying [984]. Reply [771]. Report [97, 96, 95, 195, 194, 286, 285, 284, 383, 382, 533, 532, 531, 621, 620, 619, 683, 682, 681, 646, 730, 729, 728, 856, 855, 854, 980, 703]. Reported [976]. Requirements [772]. Research [351, 960, 443, 81, 174, 263, 577, 520, 191, 246, 785, 599, 19, 927, 352, 321, 814, 19]. Reserving [720]. Residuals [407]. Response [780]. Responses [208].

Responsibility [697, 973]. Results [258, 282]. Retrieval [399]. Reuter [416]. Revenue [213, 120]. Reversibility [128]. Review [294, 886, 130, 955, 967, 297, 348, 351, 793, 537, 600, 372, 816, 669, 667, 878, 197, 960, 713, 797, 866, 828, 589, 262, 472, 668, 834, 836, 84, 216, 558, 81, 465, 756, 481, 839, 75, 392, 416, 658, 903, 950, 47, 884, 182, 199, 947, 965, 989, 790, 73, 869, 835, 511, 782, 23, 860, 29, 54, 41, 249, 518, 344, 914, 540, 542, 633, 516, 911, 912, 418, 817, 167, 560, 604, 720, 319, 750, 390, 15, 125, 326, 660, 672, 707, 714, 832, 28, 176, 177, 225, 300, 838, 17, 53, 126]. Review [150, 153, 250, 290, 85, 789, 868, 22, 13, 414, 711, 859, 48, 969, 824, 825, 149, 224, 752, 829, 717, 843, 226, 671, 910, 466, 630, 719, 946, 174, 468, 880, 342, 881, 959, 46, 263, 478, 934, 355, 521, 661, 670, 961, 879, 461, 492, 538, 957, 515, 536, 801, 889, 410, 786, 554, 335, 791, 920, 904, 792, 70, 143, 168, 512, 795, 522, 18, 870, 555, 887, 288, 152, 708, 82, 181, 146, 890, 402, 583, 77, 196, 267, 19, 605, 954, 901, 514, 129, 577, 580, 520, 151, 232, 404, 67, 229, 963]. Review [330, 584, 976, 991, 915, 21, 951, 324, 788, 927, 154, 498, 111, 731, 876, 541, 350, 405, 517, 779, 926, 944, 702, 757, 248, 220, 16, 802, 925, 818, 907, 942, 49, 50, 80, 148, 171, 215, 781, 218, 415, 408, 821, 826, 918, 391, 544, 24, 347, 491, 687, 747, 45, 101, 420, 463, 479, 493, 477, 489, 127, 217, 385, 585, 631, 110, 247, 749, 979, 928, 949, 956, 222, 238, 990, 99, 943, 27, 612, 141, 359, 51, 356, 763, 495, 718, 885, 732, 770, 666, 873, 902, 924, 945, 203, 623, 744]. Review [820, 844, 906, 908, 735, 208, 882, 386, 601, 799, 576, 603, 733, 74, 106, 108, 228, 327, 921, 626, 464, 900, 628, 322, 665, 104, 663, 575, 691, 917, 102, 251, 475, 688, 819, 629, 964, 706, 389, 12, 40, 52, 107, 221, 494, 622, 867, 79, 864, 783, 837, 877, 191, 875, 69, 384, 659, 863, 132, 118, 293, 710, 840, 539, 940, 689, 403, 166, 175, 299, 470, 11, 952, 823, 909, 246, 473, 550, 144, 298, 346, 388, 487, 409, 513, 794, 784, 325, 353, 871, 557, 899, 923, 888, 827, 60, 169, 745, 822]. Review [874, 715, 490, 780, 485, 581, 83, 291, 712, 803, 992, 686, 343, 523, 579, 716, 219, 754, 178, 597, 891, 958, 460, 587, 71, 72, 86, 142, 200, 214, 227, 252, 253, 265, 269, 292, 295, 301, 352, 459, 476, 480, 482, 484, 535, 553, 559, 586, 598, 664, 673, 690, 751, 862, 872, 966, 321, 323, 245, 421, 734, 412, 145, 497, 800, 25, 407, 543, 778, 919, 905, 231, 831, 865, 922, 320, 329, 76, 939, 941, 582, 953, 349, 296, 230, 916, 968, 984, 987, 14, 128, 471, 574, 578, 796, 354, 798, 913, 98, 287, 302]. Review [345, 357, 411, 462, 474, 496, 624, 685, 519, 20, 103, 588, 358, 962, 109, 985, 815, 26, 172, 406, 709, 534, 556, 467, 841, 328, 883, 266, 602, 510, 627, 173, 413, 483, 488, 746, 202, 223, 387, 830, 170, 986, 632, 268, 289, 469, 748, 753, 988, 147, 755, 948, 842, 861, 131, 662, 785, 721, 573, 486, 814, 833, 42, 44, 68, 100, 105, 179, 198, 201, 417, 625, 78, 180, 43, 264, 419, 787, 599]. Reviews [303, 458, 703, 67, 878, 836, 914, 879, 67, 913]. Revised [931]. Revision [209]. Revolution [384]. Richard [216, 470]. Rights [836]. Rindskopf [902]. Ripley [882, 494]. Risk [418, 312, 161, 979, 79, 864, 504, 977, 489, 604, 477]. Risk-Related [312]. Risks [348, 38, 701]. Rivett [132]. Road [653]. Roads [952]. Robert [216, 499, 361, 229, 346, 295, 103, 572]. Roberts [813]. Robin [346]. Robinson [992, 670, 495, 622, 992]. Robust [414, 843, 732, 922]. Robustness [90, 616, 595, 554, 733]. Rogers [228]. Rohatgi [538, 350]. Role [85, 276, 677]. Roll [974]. Rolph [354]. Romaniuc [840]. Ronald [80, 710]. Ronchetti [732]. Rootzen [537]. Ross [418, 522, 841]. Rousseeuw [732, 922]. Roy [582, 545, 110]. Royal [761, 427, 94, 93, 192, 193, 283, 380, 318, 381, 530, 529, 618, 571, 617, 684, 646, 758, 975, 428, 449]. Rubin [832, 883]. Rubinstein [330]. Rudolf [102]. Rules [75]. Run [340]. Russia [568]. Russwurm [520]. Rustagi [689].

S [955, 967, 297, 465, 756, 481, 903, 54, 540, 418, 167, 672, 707, 17, 150, 22, 825, 717, 881, 791, 232, 963, 21, 907, 45, 101, 463, 949, 990, 155, 623, 228, 79, 689, 11, 794, 557, 923, 822, 686, 86, 476, 484, 497, 800, 905, 349, 230, 984, 471, 578, 798, 474, 588, 815, 172, 556, 841, 147, 662, 785, 264]. Sakamoto [884]. Saleh [408]. Salehi [486]. Salsburg [798]. Salt [886]. Sambandham [731]. Sameness [936]. Sampford [361]. Sample [797, 379, 434, 611, 91]. Samples [754]. Sampling [606, 90, 8, 560, 355, 521, 168, 876, 641, 421, 568, 679, 91, 131, 726, 541, 359, 328]. Sandon [156]. Sanford [407]. Sarndal [796]. Sator [889]. Savage [262]. Scaling [297, 468, 257, 468]. Schilling [355]. Schneider [754]. School [637, 974]. Schools [505, 640]. Schriever [719]. Schubert [78]. Schulmeister [21]. Schwartz [229]. Science [512, 724, 918, 141, 659, 645, 745, 948, 792, 51, 152]. Sciences [481, 776, 792, 876, 438, 877, 445, 685, 78, 180, 288, 49, 689, 409]. Scientific [90, 918, 11, 266]. Scientist [935]. Scientists [623, 292, 301, 841]. Scoring [312]. Scotland [639]. Scott [345, 904]. Scottish [163]. Screening [131]. Seal [562]. Sealy [177]. Searle [923]. Seasonal [9, 63, 207, 209, 203]. Seat [653, 806]. Seber [153, 539]. Second [678, 779]. Secondary [351]. Sector [698]. Sedransk [588]. Seidenfeld [40]. Selected [558, 50, 713, 962]. Selection [708, 196, 452, 262, 831]. Selective [808]. Sellars [51]. Seminar [717, 174, 791, 942]. Sen [669, 715, 496]. Seneta [356]. Sengupta [965]. September [779]. Sequence [934]. Sequences [537, 65, 237].

Sequential [392, 658, 196, 721]. Serfling [267]. Series [761, 326, 430, 687, 653, 444, 990, 718, 207, 622, 867, 343, 579, 597, 649, 214, 252, 253, 265, 292, 301, 484, 535, 553, 664, 673, 751, 534, 556, 170, 455, 125, 146, 16, 495, 203, 891, 71, 72, 142, 200, 252, 480]. Server [325]. Service [316, 120]. Services [372, 23, 335, 740, 705, 238, 657, 263]. Servicing [504]. Session [980, 773]. Set [931, 136]. Sets [628]. Seventeenth [745]. Seventh [192]. Sexual [769]. Shapes [786]. Shapiro [540]. Sheila [473]. Sheldon [522]. Shifts [679]. Shimizu [943]. Shipbuilding [878]. Shiryaev [75]. Shiryayev [589]. Shorack [799]. Short [584, 21, 768, 340, 417]. Short-Run [340]. Short-Term [21, 768]. Shubert [180]. Sickness [594]. Siegel [389]. Sigmoids [853]. Signal [63, 559]. Signals [180]. Significance [453]. Significant [936]. Silverman [631, 755]. Silvey [230]. Simar [579]. Simon [784, 634, 19]. Simplified [764]. Simulating [35]. Simulation [830, 882, 586, 787, 330]. since [985]. Singer [50, 712]. Singh [796]. Single [325]. Sint [598]. Sir [30, 305, 270, 552, 423]. Sixteenth [659]. Sixth [94]. Sixty [406]. Sixty-Fifth [406]. Size [797, 123, 370]. Sizes [611]. Sketch [568]. Skills [27]. Skipworth [201]. Slower [487]. Small [796]. Smith [885, 734, 290, 779, 559, 948]. Smoking [957]. Smoothed [848]. Smoothing [906]. Social [351, 372, 81, 344, 48, 335, 740, 288, 438, 141, 770, 657, 207, 83, 292, 301, 742, 685, 20, 985, 266]. Society [761, 857, 898, 938, 981, 812, 427, 231, 94, 93, 192, 193, 283, 380, 318, 381, 530, 529, 618, 571, 617, 684, 646, 758, 975, 427, 428, 449]. Socio [377]. Socio-Economic [377]. Software [928, 127, 691]. Solomon [871, 871]. Solutions [816]. Some [443, 431, 442, 446, 440, 430, 438, 851, 444, 806, 436, 118, 439, 429, 432, 257, 894, 445, 433, 434, 437, 647, 369]. Someone [7]. Soto [100]. Source [979]. Sources [351, 334, 878, 836, 914, 67, 703, 166, 913]. Spaces [844]. Spath [605]. Spatial [549, 678, 70, 143, 705, 867, 109, 671, 466, 494, 867, 863, 71, 145]. Special [577, 93, 193, 318, 381, 529, 571, 617]. Specifying [411]. Spectral [885, 891, 252]. Spheres [497]. Spline [906]. Spohn [384]. Sprent [924, 268]. SPSS [951]. SPSS/PC [951]. Square [946]. Squared [62]. Squares [950, 523]. Sri [108]. Srivastava [45]. Stablein [513]. Stafford [331]. Stahel [732]. Starting [613]. State [844, 506]. Statement [773]. States [348]. Static [716]. Stationary [495, 74, 484]. Statistic [164, 2, 185, 165]. Statistical [886, 955, 190, 551, 637, 94, 93, 192, 193, 283, 372, 380, 318, 381, 530, 529, 618, 571, 617, 684, 646, 758, 761, 878, 828, 262, 836, 481, 399, 549, 790, 189, 860, 29, 914, 560, 176, 300, 825, 752, 717, 776, 174, 342, 881, 934, 661, 879, 975, 440, 335, 740, 792, 143, 512, 555, 724, 367, 278, 67, 876, 517, 741, 127, 427, 979, 928, 238, 141, 666, 744, 576, 733, 772, 973, 106, 665, 378, 691, 819, 61, 40, 703, 570, 877, 191, 118, 710, 689, 429, 144, 794, 353, 169, 490, 428]. Statistical [655, 257, 323, 894, 449, 574, 354, 913, 411, 398, 709, 746, 223, 387, 986, 469, 121, 787, 427, 23, 832, 538, 18, 890, 514, 129, 577, 580, 757, 51, 623, 603, 246, 291, 459, 559, 321, 734, 14, 474, 26, 709, 488, 202, 830, 814, 833]. Statistician [15, 276]. Statisticians [646, 279, 646, 472]. Statistics [334, 761, 857, 898, 938, 981, 376, 443, 774, 756, 416, 199, 566, 431, 812, 250, 868, 640, 160, 630, 959, 446, 801, 377, 791, 275, 724, 181, 324, 438, 779, 775, 448, 732, 902, 464, 251, 163, 811, 629, 12, 52, 935, 659, 439, 4, 645, 432, 11, 144, 871, 745, 992, 338, 178, 295, 480, 482, 535, 445, 25, 796, 302, 496, 437, 933, 677, 841, 602, 268, 369, 696, 755, 44, 68, 262, 507, 919, 130, 713, 866, 84, 884, 817, 28, 177, 838, 48, 46, 492, 920, 82, 77, 267, 584, 405, 926].

Statistics

[248, 802, 80, 171, 215, 45, 949, 27, 799, 733, 900, 494, 837, 952, 219, 476, 598, 953, 588, 962, 467, 173, 413, 753, 721, 201, 816, 558, 54, 249, 225, 927, 408, 464, 473, 497, 798, 266]. Statistics. [49]. Statistique [783]. Statistische [818]. Statutory [59]. Steel [878]. Stegmüller [384]. Step [631]. Stephens [782]. Stereology [402]. Steven [319]. Stigler [756]. Stirling [886]. Stiteler [226]. Stochastic [876, 942, 391, 544, 877, 128, 78, 344, 859, 410, 844, 882, 227, 352, 690, 800]. Stoppages [314]. Stopping [75]. Storage [227, 859]. Store [968, 968]. Stoyan [800]. Stoyanov [925]. Straszak [43]. Strategies [850]. Strathclyde [550]. Strigel [269]. Strike [123, 614]. Strikes [10, 454]. Structural [704, 653, 861]. Structure [549, 17, 370, 594, 985]. Structures [989, 912, 958]. Stuart [801, 480, 558, 541, 178]. Student [592]. Students [49]. Studies [637, 85, 514, 246, 84]. Study [653, 240, 921, 628, 917, 231, 100]. Stuvel [842]. Style [612]. Styles [190]. Subba [718]. Subjective [803]. Subrahmaniam [53]. Subsets [452]. Suicide [38]. Sukhatme [420]. Sunspot [371]. Supply [646, 550]. Support [398]. Surfaces [780]. Survey [313, 904, 181, 583, 62, 208, 403, 616, 595, 919, 568, 91, 551, 359, 14]. Surveying [803]. Surveys [727, 92, 434, 83, 269, 20, 883, 839, 103]. Survival [513]. Symposium [659, 745, 796]. Syski [111]. System [189, 741, 44]. Systems [368, 575, 211, 398, 556, 316, 70, 986, 180]. Szekely [802].

T [465, 540, 604, 13, 717, 731, 99, 718, 623, 40, 864, 783, 875, 299, 952, 822, 460, 586, 321, 574, 746, 202]. T.-P [465]. Table [506]. Tables [713, 521, 786, 60, 962, 833, 453, 131].

Tabulation [583]. Takeuchi [342]. Tamhane [874]. Tampere [717]. Tan [843]. Tashman [82]. Tatham [873]. Tauberian [967]. Tax [124]. Taxonomy [349]. Taylor [720]. Teaching [886, 190, 900, 338, 120, 791]. Technical [704]. Techniques [600, 555, 136, 782, 785, 167]. Technology [464, 645]. Telephone [727, 103]. Temperature [649]. Temporal [145]. Ten [23]. Tensor [837]. Term [21, 768, 417]. Terms [225, 897]. Testing [18, 642, 790, 175]. Tests [457, 521, 62, 169, 64, 210, 453, 946, 961, 905, 147]. Textbook [584]. texts [468]. Th [353]. Their [828, 85, 717, 420, 974, 903, 949, 575]. Theology [774]. Theorem [386]. Theorems [967, 267]. Theoretical [823, 492]. Theories [685, 105]. Theory [294, 955, 667, 960, 558, 839, 542, 319, 153, 160, 881, 670, 801, 440, 605, 404, 139, 517, 802, 80, 415, 943, 666, 603, 475, 622, 388, 784, 871, 169, 343, 178, 480, 553, 321, 230, 984, 474, 534, 986, 121, 486, 456, 599, 600, 604, 711, 801, 876, 944, 781, 601, 74, 214, 862, 966, 289, 518, 387]. Therapeutic [390]. Third [642]. Thomas [785, 417, 542, 477, 242]. Thompson [887, 667]. Thought [11]. Three [457]. Thrift [70]. Thrombosis [743]. Tijms [690]. Tiku [843, 733]. Time [600, 767, 125, 326, 430, 16, 391, 544, 687, 653, 444, 990, 718, 203, 207, 867, 579, 597, 891, 649, 71, 72, 200, 252, 301, 480, 484, 535, 664, 673, 170, 455, 23, 495, 175, 687, 222, 622, 343, 142, 214, 265, 292, 553, 751, 534, 556]. Time-Series [444, 495]. Tintner [546]. Tippett [244, 638]. Titterington [734]. Toby [576]. Tocher [307]. Today [348]. Tolerance [131]. Tong [182]. Tool [102]. Tools [830]. Topic [441, 435]. Topics [869, 13, 99, 872, 474, 839, 408]. Total [317, 211]. Tourism [21, 21]. Toutenberg [357]. Townsend [27]. Toxicologists [798]. Trade [897, 69, 245]. Traditions [739].

Training [337]. Tran [987]. Transect [168]. Transformations [901, 706]. Transmission [763]. Treasurer [96, 194, 285, 382, 532, 620, 682, 729, 855]. Treatment [38, 528]. Trend [207]. Trends [895, 786, 479, 476, 985, 64]. Trial [390, 611, 210]. Trials [793, 376, 540, 378, 557, 59, 392, 229, 833]. Triangle [947]. Tribus [47]. Tribute [364, 244, 243, 365, 552]. Truncated [754]. Tsurumi [861]. Tuckwell [966]. Tufte [543]. Tukey [597, 414, 786, 463]. Turner [803]. Twenty [791]. Two [969]. Tye [791]. Tygstrup [390]. Types [505]. Tzannatos [735].

U [327]. UK [189, 976, 259]. Ullah [358]. Umphrey [792, 876, 877, 751]. Unbalanced [851, 923]. Uncertainty [756, 944, 864]. Unconditional [611]. Underestimating [161]. Undergraduates [742]. Understanding [566, 918, 403, 809, 414, 924]. Unemployed [36]. Unemployment [551, 594]. Unequal [679]. Unified [138, 493]. United [348, 878, 836, 914, 879, 67, 913, 773, 703, 59]. Univariate [252]. Universities [742]. University [174]. Unknown [631]. Unlisted [727]. untitled [834, 835, 511, 510]. Up-Dating [312]. Updating [7]. Uppuluri [864]. Upton [671]. Urban [100]. Use [569, 312, 698, 931, 509, 853, 726]. Used [979]. User [507, 468, 301]. Uses [974]. Using [259, 455, 411]. Utilising [240]. Utilitas [478]. Utility [287].

V [967, 668, 84, 868, 538, 402, 963, 350, 779, 215, 420, 79, 864, 227, 295, 462, 486, 264, 107]. Valencia [779]. Validation [39]. Validity [379]. Values [777]. Variable [769, 528, 578]. Variables [452, 78, 778]. Variance [868, 964, 250]. Variation [678, 258]. Variations [697, 282]. Varieties [810, 59]. Variety [179]. Vasić [903]. Versus [401, 615, 654]. Veterinary [51]. VI [714]. Vic [576, 322, 323]. Vicinity [976, 977]. Viertl [926]. View [776, 935, 627]. Views [443, 431, 442, 446, 440, 430, 438, 444, 436, 439, 429, 432, 445, 433, 434, 437]. Vinod [358]. Visual [725, 543]. Vital [317, 211]. Vlidimirov [967]. Vogt [889]. Vol [481, 914, 17, 150, 13, 12, 689, 246, 178, 252, 480, 482, 535, 320, 296, 913, 709, 180]. Volume [600, 88, 187, 273, 374, 525, 608, 675, 737, 846, 971, 878, 836, 250, 671, 879, 536, 801, 792, 67, 876, 99, 251, 877, 871, 597, 78, 253]. Volumes [478, 24, 598, 777]. Voting [140]. Vrieze [839, 844].

W [294, 886, 878, 713, 836, 950, 947, 23, 29, 914, 516, 596, 660, 414, 149, 224, 843, 226, 879, 889, 786, 887, 146, 954, 67, 148, 171, 218, 463, 631, 928, 732, 900, 691, 384, 299, 409, 325, 992, 219, 597, 800, 831, 14, 574, 913, 746, 755, 662, 599]. W. [588]. Wage [44]. Wages [735]. Wales [639, 647, 435]. Walks [74]. Walter [513]. Wampler [513]. Waring [456]. Warren [580]. Warwick [555]. Wassily [169]. Water [441, 550]. Waterways [67]. Watson [497]. Wayne [409]. Ways [82]. Weeda [154]. Wegman [757, 559]. Weights [807]. Weisberg [672, 232, 407]. Wellner [799]. Welsch [166]. Werner [269]. Wertz [14]. Western [487]. Wetherill [353, 721, 300]. Wetz [401, 615]. Wheatley [27]. Wheelwright [86]. Where [448]. White [421, 498, 785]. Whitehead [392]. Whittemore [418]. Whittle [391, 544, 523]. Wickens [24]. Wilburn [560]. Wilheim [563]. Willemain [202]. William [129, 49, 403, 467, 304]. Williams [301, 633]. Wilrich [353]. Wilson [118]. Winkler [563]. within [698, 897, 726]. Witness [276]. Wold [404]. Wolf [371]. Wolfe [80]. Wolfer [371]. Women [248, 735]. Wood [101]. Woodward [927, 928, 691]. Woolson [890]. Work [226]. Worked [201]. Working [646]. Works [597]. World [904, 91, 487]. Worthington [961]. Wrigley [968]. Writings [262]. Wu [556].

X [67, 209]. X-11 [209]. XIX [836]. XVI [878]. XX [913]. XXI [914]. XXII [879].

Yadolah [748]. Yates [478, 243, 359]. Year [23]. Yearly [649]. Years [427, 447, 428, 807]. Yook [791]. Young [673]. Youngman [81]. Ysander [969]. Yu [967].

Z [735]. Zabalza [735]. Zavialov [967]. Zellner [203, 302]. Zucchini [831]. Zurbenko [891]. Zurich [371]. Zvi [750, 536].

References

Moran:1947:RDI

[1] P. A. P. Moran. The random division of an interval. Supplement to the Journal of the Royal Statistical Society, 9(1):92– 98, ???? 1947. CODEN ???? ISSN

                                             1466-6162. URL http://www.jstor.
                                             org/stable/2983572. See corrigendum

[184].

Hill:1979:ADG

[2] I. D. Hill. Approximating the distribution of Greenwood’s statistic with Johnson distributions. Journal of the Royal Statistical Society. Series A (General), 142(3):378–380, ???? 1979. CODEN JS-SAEF. ISSN 0035-9238. URL http:/

                                             /www.jstor.org/stable/2982489. See

corrigenda [185].

Anonymous:1980:FMa

[3] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 143(1):??, ???? 1980. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981764.

Moser:1980:SPP

[4] Claus Moser. Statistics and public policy. Journal of the Royal Statistical Society. Series A (General), 143(1):1–32, ???? 1980. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981765.

Anderson:1980:GRC

[5] L. R. Anderson and R. A. Fontenot. On the gain ratio as a criterion for betting in casino games. Journal of the Royal Statistical Society. Series A (General), 143(1):33–40, ???? 1980. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981766.

Downton:1980:MGR

[6] F. Downton. Monotonicity and the gain ratio. Journal of the Royal Statistical Society. Series A (General), 143(1):41– 42, ???? 1980. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981767.

French:1980:UBL

[7] Simon French. Updating of belief in the light of someone Else’s opinion. Journal of the Royal Statistical Society. Series A (General), 143(1):43–48, ???? 1980. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981768.

Bravo:1980:MSP

[8] P. C. Bravo and G. B. Wetherill. The matching of sampling plans and the design of double sampling plans. Journal of the Royal Statistical Society. Series A (General), 143(1):49–67, ???? 1980. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981769.

Balk:1980:MCP

[9] B. M. Balk. A method for constructing price indices for seasonal commodities. Journal of the Royal Statistical Society. Series A (General), 143(1):68–75, ???? 1980. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981770.

Geary:1980:PS

[10] R. C. Geary. Prais on strikes. Journal of the Royal Statistical Society. Series A (General), 143(1):76–77, ???? 1980. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981771.

Newman:1980:BRB

[11] Dennis Newman. Book review: The History of Statistics in the 17th and 18th Centuries, Against the Changing Background of Intellectual, Scientific and Religious Thought: Lectures by Karl Pearson, 1921–1933, by E. S. Pearson. Journal of the Royal Statistical Society. Series A (General), 143(1):78–79, ???? 1980. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981772.

Mardia:1980:BRB

[12] K. V. Mardia. Book review: Developments in Statistics, Vol. 2, by P. R. Krishnaiah. Journal of the Royal Statistical Society. Series A (General), 143(1): 79, ???? 1980. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981773.

Coleman:1980:BRB

[13] Rodney Coleman. Book review: Probabilistic Analysis and Related Topics, Vol. 1, by A. T. Bharucha-Reid. Journal of the Royal Statistical Society. Series A (General), 143(1):79–80, ???? 1980. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981774.

Silverman:1980:BRB

[14] B. W. Silverman. Book review: Statistical Density Estimation: A Survey, by W. Wertz. Journal of the Royal Statistical Society. Series A (General), 143(1): 80, ???? 1980. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981775.

Chatfield:1980:BRB

[15] C. Chatfield. Book review: Forecasting. Proceedings of the Institute of Statisticians’ Annual Conference at Cambridge in 1976, by O. D. Anderson. Journal of the Royal Statistical Society. Series A (General), 143(1):80–81, ???? 1980. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981776.

Gregg:1980:BRB

[16] D. P. Gregg. Book review: Practical Experiences with Modelling and Forecasting Time Series, by G. M. Jenkins. Journal of the Royal Statistical Society. Series A (General), 143(1):81–82, ???? 1980. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2981777.

Clarke:1980:BRBa

[17] G. M. Clarke. Book review: Analysing Qualitative/Categorical Data: Loglinear Models and Latent Structure Analysis, by L. A. Goodman; J. A. Davis; J. Magidson; Analysis of Qualitative Data. Vol. I, by S. J. Haberman. Journal of the Royal Statistical Society. Series A (General), 143(1):82–83, ???? 1980. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2981778.

Downton:1980:BRB

[18] F. Downton. Book review: Statistical Analysis of Reliability and Life Testing Models, by Lee J. Bain. Journal of the Royal Statistical Society. Series A (General), 143(1):83–84, ???? 1980. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2981779.

Ermisch:1980:BRB

[19] J. F. Ermisch. Book review: Research in Population Economics: An Annual Compilation of Research, by J. L. Simon. Journal of the Royal Statistical Society. Series A (General), 143(1):84– 85, ???? 1980. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2981780.

Spitz:1980:BRBa

[20] K. E. Spitz. Book review: The Recall Method in Social Surveys, by Louis Moss; Harvey Goldstein. Journal of the Royal Statistical Society. Series A (General), 143(1):85, ???? 1980. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2981781.

Gibson:1980:BRB

[21] J. G. Gibson. Book review: Tourism and the Business Cycle: Econometric Models for the Purpose of Analysis and Forecasting of Short-Term Changes in the Demand for Tourism, by S. Schulmeister. Journal of the Royal Statistical Society. Series A (General), 143(1):85–86, ???? 1980. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981782.

Cohen:1980:BRB

[22] G. Cohen. Book review: Issues in Demographic Data Collection in Pakistan, by K. J. Krotki; S. S. Hashmi. Journal of the Royal Statistical Society. Series A (General), 143(1):86–87, ???? 1980. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981783.

Bowyer:1980:BRB

[23] Adrian Bowyer. Book review: Statistical Services in Ten Years’ Time, by J. W. Duncan. Journal of the Royal Statistical Society. Series A (General), 143(1): 87, ???? 1980. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981784.

Harvey:1980:BRB

[24] A. C. Harvey. Book review: Exercises in Econometrics, Volumes 1 and 2, by P. C. B. Phillips; M. R. Wickens. Journal of the Royal Statistical Society. Series A (General), 143(1):87–88, ???? 1980. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981785.

Roberts:1980:BRB

[25] Carole A. Roberts. Book review: Basic Statistics: A Modern Approach, 2nd edition, by Morris Hamburg. Journal of the Royal Statistical Society. Series A (General), 143(1):88–89, ???? 1980. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2981786.

Stone:1980:BRB

[26] M. Stone. Book review: Statistical Analysis: A Decision-Making Approach, 2nd edition, by R. Parsons. Journal of the Royal Statistical Society. Series A (General), 143(1):89, ???? 1980. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2981787.

Holmes:1980:BRB

[27] P. Holmes. Book review: Developing Skills in Statistics, by N. R. Townsend; G. H. Wheatley. Journal of the Royal Statistical Society. Series A (General), 143(1):89–90, ???? 1980. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2981788.

Chedzoy:1980:BRB

[28] O. B. Chedzoy. Book review: Introductory Statistics, by A. M. Plews. Journal of the Royal Statistical Society. Series A (General), 143(1):90, ???? 1980. CODEN JSSAEF. ISSN 0035-9238.

                                        URL http://www.jstor.org/stable/
                                        2981789.

Brook:1980:BRBa

[29] D. Brook. Book review: Elementary Statistical Quality Control, by I. W. Burr. Journal of the Royal Statistical Society. Series A (General), 143(1):91, ???? 1980. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981790.

Yame:1980:PSA

[30] B. S. Yame. Professor Sir Arnold Plant, 1898–1978. Journal of the Royal Statistical Society. Series A (General), 143(1): 92, ???? 1980. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981791.

Finney:1980:CIB

[31] D. J. Finney. Chester Ittner Bliss, 1899– 1979. Journal of the Royal Statistical Society. Series A (General), 143(1):92–93, ???? 1980. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981792.

Anonymous:1980:RPa

[32] Anonymous. Recent periodicals. Journal of the Royal Statistical Society. Series A (General), 143(1):94–108, ???? 1980. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981793.

Anonymous:1980:BMa

[33] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 143(1):??, ???? 1980. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981794.

Anonymous:1980:FMb

[34] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 143(2):??, ???? 1980. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981984.

Broadbent:1980:SLH

[35] Simon Broadbent. Simulating the Ley Hunter. Journal of the Royal Statistical Society. Series A (General), 143(2):109– 140, ???? 1980. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981985.

Lancaster:1980:ARE

[36] Tony Lancaster and Stephen Nickell. The analysis of re-employment probabilities for the unemployed. Journal of the Royal Statistical Society. Series A (General), 143(2):141–165, ???? 1980. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981986.

Anonymous:1980:OP

[37] Anonymous. Opinion polls. Journal of the Royal Statistical Society. Series A (General), 143(2):166, ???? 1980. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981987.

Copas:1980:DES

[38] J. B. Copas and M. J. Fryer. Density estimation and suicide risks in psychiatric treatment. Journal of the Royal Statistical Society. Series A (General), 143(2):167–176, ???? 1980. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981988.

Clark:1980:CCV

[39] R. M. Clark. Calibration, crossvalidation and carbon-14. II. Journal of the Royal Statistical Society. Series A (General), 143(2):177–194, ???? 1980. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981989.

Marriott:1980:BRBa

[40] F. H. C. Marriott. Book review: Philosophical Problems of Statistical Inference. Learning from R. A. Fisher, by T. Seidenfeld. Journal of the Royal Statistical Society. Series A (General), 143(2): 195, ???? 1980. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981990.

Brown:1980:BRB

[41] A. J. Brown. Book review: Causality in Economics, by John Hicks. Journal of the Royal Statistical Society. Series A (General), 143(2):196, ???? 1980. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981991.

Wilkinson:1980:BRBa

[42] Margaret Wilkinson. Book review: The Economics of Modern Britain. An Introduction to Macroeconomics, by John Black. Journal of the Royal Statistical Society. Series A (General), 143(2):196– 197, ???? 1980. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981992.

Yates:1980:BRB

[43] Judith Yates. Book review: Models and Decision Making in National Economies, by J. M. L. Janssen; L. F. Pau; A. Straszak. Journal of the Royal Statistical Society. Series A (General), 143(2):197–198, ???? 1980. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981993.

Wilkinson:1980:BRBb

[44] R. K. Wilkinson. Book review: An Integrated System of Wage Statistics. A Manual on Methods. Journal of the Royal Statistical Society. Series A (General), 143(2):198, ???? 1980. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2981994.

Healy:1980:BRBa

[45] M. J. R. Healy. Book review: An Introduction to Multivariate Statistics, by M. S. Srivastava; C. G. Khatri. Journal of the Royal Statistical Society. Series A (General), 143(2):198–199, ???? 1980. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981995.

Curnow:1980:BRB

[46] R. N. Curnow. Book review: Basic Applied Statistics, by B. L. Raktoe; J. J. Hubert. Journal of the Royal Statistical Society. Series A (General), 143(2): 199, ???? 1980. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981996.

Bernardo:1980:BRB

[47] J. M. Bernardo. Book review: The Maximum Entropy Formalism, by R. D. Levine; M. Tribus. Journal of the Royal Statistical Society. Series A (General), 143(2):199–200, ???? 1980. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981997.

Conway:1980:BRB

[48] Freda Conway. Book review: Demystifying Social Statistics, by John Irvine. Journal of the Royal Statistical Society. Series A (General), 143(2):200–201, ???? 1980. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981998.

Haigh:1980:BRBa

[49] John Haigh. Book review: Statistics. For Students in the Behavioural Sciences, by William A. Lindner. Journal of the Royal Statistical Society. Series A (General), 143(2):201–202, ???? 1980. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981999.

Haigh:1980:BRBb

[50] John Haigh. Book review: Distribution-Free Methods for Non-Parametric Problems: A Classified and Selected Bibliography, by Bernard Singer. Journal of the Royal Statistical Society. Series A (General), 143(2):202, ???? 1980. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2982000.

Hunter:1980:BRB

[51] E. A. Hunter. Book review: Statistical Epidemiology in Veterinary Science, by F. B. Leech; K. C. Sellars. Journal of the Royal Statistical Society. Series A (General), 143(2):202–203, ???? 1980. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982001.

Marriott:1980:BRBb

[52] F. H. C. Marriott. Book review: Advanced Level Statistics: An Integrated Course, by A. Francis. Journal of the Royal Statistical Society. Series A (General), 143(2):203, ???? 1980. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2982002.

Clarke:1980:BRBb

[53] L. E. Clarke. Book review: A Primer in Probability, by Kathleen Subrahmaniam. Journal of the Royal Statistical Society. Series A (General), 143(2):203– 204, ???? 1980. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2982003.

Brook:1980:BRBb

[54] D. Brook. Book review: Statistics: Concepts and Controversies, by D. S. Moore. Journal of the Royal Statistical Society. Series A (General), 143(2):204, ???? 1980. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2982004.

Armitage:1980:JC

[55] Peter Armitage. Jerome Cornfield, 1912–1979. Journal of the Royal Statistical Society. Series A (General), 143(2): 205, ???? 1980. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2982005.

Anonymous:1980:RPb

[56] Anonymous. Recent periodicals. Journal of the Royal Statistical Society. Series A (General), 143(2):206–218, ???? 1980. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2982006.

Anonymous:1980:BMb

[57] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 143(2):??, ???? 1980. CODEN JSSAEF. ISSN 0035-9238.

                                        URL http://www.jstor.org/stable/
                                        2982007.

Anonymous:1980:FMc

[58] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 143(3):??, ???? 1980. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2982127.

Patterson:1980:SRL

[59] H. D. Patterson and Valerie Silvey. Statutory and recommended list trials of crop varieties in the United Kingdom. Journal of the Royal Statistical Society. Series A (General), 143(3):219– 252, ???? 1980. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2982128.

Payne:1980:IKD

[60] R. W. Payne and D. A. Preece. Identification keys and diagnostic tables: A review. Journal of the Royal Statistical Society. Series A (General), 143(3):253– 292, ???? 1980. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2982129.

Mardia:1980:SAM

[61] K. V. Mardia and Dorothy Holmes. A statistical analysis of megalithic data under elliptic pattern. Journal of the Royal Statistical Society. Series A (General), 143(3):293–302, ???? 1980. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2982130.

Holt:1980:CST

[62] D. Holt, A. J. Scott, and P. D. Ewings. Chi-squared tests with survey data. Journal of the Royal Statistical Society. Series A (General), 143(3):303–320, ???? 1980. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2982131.

Burman:1980:SAS

[63] J. P. Burman. Seasonal adjustment by signal extraction. Journal of the Royal Statistical Society. Series A (General), 143(3):321–337, ???? 1980. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982132.

Taylor:1980:CMT

[64] Stephen J. Taylor. Conjectured models for trends in financial prices, tests and forecasts. Journal of the Royal Statistical Society. Series A (General), 143 (3):338–362, ???? 1980. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982133.

Downton:1980:NLS

[65] F. Downton. A note on Labouchere sequences. Journal of the Royal Statistical Society. Series A (General), 143 (3):363–366, ???? 1980. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982134.

Webb:1980:NAF

[66] Norman L. Webb. A note on the accuracy of forecasting: Opinion polls. Journal of the Royal Statistical Society. Series A (General), 143(3):367–368, ???? 1980. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982135.

Flowerdew:1980:BRB

[67] A. D. J. Flowerdew. Book review: Reviews of United Kingdom Statistical Sources: Volume X: Ports and Inland Waterways, by W. F. Maunder; R. E. Baxter; Reviews of United Kingdom Statistical Sources.: Civil Aviation, by W. F. Maunder; Celia M. Phillips. Journal of the Royal Statistical Society. Series A (General), 143(3):369–370, ???? 1980. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982136.

Wilkinson:1980:BRBc

[68] R. K. Wilkinson. Book review: British Regional Employment Statistics, 1841– 1971, by C. H. Lee. Journal of the Royal Statistical Society. Series A (General), 143(3):370, ???? 1980. CODEN JS-SAEF. ISSN 0035-9238. URL http:/

/www.jstor.org/stable/2982137.

Meyer:1980:BRB

[69] F. V. Meyer. Book review: A Geometric Approach to International Trade, by M. B. Krauss. Journal of the Royal Statistical Society. Series A (General), 143 (3):370–371, ???? 1980. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982138.

Diggle:1980:BRB

[70] P. J. Diggle. Book review: Towards the Dynamic Analysis of Spatial Systems, by R. L. Martin; N. J. Thrift; R. J. Bennett. Journal of the Royal Statistical Society. Series A (General), 143(3):371, ???? 1980. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2982139.

Ray:1980:BRBa

[71] W. D. Ray. Book review: Spatial Time Series, by R. J. Bennett. Journal of the Royal Statistical Society. Series A (General), 143(3):371–372, ???? 1980. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2982140.

Ray:1980:BRBb

[72] W. D. Ray. Book review: Analysis of Economic Time Series, by Marc Nerlove; David M. Grether; Jose L. Carvalho. Journal of the Royal Statistical Society. Series A (General), 143(3):372, ???? 1980. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2982141.

Bingham:1980:BRB

[73] N. H. Bingham. Book review: Probability and Measure, by P. Billingsley. Journal of the Royal Statistical Society. Series A (General), 143(3):372–373, ???? 1980. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982142.

Kingman:1980:BRB

[74] J. F. C. Kingman. Book review: Random Walks with Stationary Increments and Renewal Theory, by H. C. P. Berbee. Journal of the Royal Statistical Society. Series A (General), 143(3):373, ???? 1980. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982143.

Bather:1980:BRB

[75] J. A. Bather. Book review: Optimal Stopping Rules, by A. N. Shiryaev. Journal of the Royal Statistical Society. Series A (General), 143(3):373–374, ???? 1980. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982144.

Shanbhag:1980:BRB

[76] D. N. Shanbhag. Book review: ‘Classifying Infinitely Divisible Distributions, by Functional Equations. by K. Van Harn. Journal of the Royal Statistical Society. Series A (General), 143(3):374, ???? 1980. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2982145.

Eplett:1980:BRB

[77] W. J. R. Eplett. Book review: Contributions to Statistics, by Jana Jureckova. Journal of the Royal Statistical Society. Series A (General), 143(3):374– 375, ???? 1980. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2982146.

Wright:1980:BRB

[78] C. K. Wright. Book review: Probabilistic Models in the Engineering Sciences, Volume 1. Random Variables and Stochastic Processes, by H. J. Larson; B. O. Schubert. Journal of the Royal Statistical Society. Series A (General), 143 (3):375–376, ???? 1980. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982147.

Matatko:1980:BRB

[79] John Matatko. Book review: Estimation Risk and Optimal Portfolio Choice, by V. S. Bawa; S. J. Brown; R. Klein. Journal of the Royal Statistical Society. Series A (General), 143(3):376, ???? 1980. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982148.

Haigh:1980:BRBc

[80] John Haigh. Book review: Introduction to the Theory of Nonparametric Statistics, by Ronald H. Randles; Douglas A. Wolfe. Journal of the Royal Statistical Society. Series A (General), 143(3):376– 377, ???? 1980. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2982149.

Bartholomew:1980:BRB

[81] D. J. Bartholomew. Book review: Analysing Social and Educational Research Data, by M. B. Youngman. Journal of the Royal Statistical Society. Series A (General), 143(3):377–378, ???? 1980. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982150.

Ehrenberg:1980:BRB

[82] A. S. C. Ehrenberg. Book review: The Ways and Means of Statistics, by L. J. Tashman; K. Lamborn. Journal of the Royal Statistical Society. Series A (General), 143(3):378, ???? 1980. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2982151.

Plewis:1980:BRB

[83] Ian Plewis. Book review: Methodological Issues in Social Surveys, by Mavis Maclean; Hazel Genn. Journal of the Royal Statistical Society. Series A (General), 143(3):378–379, ???? 1980. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982152.

Barnett:1980:BRB

[84] Vic Barnett. Book review: Studies in Statistics, by R. V. Hogg. Journal of the Royal Statistical Society. Series A (General), 143(3):379–380, ???? 1980. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982153.

Clayton:1980:BRB

[85] D. G. Clayton. Book review: The Design and Analysis of Longitudinal Studies: Their Role in The Measurement of Change, by H. Goldstein. Journal of the Royal Statistical Society. Series A (General), 143(3):380–381, ???? 1980. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982154.

Ray:1980:BRBc

[86] W. D. Ray. Book review: Forecasting, by S. Makridakis; S. C. Wheelwright. Journal of the Royal Statistical Society. Series A (General), 143(3):381, ???? 1980. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2982155.

Anonymous:1980:BMc

[87] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 143(3):i–iii, ???? 1980. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2982156.

Anonymous:1980:VI

[88] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (General), 143(4):549–552, ???? 1980. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982061.

Anonymous:1980:FMd

[89] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 143(4):??, ???? 1980. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2982062.

Box:1980:SBI

[90] George E. P. Box. Sampling and Bayes’ inference in scientific modelling and robustness. Journal of the Royal Statistical Society. Series A (General), 143 (4):383–430, ???? 1980. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982063.

Verma:1980:SDS

[91] Vijay Verma, Christopher Scott, and Colm O’Muircheartaigh. Sample designs and sampling errors for the world fertility survey. Journal of the Royal Statistical Society. Series A (General), 143 (4):431–473, ???? 1980. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2982064.

Holt:1980:RAD

[92] D. Holt, T. M. F. Smith, and P. D. Winter. Regression analysis of data from complex surveys. Journal of the Royal Statistical Society. Series A (General), 143(4):474–487, ???? 1980. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2982065.

Anonymous:1980:PSG

[93] Anonymous. Proceedings of a Special General Meeting of the Royal Statistical Society. Journal of the Royal Statistical Society. Series A (General), 143 (4):488–489, ???? 1980. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2982066.

Anonymous:1980:POH

[94] Anonymous. Proceedings of the one hundred and forty-sixth Annual General Meeting of the Royal Statistical Society. Journal of the Royal Statistical Society. Series A (General), 143(4):489–490, ???? 1980. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2982067.

Anonymous:1980:PR

[95] Anonymous. President’s report. Journal of the Royal Statistical Society. Series A (General), 143(4):491–493, ????

1980. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982068.

Anonymous:1980:HTR

[96] Anonymous. Honorary Treasurer’s report. Journal of the Royal Statistical Society. Series A (General), 143(4):494, ???? 1980. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2982069.

Anonymous:1980:ARC

[97] Anonymous. Annual report of the Council. Journal of the Royal Statistical Society. Series A (General), 143(4):495–517, ???? 1980. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2982070.

Smith:1980:BRB

[98] Cedric A. B. Smith. Book review: The Geometry of Population Genetics, by Ethan Akin. Journal of the Royal Statistical Society. Series A (General), 143(4): 518, ???? 1980. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2982071.

Holgate:1980:BRB

[99] P. Holgate. Book review: Probabilistic Analysis and Related Topics. Volume 2, by A. T. Barucha-Reid. Journal of the Royal Statistical Society. Series A (General), 143(4):518–519, ???? 1980. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982072.

Wilkinson:1980:BRBd

[100] R. K. Wilkinson. Book review: The Good City. A Study of Urban Development and Policy in Britain, by David Domison; Paul Soto. Journal of the Royal Statistical Society. Series A (General), 143(4):519–520, ???? 1980. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2982073.

Healy:1980:BRBb

[101] M. J. R. Healy. Book review: Fitting Equations to Data, 2nd Ed, by C. Daniel; F. S. Wood. Journal of the Royal Statistical Society. Series A (General), 143(4):520, ???? 1980. CODEN JS-SAEF. ISSN 0035-9238. URL http:/

/www.jstor.org/stable/2982074.

Loynes:1980:BRB

[102] R. M. Loynes. Book review: Regression Methods: A Tool for Data Analysis, by Rudolf J. Freund; Paul D. Minton. Journal of the Royal Statistical Society. Series A (General), 143(4):520–521, ???? 1980. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982075.

Spitz:1980:BRBb

[103] K. E. Spitz. Book review: Surveys, by Telephone, A National Comparison with Personal Interviews. by Robert M. Groves; Robert L. Kahn. Journal of the Royal Statistical Society. Series A (General), 143(4):521–522, ???? 1980. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2982076.

Lewes:1980:BRB

[104] Fred Lewes. Book review: Regional Demographic Development, by John Hobcraft; Phillip Rees. Journal of the Royal Statistical Society. Series A (General), 143(4):522, ???? 1980. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2982077.

Wilkinson:1980:BRBe

[105] R. K. Wilkinson. Book review: Modern Theories of Income Distribution, by M. C. Howard. Journal of the Royal Statistical Society. Series A (General), 143(4):523, ???? 1980. CODEN JS-SAEF. ISSN 0035-9238. URL http:/

/www.jstor.org/stable/2982078.

Knott:1980:BRBa

[106] M. Knott. Book review: Probability and Statistical Inference, by J. G. Kalbfleisch. Journal of the Royal Statistical Society. Series A (General), 143 (4):523–524, ???? 1980. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982079.

Marriott:1980:BRBc

[107] F. H. C. Marriott. Book review: Multivariate Analysis V, by P. R. Krishnaiah. Journal of the Royal Statistical Society. Series A (General), 143(4):524, ???? 1980. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2982080.

Knott:1980:BRBb

[108] M. Knott. Book review: Lattice Path Counting and Applications, by Sri Gopal Mohanty. Journal of the Royal Statistical Society. Series A (General), 143 (4):524–525, ???? 1980. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982081.

Stern:1980:BRB

[109] G. J. A. Stern. Book review: Multidimensional Spatial Data and Decision Analysis, by Peter Nijkamp. Journal of the Royal Statistical Society. Series A (General), 143(4):525–526, ???? 1980. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982082.

Hillmann:1980:BRB

[110] H. C. Hillmann. Book review: Capital Formation in India, by Bina Roy. Journal of the Royal Statistical Society. Series A (General), 143(4):526–527, ???? 1980. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982083.

Glazebrook:1980:BRB

[111] K. D. Glazebrook. Book review: Random Processes: A First Look, by R. Syski. Journal of the Royal Statistical Society. Series A (General), 143(4):527, ???? 1980. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2982084.

Grebenik:1980:NHC

[112] E. Grebenik. Norman Henry Carrier, 1913–1979. Journal of the Royal Statistical Society. Series A (General), 143 (4):528–529, ???? 1980. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982085.

Snell:1980:DGB

[113] E. Joyce Snell. Donald Gordon Beech, 1914–1979. Journal of the Royal Statistical Society. Series A (General), 143(4): 529, ???? 1980. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2982086.

Anonymous:1980:RPc

[114] Anonymous. Recent periodicals. Journal of the Royal Statistical Society. Series A (General), 143(4):530–548, ????

1980. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982087.

Anonymous:1980:BMd

[115] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 143(4):??, ???? 1980. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2982088.

Anonymous:1981:FMa

[116] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 144(1):??, ???? 1981. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2982157.

Meisner:1981:CMB

[117] J. Meisner and F. Demirmen. The creaming method: A Bayesian procedure to forecast future oil and gas discoveries in mature exploration provinces. Journal of the Royal Statistical Society. Series A (General), 144(1): 1–31, ???? 1981. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2982158.

Moore:1981:WCR

[118] Peter G. Moore. The Wilson Committee review of the functioning of financial institutions — some statistical aspects. Journal of the Royal Statistical Society. Series A (General), 144(1):32– 65, ???? 1981. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2982159.

Diamond:1981:NQ

[119] Marion Diamond and Mervyn Stone.

Nightingale on Quetelet. Journal of the Royal Statistical Society. Series A (General), 144(1):66–79, ???? 1981. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2982160.

Straf:1981:RAR

[120] Miron L. Straf. Revenue allocation by regression: National Health Service appropriations for teaching hospitals. Journal of the Royal Statistical Society. Series A (General), 144(1):80–84, ???? 1981. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2982161.

Weerahandi:1981:MBS

[121] S. Weerahandi and J. V. Zidek. Multi-Bayesian statistical decision theory. Journal of the Royal Statistical Society. Series A (General), 144(1):85–93, ???? 1981. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2982162.

Leach:1981:REL

[122] Donald Leach. Re-evaluation of the logistic curve for human populations. Journal of the Royal Statistical Society. Series A (General), 144(1):94–103, ???? 1981. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2982163.

Churnside:1981:SAP

[123] R. J. Churnside and S. W. Creigh. Strike activity and plant size: A note. Journal of the Royal Statistical Society. Series A (General), 144(1):104–111, ???? 1981. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2982164.

Bartholomew:1981:TPI

[124] D. J. Bartholomew. The tax and price index. Journal of the Royal Statistical Society. Series A (General), 144(1):112, ???? 1981. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2982165.

Chatfield:1981:BRB

[125] C. Chatfield. Book review: Analysing Time Series, by O. D. Anderson. Journal of the Royal Statistical Society. Series A (General), 144(1):113, ???? 1981. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2982166.

Clarke:1981:BRBa

[126] G. M. Clarke. Book review: Design and Analysis of Experiments, by M. N. Das; N. C. Giri. Journal of the Royal Statistical Society. Series A (General), 144 (1):113–114, ???? 1981. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982167.

Hill:1981:BRB

[127] I. D. Hill. Book review: A Comparative Review of Statistical Software, by Ivor Francis. Journal of the Royal Statistical Society. Series A (General), 144(1): 114, ???? 1981. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2982168.

Silverman:1981:BRB

[128] B. W. Silverman. Book review: Reversibility and Stochastic Networks, by F. P. Kelly. Journal of the Royal Statistical Society. Series A (General), 144 (1):114–115, ???? 1981. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982169.

Farlie:1981:BRB

[129] D. J. G. Farlie. Book review: Statistical Computing, by William J. Kennedy; James E. Gentle. Journal of the Royal Statistical Society. Series A (General), 144(1):115–116, ???? 1981. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2982170.

Ahmad:1981:BRB

[130] R. Ahmad. Book review: Asymptotic Statistics, by P. Mandl; M. Huskova. Journal of the Royal Statistical Society. Series A (General), 144(1):116, ???? 1981. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2982171.

Wetherill:1981:BRB

[131] G. B. Wetherill. Book review: Tables for Normal Tolerance Limits, Sampling Plans and Screening, by R. E. Odeh; D. B. Owen. Journal of the Royal Statistical Society. Series A (General), 144 (1):116–117, ???? 1981. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2982172.

Moore:1981:BRB

[132] P. G. Moore. Book review: Model Building for Decision Analysis, by B. H. P. Rivett. Journal of the Royal Statistical Society. Series A (General), 144(1):117, ???? 1981. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2982173.

Anonymous:1981:RPa

[133] Anonymous. Recent periodicals. Journal of the Royal Statistical Society. Series A (General), 144(1):118–143, ????

1981. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982174.

Anonymous:1981:BMa

[134] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 144(1):i–iv, ???? 1981. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2982175.

Anonymous:1981:FMb

[135] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 144(2):??, ???? 1981. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981917.

Titterington:1981:CDT

[136] D. M. Titterington, G. D. Murray, L. S.

Murray, D. J. Spiegelhalter, A. M. Skene, J. D. F. Habbema, and G. J. Gelpke. Comparison of discrimination techniques applied to a complex data set of head injured patients. Journal of the Royal Statistical Society. Series A (General), 144(2):145–175, ???? 1981. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981918.

Diamond:1981:NQIa

[137] Marion Diamond and Mervyn Stone.

Nightingale on Quetelet II: The marginalia. Journal of the Royal Statistical Society. Series A (General), 144 (2):176–213, ???? 1981. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981919.

Bailey:1981:UAD

[138] R. A. Bailey. A unified approach to design of experiments. Journal of the Royal Statistical Society. Series A (General), 144(2):214–223, ???? 1981. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2981920.

Forsyth:1981:TPC

[139] F. G. Forsyth and R. F. Fowler. The theory and practice of chain price index numbers. Journal of the Royal Statistical Society. Series A (General), 144 (2):224–246, ???? 1981. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2981921.

Upton:1981:LDM

[140] G. J. G. Upton and B. Sarlvik. A loyaltydistance model for voting change. Journal of the Royal Statistical Society. Series A (General), 144(2):247–259, ???? 1981. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2981922.

Holt:1981:BRB

[141] D. Holt. Book review: Discrete Statistical Models with Social Science Applications, by Erling B. Andersen. Journal of the Royal Statistical Society. Series A (General), 144(2):260, ???? 1981. CODEN JSSAEF. ISSN 0035-9238.

                                        URL http://www.jstor.org/stable/
                                        2981923.

Ray:1981:BRBa

[142] W. D. Ray. Book review: Time Series, by O. D. Anderson. Journal of the Royal Statistical Society. Series A (General), 144(2):260–261, ???? 1981. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2981924.

Diggle:1981:BRBa

[143] P. J. Diggle. Book review: Exploratory and Explanatory Statistical Analysis of Spatial Data, by C. A. P. Bartels; R. H. Ketellapper. Journal of the Royal Statistical Society. Series A (General), 144(2):261, ???? 1981. CODEN JS-SAEF. ISSN 0035-9238. URL http:/

/www.jstor.org/stable/2981925.

Oliver:1981:BRB

[144] F. R. Oliver. Book review: Guide to Official Statistics. No. 3, by Central Statistical Office. Journal of the Royal Statistical Society. Series A (General), 144(2): 262, ???? 1981. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981926.

Richardson:1981:BRB

[145] S. C. Richardson. Book review: Spatial and Temporal Analysis in Ecology, by R. M. Cormack; J. K. Ord. Journal of the Royal Statistical Society. Series A (General), 144(2):262–263, ???? 1981. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981927.

Elton:1981:BRB

[146] R. A. Elton. Book review: BMDP-79: Biomedical Computer Programs, P-Series, by W. J. Dixon; M. B. Brown. Journal of the Royal Statistical Society. Series A (General), 144(2):263–264, ???? 1981. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981928.

Walters:1981:BRB

[147] D. E. Walters. Book review: Randomization Tests, by E. S. Edgington.

Journal of the Royal Statistical Society. Series A (General), 144(2):264, ???? 1981. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981929.

Haigh:1981:BRBa

[148] John Haigh. Book review: Mathematical Population Genetics, by W. J. Ewens. Journal of the Royal Statistical Society. Series A (General), 144(2):264–265, ???? 1981. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981930.

Cooke:1981:BRB

[149] D. Cooke. Book review: Mathematical Modelling in Biology and Ecology, by W. M. Getz. Journal of the Royal Statistical Society. Series A (General), 144 (2):265–266, ???? 1981. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981931.

Clarke:1981:BRBb

[150] G. M. Clarke. Book review: Analysis of Qualitative Data: Vol. 2, New Developments, by S. J. Haberman. Journal of the Royal Statistical Society. Series A (General), 144(2):266, ???? 1981. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981932.

Finney:1981:BRB

[151] D. J. Finney. Book review: Bioassay, by J. J. Hubert. Journal of the Royal Statistical Society. Series A (General), 144(2):267, ???? 1981. CODEN JS-SAEF. ISSN 0035-9238. URL http:/

/www.jstor.org/stable/2981933.

Edwards:1981:BRB

[152] A. W. F. Edwards. Book review: Science, Belief and Behaviour: Essays in Honour of R. B. Braithwaite, by D. H. Mellor. Journal of the Royal Statistical Society. Series A (General), 144(2):267, ???? 1981. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981934.

Clarke:1981:BRBc

[153] G. M. Clarke. Book review: The Linear Hypothesis; A General Theory, 2nd Edition, by G. A. F. Seber. Journal of the Royal Statistical Society. Series A (General), 144(2):268, ???? 1981. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2981935.

Gittins:1981:BRB

[154] J. C. Gittins. Book review: Finite Generalised Markov Programming, by P. J. Weeda. Journal of the Royal Statistical Society. Series A (General), 144(2):268– 269, ???? 1981. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981936.

Johnson:1981:P

[155] N. L. Johnson. E. S. Pearson, 1895– 1980. Journal of the Royal Statistical Society. Series A (General), 144(2):270– 271, ???? 1981. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981937.

Newman:1981:FS

[156] G. B. Newman. Frank Sandon, 1891– 1979. Journal of the Royal Statistical Society. Series A (General), 144(2):271, ???? 1981. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981938.

Anonymous:1981:RPb

[157] Anonymous. Recent periodicals. Journal of the Royal Statistical Society. Series A (General), 144(2):272–288, ???? 1981. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2981939.

Anonymous:1981:BMb

[158] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 144(2):i–vi, ???? 1981. CODEN JSSAEF. ISSN 0035-9238.

                                        URL http://www.jstor.org/stable/
                                        2981940.

Anonymous:1981:FMc

[159] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 144(3):??, ???? 1981. CODEN JSSAEF. ISSN 0035-9238.

                                        URL http://www.jstor.org/stable/
                                        2981795.

Cox:1981:TGP

[160] D. R. Cox. Theory and general principle in statistics. Journal of the Royal Statistical Society. Series A (General), 144 (3):289–297, ???? 1981. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2981796.

Darby:1981:LLI

[161] S. C. Darby and J. A. Reissland. Low levels of ionizing radiation and cancer — are we underestimating the risk? Journal of the Royal Statistical Society. Series A (General), 144(3):298–331, ???? 1981. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2981797.

Diamond:1981:NQIb

[162] Marion Diamond and Mervyn Stone.

Nightingale on Quetelet III: Essay in memoriam. Journal of the Royal Statistical Society. Series A (General), 144 (3):332–351, ???? 1981. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981798.

Lythe:1981:SGD

[163] Charlotte Lythe and Madhavi Majmudar. Scottish Gross Domestic Product statistics for 1961–71. Journal of the Royal Statistical Society. Series A (General), 144(3):352–359, ???? 1981. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981799.

Currie:1981:FPP

[164] Iain D. Currie. Further percentage points of Greenwood’s statistic. Journal of the Royal Statistical Society. Series A (General), 144(3):360–363, ???? 1981. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981800.

Stephens:1981:FPP

[165] M. A. Stephens. Further percentage points for Greenwood’s statistic. Journal of the Royal Statistical Society. Series A (General), 144(3):364–366, ???? 1981. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981801.

Muir:1981:BRBa

[166] W. W. Muir. Book review: Regression Diagnostics: Identifying Influential Data and Sources of Collinearity, by D. A. Belsley; E. Kuh; R. E. Welsch. Journal of the Royal Statistical Society. Series A (General), 144(3):367–368, ???? 1981. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981802.

Carroll:1981:BRB

[167] P. S. Carroll. Book review: Techniques for Making Functional Population Projections, by D. J. Bogue; S. Bischler. Journal of the Royal Statistical Society. Series A (General), 144(3):368, ???? 1981. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981803.

Diggle:1981:BRBb

[168] P. J. Diggle. Book review: Estimation of Density from Line Transect Sampling of Biological Populations, by K. P. Burnham; D. R. Anderson; J. L. Laake. Journal of the Royal Statistical Society. Series A (General), 144(3):369, ???? 1981. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981804.

Peers:1981:BRB

[169] H. W. Peers. Book review: Asymptotic Theory of Statistical Tests and Estimation: In Honor of Wassily Hoeffding, by I. M. Chakravarti. Journal of the Royal Statistical Society. Series A (General), 144(3):369–370, ???? 1981. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981805.

Tong:1981:BRB

[170] H. Tong. Book review: The Analysis of Time Series: An Introduction, by C. Chatfield. Journal of the Royal Statistical Society. Series A (General), 144(3): 370, ???? 1981. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981806.

Haigh:1981:BRBb

[171] John Haigh. Book review: Practical Nonparametric Statistics, by W. J. Conover. Journal of the Royal Statistical Society. Series A (General), 144 (3):370–371, ???? 1981. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981807.

Stone:1981:BRB

[172] M. Stone. Book review: Inference and Linear Models, by D. A. S. Fraser. Journal of the Royal Statistical Society. Series A (General), 144(3):371, ???? 1981. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981808.

Thomas:1981:BRB

[173] D. A. H. Thomas. Book review: A First Course in Statistics, by A. Greer. Journal of the Royal Statistical Society. Series A (General), 144(3):372, ???? 1981. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981809.

Critchley:1981:BRB

[174] F. Critchley. Book review: Multivariate Statistical Analysis: Proceedings of the Research Seminar at Dalhousie University, Halifax, October 5th–7th, 1979, by R. P. Gupta. Journal of the Royal Statistical Society. Series A (General), 144 (3):372–373, ???? 1981. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981810.

Muir:1981:BRBb

[175] W. W. Muir. Book review: Testing the Constancy of Regression Models Over Time, by P. Hackl. Journal of the Royal Statistical Society. Series A (General), 144(3):373–374, ???? 1981. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2981811.

Chedzoy:1981:BRBa

[176] O. B. Chedzoy. Book review: Introductory Statistical Analysis, by D. L. Harnett; J. L. Murphy. Journal of the Royal Statistical Society. Series A (General), 144(3):374, ???? 1981. CODEN JS-SAEF. ISSN 0035-9238. URL http:/

                                        /www.jstor.org/stable/2981812.

Chedzoy:1981:BRBb

[177] O. B. Chedzoy. Book review: Introduction to Statistics, by M. R. Heyworth; J. R. Sealy. Journal of the Royal Statistical Society. Series A (General), 144 (3):374–375, ???? 1981. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2981813.

Priestley:1981:BRB

[178] M. B. Priestley. Book review: The Advanced Theory of Statistics, Vol. 2: Inference and Decision, by M. Kendal; A. Stuart. Journal of the Royal Statistical Society. Series A (General), 144(3):375, ???? 1981. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2981814.

Wilkinson:1981:BRBa

[179] R. K. Wilkinson. Book review: Variety, Equity and Efficiency, by K. Lancaster. Journal of the Royal Statistical Society. Series A (General), 144(3):375– 376, ???? 1981. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2981815.

Wright:1981:BRB

[180] C. K. Wright. Book review: Probabilistic Models in Engineering Sciences, Vol. II: Random Noise, Signals and Dynamic Systems, by H. J. Larson; B. O. Shubert. Journal of the Royal Statistical Society. Series A (General), 144(3):376, ???? 1981. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981816.

Ehrenberg:1981:BRB

[181] A. S. C. Ehrenberg. Book review: Foundations of Statistics: A Survey for Managers, by P. Montagnon. Journal of the Royal Statistical Society. Series A (General), 144(3):377, ???? 1981. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2981817.

Bibby:1981:BRBa

[182] J. Bibby. Book review: Probability: Inequalities in Multivariate Distributions, by Y. L. Tong. Journal of the Royal Statistical Society. Series A (General), 144(3):377, ???? 1981. CODEN JS-SAEF. ISSN 0035-9238. URL http:/

/www.jstor.org/stable/2981818.

Anonymous:1981:RPc

[183] Anonymous. Recent periodicals. Journal of the Royal Statistical Society. Series A (General), 144(3):378–387, ???? 1981. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981819.

Moran:1981:CRD

[184] P. A. P. Moran. Corrigenda: The random division of an interval. Journal of the Royal Statistical Society. Series A (General), 144(3):388, ???? 1981.

CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981820.

Hill:1981:CAD

[185] I. D. Hill. Corrigenda: Approximating the distribution of Greenwood’s statistic with Johnson distributions. Journal of the Royal Statistical Society. Series A (General), 144(3):388, ???? 1981. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981821.

Anonymous:1981:BMc

[186] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 144(3):??, ???? 1981. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981822.

Anonymous:1981:VI

[187] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (General), 144(4):557–560, ???? 1981. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981823.

Anonymous:1981:FMd

[188] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 144(4):??, ???? 1981. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981824.

Bissell:1981:UAQ

[189] A. F. Bissell and W. A. Pridmore. The UK average quantity system and its statistical implications. Journal of the Royal Statistical Society. Series A (General), 144(4):389–418, ???? 1981.

CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981825.

Aitkin:1981:SMD

[190] Murray Aitkin, Dorothy Anderson, and John Hinde. Statistical modelling of data on teaching styles. Journal of the Royal Statistical Society. Series A (General), 144(4):419–461, ???? 1981. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981826.

Mead:1981:RSI

[191] R. Mead and Janet Riley. A review of statistical ideas relevant to intercropping research. Journal of the Royal Statistical Society. Series A (General), 144 (4):462–509, ???? 1981. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981827.

Anonymous:1981:POH

[192] Anonymous. Proceedings of the one hundred and forty-seventh Annual General Meeting of the Royal Statistical Society. Journal of the Royal Statistical Society. Series A (General), 144(4):510, ???? 1981. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981828.

Anonymous:1981:PSG

[193] Anonymous. Proceedings of a Special General Meeting of the Royal Statistical Society. Journal of the Royal Statistical Society. Series A (General), 144 (4):510–511, ???? 1981. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981829.

Anonymous:1981:HTR

[194] Anonymous. Honorary Treasurer’s report. Journal of the Royal Statistical Society. Series A (General), 144(4):512, ???? 1981. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2981830.

Anonymous:1981:ARC

[195] Anonymous. Annual report of the Council. Journal of the Royal Statistical Society. Series A (General), 144(4):513–536, ???? 1981. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2981831.

Eplett:1981:BRB

[196] W. J. R. Eplett. Book review: Nonparametric Sequential Selection Procedures, by H. Buringer. Journal of the Royal Statistical Society. Series A (General), 144(4):537, ???? 1981. CODEN JS-SAEF. ISSN 0035-9238. URL http:/

                                        /www.jstor.org/stable/2981832.

Baczkowski:1981:BRB

[197] A. J. Baczkowski. Book review: Practical Geostatistics, by Isobel Clark. Journal of the Royal Statistical Society. Series A (General), 144(4):537, ???? 1981. CODEN JSSAEF. ISSN 0035-9238.

                                        URL http://www.jstor.org/stable/
                                        2981833.

Wilkinson:1981:BRBb

[198] R. K. Wilkinson. Book review: Economics and Consumer Behaviour, by A. Deaton; J. Muellbauer. Journal of the Royal Statistical Society. Series A (General), 144(4):538, ???? 1981. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2981834.

Bibby:1981:BRBb

[199] John Bibby. Book review: The Readable Maths and Statistics Book, by B. Edwards. Journal of the Royal Statistical Society. Series A (General), 144 (4):538–539, ???? 1981. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981835.

Ray:1981:BRBb

[200] W. D. Ray. Book review: The Econometric Analysis of Time Series, by A. C. Harvey. Journal of the Royal Statistical Society. Series A (General), 144(4): 539, ???? 1981. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981836.

Wilkinson:1981:BRBc

[201] R. K. Wilkinson. Book review: Exercises and Worked Examples in Statistics, by G. E. Skipworth. Journal of the Royal Statistical Society. Series A (General), 144(4):539–540, ???? 1981. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981837.

Thunhurst:1981:BRB

[202] Colin Thunhurst. Book review: Statistical Methods for Planners, by T. R. Willemain. Journal of the Royal Statistical Society. Series A (General), 144(4): 540, ???? 1981. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981838.

Jones:1981:BRB

[203] G. C. Jones. Book review: Seasonal Analysis of Economic Time Series, by A. Zellner. Journal of the Royal Statistical Society. Series A (General), 144(4): 541, ???? 1981. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2981839.

Anonymous:1981:RPd

[204] Anonymous. Recent periodicals. Journal of the Royal Statistical Society. Series A (General), 144(4):542–556, ???? 1981. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2981840.

Anonymous:1981:BMd

[205] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 144(4):??, ???? 1981. CODEN JSSAEF. ISSN 0035-9238.

                                        URL http://www.jstor.org/stable/
                                        2981841.

Anonymous:1982:FMa

[206] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 145(1):??, ???? 1982. CODEN JSSAEF. ISSN 0035-9238.

                                        URL http://www.jstor.org/stable/
                                        2981419.

Kenny:1982:LTE

[207] P. B. Kenny and J. Durbin. Local trend estimation and seasonal adjustment of economic and social time series. Journal of the Royal Statistical Society. Series A (General), 145(1):1–41, ???? 1982. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2981420.

Kalton:1982:EQS

[208] Graham Kalton and Howard Schuman. The effect of the question on survey responses: A review. Journal of the Royal Statistical Society. Series A (General), 145(1):42–73, ???? 1982. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981421.

Wallis:1982:SAR

[209] Kenneth F. Wallis. Seasonal adjustment and revision of current data: Linear filters for the x-11 method. Journal of the Royal Statistical Society. Series A (General), 145(1):74–85, ???? 1982. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2981422.

Upton:1982:CAT

[210] Graham J. G. Upton. A comparison of alternative tests for the 2 × 2 comparative trial. Journal of the Royal Statistical Society. Series A (General), 145 (1):86–105, ???? 1982. CODEN JS-SAEF. ISSN 0035-9238. URL http:/

/www.jstor.org/stable/2981423.

Nour:1982:ETN

[211] El-Sayed Nour. On the estimation of the total number of vital events with data from dual collection systems. Journal of the Royal Statistical Society. Series A (General), 145(1):106–116, ???? 1982. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981424.

Lindley:1982:IPJ

[212] D. V. Lindley. The improvement of probability judgements. Journal of the Royal Statistical Society. Series A (General), 145(1):117–126, ???? 1982. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981425.

Culyer:1982:RAR

[213] A. J. Culyer, J. Wiseman, M. F. Drummond, and P. A. West. Revenue allocation by regression: A rejoinder. Journal of the Royal Statistical Society. Series A (General), 145(1):127–133, ???? 1982. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981426.

Ray:1982:BRBa

[214] W. D. Ray. Book review: Time Series: Data Analysis and Theory, by David Brillinger. Journal of the Royal Statistical Society. Series A (General), 145 (1):134–135, ???? 1982. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981427.

Haigh:1982:BRB

[215] John Haigh. Book review: How to Pass Examinations in Statistics, by D. V. Friend. Journal of the Royal Statistical Society. Series A (General), 145 (1):135–136, ???? 1982. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981428.

Barnett:1982:BRB

[216] Vic Barnett. Book review: Regression Analysis and Its Application: A Data-Oriented Approach, by Richard F. Gunst; Robert L. Mason. Journal of the Royal Statistical Society. Series A (General), 145(1):136–137, ???? 1982. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981429.

Hill:1982:BRB

[217] I. D. Hill. Book review: Godel, Escher, Bach: An Eternal Golden Braid, by D. R. Hofstadter. Journal of the Royal Statistical Society. Series A (General), 145 (1):137–138, ???? 1982. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981430.

Hall:1982:BRB

[218] David J. Hall. Book review: Measurement of Levels of Health, by W. Holland. Journal of the Royal Statistical Society. Series A (General), 145(1):138– 139, ???? 1982. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981431.

Preece:1982:BRB

[219] D. A. Preece. Book review: Incomplete Block Designs. Lecture Notes in Statistics, by P. W. M. John. Journal of the Royal Statistical Society. Series A (General), 145(1):139, ???? 1982. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2981432.

Gray:1982:BRB

[220] J. R. Gray. Book review: Computational Probability, by P. M. Kahn. Journal of the Royal Statistical Society. Series A (General), 145(1):139–140, ???? 1982. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981433.

Marriott:1982:BRB

[221] F. H. C. Marriott. Book review: Multivariate Analysis, by Maurice Kendall. Journal of the Royal Statistical Society. Series A (General), 145(1):140, ???? 1982. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981434.

Hiorns:1982:BRB

[222] R. W. Hiorns. Book review: Time Lags in Biological Models, by Norman Mac-Donald. Journal of the Royal Statistical Society. Series A (General), 145 (1):140–141, ???? 1982. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981435.

Titterington:1982:BRB

[223] D. M. Titterington. Book review: Recent Developments in Statistical Inference and Data Analysis, by K. Matusita. Journal of the Royal Statistical Society. Series A (General), 145(1):141, ???? 1982. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981436.

Cooke:1982:BRB

[224] D. Cooke. Book review: Biostatistics Casebook, by R. G. Miller,; B. Efron; B. W. Brown,; L. E. Moses. Journal of the Royal Statistical Society. Series A (General), 145(1):141–142, ???? 1982. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981437.

Chedzoy:1982:BRBa

[225] O. B. Chedzoy. Book review: Statistics — A Dictionary of Terms and Ideas, by Colm O’Muircheartaigh; David Pitt Francis. Journal of the Royal Statistical Society. Series A (General), 145(1): 142, ???? 1982. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981438.

Cormack:1982:BRB

[226] R. M. Cormack. Book review: Multivariate Methods in Ecological Work, by L. Orloci; C. R. Rao; W. M. Stiteler. Journal of the Royal Statistical Society. Series A (General), 145(1):142–143, ???? 1982. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981439.

Ray:1982:BRBb

[227] W. D. Ray. Book review: Stochastic Storage Processes, by N. V. Prabhu. Journal of the Royal Statistical Society. Series A (General), 145(1):143–144, ???? 1982. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981440.

Knott:1982:BRB

[228] M. Knott. Book review: Matrix Derivatives, by G. S. Rogers. Journal of the Royal Statistical Society. Series A (General), 145(1):144, ???? 1982. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2981441.

Freeman:1982:BRB

[229] P. Freeman. Book review: Clinical Trials, by Daniel Schwartz; Robert Flamant; Joseph Lallouch; M. J. R. Healy. Journal of the Royal Statistical Society. Series A (General), 145(1):144–145, ???? 1982. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981442.

Sibson:1982:BRB

[230] Robin Sibson. Book review: Optimal Design: An Introduction to the Theory for Parameter Estimation, by S. D. Silvey. Journal of the Royal Statistical Society. Series A (General), 145(1):145, ???? 1982. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981443.

Scott:1982:BRB

[231] W. F. Scott. Book review: Build Study 1979, by Society of Actuaries and the Association of Life Insurance Medical Directors of America. Journal of the Royal Statistical Society. Series A (General), 145(1):145–146, ???? 1982. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981444.

Fisk:1982:BRB

[232] P. R. Fisk. Book review: Applied Linear Regression, by S. Weisberg. Journal of the Royal Statistical Society. Series A (General), 145(1):146–147, ???? 1982. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981445.

Box:1982:HOH

[233] G. E. P. Box. Herman Otto Hartley, 1912–1980. Journal of the Royal Statistical Society. Series A (General), 145(1): 148, ???? 1982. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981446.

Anonymous:1982:RPa

[234] Anonymous. Recent periodicals. Journal of the Royal Statistical Society. Series A (General), 145(1):149–168, ???? 1982. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981447.

Anonymous:1982:BMa

[235] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 145(1):??, ???? 1982. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981448.

Anonymous:1982:FMb

[236] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 145(2):??, ???? 1982. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981532.

vanderPloeg:1982:RAS

[237] F. van der Ploeg. Reliability and the adjustment of sequences of large economic accounting matrices. Journal of the Royal Statistical Society. Series A (General), 145(2):169–194, ???? 1982. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981533.

Hoinville:1982:RRG

[238] G. Hoinville and T. M. F. Smith. The Rayner review of government statistical services. Journal of the Royal Statistical Society. Series A (General), 145 (2):195–207, ???? 1982. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981534.

Leech:1982:GRE

[239] Dennis Leech and Keith Cowling. Generalized regression estimation from grouped observations: A generalization and an application to the relationship between diet and mortality. Journal of the Royal Statistical Society. Series A (General), 145(2):208–223, ???? 1982. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981535.

Jolliffe:1982:UCC

[240] I. T. Jolliffe, B. Jones, and B. J. T. Morgan. Utilising clusters: A case-study involving the elderly. Journal of the Royal Statistical Society. Series A (General), 145(2):224–236, ???? 1982. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981536.

Gokhale:1982:APD

[241] D. V. Gokhale and S. James Press. Assessment of a prior distribution for the correlation coefficient in a bivariate normal distribution. Journal of the Royal Statistical Society. Series A (General), 145(2):237–249, ???? 1982. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2981537.

Stigler:1982:TBB

[242] Stephen M. Stigler. Thomas Bayes’s Bayesian inference. Journal of the Royal Statistical Society. Series A (General), 145(2):250–258, ???? 1982. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2981538.

Finney:1982:TFY

[243] D. J. Finney. A tribute to Frank yates. Journal of the Royal Statistical Society. Series A (General), 145(2):259–260, ???? 1982. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2981539.

Daniels:1982:THC

[244] H. E. Daniels. A tribute to L. H. C. Tippett. Journal of the Royal Statistical Society. Series A (General), 145 (2):261–262, ???? 1982. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2981540.

Reddaway:1982:BRB

[245] W. B. Reddaway. Book review: Industrialisation and the Basis for Trade, by R. A. Batchelor; R. L. Major; A. D. Morgan. Journal of the Royal Statistical Society. Series A (General), 145 (2):263–264, ???? 1982. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2981541.

Oakes:1982:BRB

[246] David Oakes. Book review: Statistical Methods in Cancer Research, Vol 1: The Analysis of Case-Control Studies, by N. E. Breslow; N. E. Day. Journal of the Royal Statistical Society. Series A (General), 145(2):264–265, ???? 1982. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981542.

Hillmann:1982:BRB

[247] H. C. Hillmann. Book review: Growth, Innovation and Employment, by J. C. Cox; H. Kriegbaum. Journal of the Royal Statistical Society. Series A (General), 145(2):265, ???? 1982. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2981543.

Graham:1982:BRB

[248] Hilary Graham. Book review: Women and Government Statistics, by Equal Opportunities Commission. Journal of the Royal Statistical Society. Series A (General), 145(2):265–266, ???? 1982. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981544.

Brown:1982:BRB

[249] D. Brown. Book review: Statistics for Biologists, by D. J. Finney. Journal of the Royal Statistical Society. Series A (General), 145(2):266–267, ???? 1982. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981545.

Clarke:1982:BRBa

[250] G. M. Clarke. Book review: Handbook of Statistics, Volume 1: Analysis of Variance, by P. R. Krishnaiah. Journal of the Royal Statistical Society. Series A (General), 145(2):267–268, ???? 1982. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981546.

Loynes:1982:BRB

[251] R. M. Loynes. Book review: Developments in Statistics, Volume 3, by P. R. Krishnaiah. Journal of the Royal Statistical Society. Series A (General), 145 (2):268–269, ???? 1982. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981547.

Ray:1982:BRBc

[252] W. D. Ray. Book review: Spectral Analysis and Time Series. Vol. 1: Univariate Series, by M. B. Priestley. Journal of the Royal Statistical Society. Series A (General), 145(2):269, ???? 1982. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2981548.

Ray:1982:BRBd

[253] W. D. Ray. Book review: Volume 2: Multivariate Series, Prediction and Control, by M. B. Priestley. Journal of the Royal Statistical Society. Series A (General), 145(2):269–270, ???? 1982. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981549.

Anonymous:1982:RPb

[254] Anonymous. Recent periodicals. Journal of the Royal Statistical Society. Series A (General), 145(2):271–283, ???? 1982. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981550.

Anonymous:1982:BMb

[255] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 145(2):??, ???? 1982. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981551.

Anonymous:1982:FMc

[256] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 145(3):??, ???? 1982. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981864.

Ramsay:1982:SSA

[257] J. O. Ramsay. Some statistical approaches to multidimensional scaling data. Journal of the Royal Statistical Society. Series A (General), 145(3):285– 312, ???? 1982. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981865.

Pocock:1982:AGV

[258] S. J. Pocock, D. G. Cook, and A. G.

Shaper. Analysing geographic variation in cardiovascular mortality: Methods and results. Journal of the Royal Statistical Society. Series A (General), 145 (3):313–341, ???? 1982. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981866.

Taffler:1982:FCF

[259] R. J. Taffler. Forecasting company failure in the UK using discriminant analysis and financial ratio data. Journal of the Royal Statistical Society. Series A (General), 145(3):342–358, ???? 1982. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981867.

Oliver:1982:NLC

[260] F. R. Oliver. Notes on the logistic curve for human populations. Journal of the Royal Statistical Society. Series A (General), 145(3):359–363, ???? 1982. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981868.

Ehrenberg:1982:HGB

[261] A. S. C. Ehrenberg. How good is best? Journal of the Royal Statistical Society. Series A (General), 145(3):364–366, ???? 1982. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981869.

Barnard:1982:BRB

[262] G. A. Barnard. Book review: The Writings of Leonard Jimmie Savage — A Memorial Selection, by Leonard Jimmie Savage; American Statistical Association; The Institute of Mathematical Statistics. Journal of the Royal Statistical Society. Series A (General), 145(3): 367, ???? 1982. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981870.

Curnow:1982:BRB

[263] R. N. Curnow. Book review: Operational Research Applied to the Health Services, by D. Boldy. Journal of the Royal Statistical Society. Series A (General), 145(3):367–368, ???? 1982. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981871.

Yates:1982:BRB

[264] F. Yates. Book review: R. A. Fisher:

An Appreciation, by R. A. Fisher; S. E. Fienberg; D. V. Hinckley. Journal of the Royal Statistical Society. Series A (General), 145(3):368–370, ???? 1982. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981872.

Ray:1982:BRBe

[265] W. D. Ray. Book review: Time Series Models, by A. C. Harvey. Journal of the Royal Statistical Society. Series A (General), 145(3):370, ???? 1982. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2981873.

Sutherland:1982:BRB

[266] Gillian Sutherland. Book review: Statistics in Britain 1865–1930. The Social Construction of Scientific Knowledge, by Donald A. MacKenzie. Journal of the Royal Statistical Society. Series A (General), 145(3):370–371, ???? 1982. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981874.

Eplett:1982:BRB

[267] W. J. R. Eplett. Book review: Approximation Theorems of Mathematical Statistics, by R. J. Serfling. Journal of the Royal Statistical Society. Series A (General), 145(3):371–372, ???? 1982. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981875.

Trustrum:1982:BRB

[268] K. Trustrum. Book review: Quick Statistics: An Introduction to Non-Parametric Methods, by Peter Sprent.

Journal of the Royal Statistical Society. Series A (General), 145(3):372, ???? 1982. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2981876.

Ray:1982:BRBf

[269] G. F. Ray. Book review: Business Cycle Surveys, by Werner H. Strigel. Journal of the Royal Statistical Society. Series A (General), 145(3):372–373, ???? 1982. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2981877.

Paine:1982:SPC

[270] Toby Paine. Sir Paul Chambers, 1904– 1981. Journal of the Royal Statistical Society. Series A (General), 145(3):374– 375, ???? 1982. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2981878.

Anonymous:1982:RPc

[271] Anonymous. Recent periodicals. Journal of the Royal Statistical Society. Series A (General), 145(3):376–394, ???? 1982. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2981879.

Anonymous:1982:BMc

[272] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 145(3):??, ???? 1982. CODEN JSSAEF. ISSN 0035-9238.

                                        URL http://www.jstor.org/stable/
                                        2981880.

Anonymous:1982:VI

[273] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (General), 145(4):550–553, ????

1982. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982089.

Anonymous:1982:FMd

[274] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 145(4):??, ???? 1982. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2982090.

Downton:1982:LPS

[275] F. Downton. Legal probability and statistics. Journal of the Royal Statistical Society. Series A (General), 145 (4):395–402, ???? 1982. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982091.

Newell:1982:RSE

[276] David Newell. The role of the statistician as an expert witness. Journal of the Royal Statistical Society. Series A (General), 145(4):403–409, ???? 1982. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2982092.

Anonymous:1982:A

[277] Anonymous. Addendum. Journal of the Royal Statistical Society. Series A (General), 145(4):409, ???? 1982. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2982093.

Fienberg:1982:SAE

[278] Stephen E. Fienberg and Miron L.

Straf. Statistical assessments as evidence. Journal of the Royal Statistical Society. Series A (General), 145(4):410– 421, ???? 1982. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2982094.

Napley:1982:LS

[279] David Napley. Lawyers and statisticians. Journal of the Royal Statistical Society. Series A (General), 145(4):422– 438, ???? 1982. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2982095.

Burrell:1982:ALD

[280] Quentin L. Burrell and Violet R. Cane. The analysis of library data. Journal of the Royal Statistical Society. Series A (General), 145(4):439–471, ???? 1982. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2982096.

Andersen:1982:EEB

[281] Allan Andersen. An empirical examination of Box-Jenkins forecasting. Journal of the Royal Statistical Society. Series A (General), 145(4):472–478, ???? 1982. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2982097.

Pocock:1982:DAG

[282] S. J. Pocock, D. G. Cook, and A. G. Shaper. In discussion of “analysing geographic variations in cardiovascular mortality: Methods and results”. Journal of the Royal Statistical Society. Series A (General), 145(4):479–480, ???? 1982. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2982098.

Anonymous:1982:POH

[283] Anonymous. Proceedings of the one hundred and forty-eighth Annual General Meeting of the Royal Statistical Society. Journal of the Royal Statistical Society. Series A (General), 145(4):481, ???? 1982. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2982099.

Anonymous:1982:PR

[284] Anonymous. President’s report. Journal of the Royal Statistical Society. Series A (General), 145(4):482–483, ???? 1982. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982100.

Anonymous:1982:HTR

[285] Anonymous. Honorary Treasurer’s report. Journal of the Royal Statistical Society. Series A (General), 145(4):484, ???? 1982. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2982101.

Anonymous:1982:ARC

[286] Anonymous. Annual report of the Council. Journal of the Royal Statistical Society. Series A (General), 145(4):485–507, ???? 1982. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2982102.

Smith:1982:BRBa

[287] A. F. M. Smith. Book review: Expected Utility Hypotheses and the Allais Paradox, by M. Allais; O. Hagen. Journal of the Royal Statistical Society. Series A (General), 145(4):508, ???? 1982. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982103.

Ecob:1982:BRB

[288] Russell Ecob. Book review: Mathematical Methods in Social Sciences, by D. J. Bartholomew. Journal of the Royal Statistical Society. Series A (General), 145(4):509, ???? 1982. CODEN JS-SAEF. ISSN 0035-9238. URL http:/

                                        /www.jstor.org/stable/2982104.

Unwin:1982:BRB

[289] Antony R. Unwin. Book review: Introduction to Queueing Theory, by R. B. Cooper. Journal of the Royal Statistical Society. Series A (General), 145(4): 509, ???? 1982. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2982105.

Clarke:1982:BRBb

[290] G. M. Clarke. Book review: Applied Regression Analysis, by N. R. Draper; H. Smith. Journal of the Royal Statistical Society. Series A (General), 145(4):510, ???? 1982. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2982106.

Plewis:1982:BRB

[291] Ian Plewis. Book review: Statistical Methods for Rates and Proportions, by J. L. Fleiss. Journal of the Royal Statistical Society. Series A (General), 145 (4):510–511, ???? 1982. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2982107.

Ray:1982:BRBg

[292] W. D. Ray. Book review: Time Series Analysis: A Comprehensive Introduction for Social Scientists, by John M. Gottman. Journal of the Royal Statistical Society. Series A (General), 145(4): 511, ???? 1982. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2982108.

Moore:1982:BRB

[293] B. A. Moore. Book review: Discrimination and Classification, by D. J. Hand. Journal of the Royal Statistical Society. Series A (General), 145(4):511, ???? 1982. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982109.

Abakuks:1982:BRB

[294] Andris Abakuks. Book review: The Mathematical Theory of the Dynamics of Biological Populations II, by R. W. Hiorns; D. Cooke. Journal of the Royal Statistical Society. Series A (General), 145(4):512, ???? 1982. CODEN JS-SAEF. ISSN 0035-9238. URL http:/

/www.jstor.org/stable/2982110.

Ray:1982:BRBh

[295] W. D. Ray. Book review: Modern Statistics–Methods and Applications, by Robert V. Hogg. Journal of the Royal Statistical Society. Series A (General), 145(4):512, ???? 1982. CODEN JS-SAEF. ISSN 0035-9238. URL http:/

/www.jstor.org/stable/2982111.

Shotter:1982:BRB

[296] Margaret Shotter. Book review: Biostatistics: Vol. 1 — 1064 Questions, by J. J. Hubert; E. M. Carter; Biostatistics: Vol. 2 — 1064 Answers, by J. J. Hubert; E. M. Carter. Journal of the Royal Statistical Society. Series A (General), 145 (4):512–513, ???? 1982. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982112.

Aitkin:1982:BRB

[297] Murray Aitkin. Book review: Analysis of Categorical Data: Dual Scaling and its Applications, by S. Nishisato. Journal of the Royal Statistical Society. Series A (General), 145(4):513–516, ???? 1982. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982113.

Oliver:1982:BRB

[298] F. R. Oliver. Book review: Prices and Quantities, by Arthur M. Okun. Journal of the Royal Statistical Society. Series A (General), 145(4):516–517, ???? 1982. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982114.

Muir:1982:BRB

[299] W. W. Muir. Book review: Factorial Designs, by B. L. Raktoe; A. Hedayat; W. T. Federer. Journal of the Royal Statistical Society. Series A (General), 145(4):517, ???? 1982. CODEN JS-SAEF. ISSN 0035-9238. URL http:/

/www.jstor.org/stable/2982115.

Chedzoy:1982:BRBb

[300] O. B. Chedzoy. Book review: Intermediate Statistical Methods, by G. Barrie Wetherill. Journal of the Royal Statistical Society. Series A (General), 145 (4):517–518, ???? 1982. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982116.

Ray:1982:BRBi

[301] W. D. Ray. Book review: A User’s Guide to the Gottman-Williams Time Series Analysis: Computer Programs for Social Scientists, by Esther A. Williams; John M. Gottman. Journal of the Royal Statistical Society. Series A (General), 145(4):518, ???? 1982. CODEN JS-SAEF. ISSN 0035-9238. URL http:/

/www.jstor.org/stable/2982117.

Smith:1982:BRBb

[302] A. F. M. Smith. Book review: Bayesian Analysis in Econometrics and Statistics: Essays in Honour of Harold Jeffreys, by A. Zellner. Journal of the Royal Statistical Society. Series A (General), 145 (4):518–520, ???? 1982. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982118.

Anonymous:1982:CR

[303] Anonymous. Correction: Reviews. Journal of the Royal Statistical Society. Series A (General), 145(4):520, ???? 1982. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2982119.

Yates:1982:WGC

[304] F. Yates. William Gemmell Cochran, 1909–1980. Journal of the Royal Statistical Society. Series A (General), 145 (4):521–523, ???? 1982. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982120.

Cox:1982:SGM

[305] Peter R. Cox. Sir George Maddex, 1895– 1982. Journal of the Royal Statistical Society. Series A (General), 145(4):523, ???? 1982. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2982121.

Hammersley:1982:JN

[306] J. M. Hammersley. J. Neyman, 1894– 1981. Journal of the Royal Statistical Society. Series A (General), 145(4):523– 524, ???? 1982. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2982122.

Rivett:1982:PKD

[307] B. H. P. Rivett. Professor K. D. Tocher, 1921–1981. Journal of the Royal Statistical Society. Series A (General), 145 (4):525–526, ???? 1982. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2982123.

Armitage:1982:JOI

[308] Peter Armitage. Joseph Oscar Irwin, 1898–1982. Journal of the Royal Statistical Society. Series A (General), 145 (4):526–528, ???? 1982. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2982124.

Anonymous:1982:RPd

[309] Anonymous. Recent periodicals. Journal of the Royal Statistical Society. Series A (General), 145(4):529–549, ???? 1982. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2982125.

Anonymous:1982:BMd

[310] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 145(4):??, ???? 1982. CODEN JSSAEF. ISSN 0035-9238.

                                        URL http://www.jstor.org/stable/
                                        2982126.

Anonymous:1983:FMa

[311] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 146(1):??, ???? 1983. CODEN JSSAEF. ISSN 0035-9238.

                                        URL http://www.jstor.org/stable/
                                        2981485.

Carpenter:1983:SPR

[312] R. G. Carpenter. Scoring to provide riskrelated primary health care: Evaluation and up-dating during use. Journal of the Royal Statistical Society. Series A (General), 146(1):1–32, ???? 1983. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2981486.

Atkinson:1983:RID

[313] A. B. Atkinson and J. Micklewright. On the reliability of income data in the family expenditure survey 1970–1977. Journal of the Royal Statistical Society. Series A (General), 146(1):33–61, ???? 1983. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981487.

Newby:1983:DIS

[314] Martin Newby and Jonathan Winterton. The duration of industrial stoppages. Journal of the Royal Statistical Society. Series A (General), 146(1):62– 70, ???? 1983. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981488.

Woodward:1983:CAH

[315] Wayne A. Woodward and H. L. Gray.

A comparison of autoregressive and harmonic component models for the Lynx data. Journal of the Royal Statistical Society. Series A (General), 146(1):71–73, ???? 1983. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981489.

Woodward:1983:FGA

[316] M. Woodward. On forecasting grade, age and length of service distributions in manpower systems. Journal of the Royal Statistical Society. Series A (General), 146(1):74–84, ???? 1983. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2981490.

Macarthur:1983:NEV

[317] Ewan W. Macarthur. A note on the estimation of vital events: Total number and proportion. Journal of the Royal Statistical Society. Series A (General), 146(1):85–86, ???? 1983. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2981491.

Anonymous:1983:PSGa

[318] Anonymous. Proceedings of a Special General Meeting of the Royal Statistical Society. Journal of the Royal Statistical Society. Series A (General), 146(1): 87, ???? 1983. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2981492.

Chalmers:1983:BRB

[319] C. P. Chalmers. Book review: The Theory of Linear Models and Multivariate Analysis, by Steven F. Arnold. Journal of the Royal Statistical Society. Series A (General), 146(1):88–89, ???? 1983. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2981493.

Selby:1983:BRBa

[320] M. J. P. Selby. Book review: Handbook of Mathematical Economics, Vol. 1, by K. J. Arrow; M. D. Intriligator. Journal of the Royal Statistical Society. Series A (General), 146(1):89–90, ???? 1983. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2981494.

Read:1983:BRBa

[321] K. Read. Book review: Statistical Theory and Inference in Research, by T. A. Bancroft; C. P. Han. Journal of the Royal Statistical Society. Series A (General), 146(1):90, ???? 1983. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981495.

Laycock:1983:BRB

[322] P. J. Laycock. Book review: Interpreting Multivariate Data, by Vic Barnett. Journal of the Royal Statistical Society. Series A (General), 146(1):90–91, ???? 1983. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981496.

Read:1983:BRBb

[323] K. Read. Book review: Comparative Statistical Inference, by Vic Barnett. Journal of the Royal Statistical Society. Series A (General), 146(1):91, ???? 1983. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981497.

Gill:1983:BRB

[324] Christine Gill. Book review: Circular Statistics in Biology, by E. Batschelet. Journal of the Royal Statistical Society. Series A (General), 146(1):91–92, ???? 1983. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981498.

Page:1983:BRBa

[325] E. Page. Book review: The Single Server Queue, by J. W. Cohen. Journal of the Royal Statistical Society. Series A (General), 146(1):92, ???? 1983. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981499.

Chatfield:1983:BRB

[326] C. Chatfield. Book review: Applied Time Series Analysis II, by David F.

Findley. Journal of the Royal Statistical Society. Series A (General), 146(1): 93, ???? 1983. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981500.

Knott:1983:BRB

[327] Martin Knott. Book review: Abstract Inference, by U. Grenander. Journal of the Royal Statistical Society. Series A (General), 146(1):93–94, ???? 1983. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981501.

Sugden:1983:BRB

[328] R. A. Sugden. Book review: Sampling From a Finite Population, by J. Hajek. Journal of the Royal Statistical Society. Series A (General), 146(1):94–95, ???? 1983. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981502.

Selby:1983:BRBb

[329] M. J. P. Selby. Book review: Handbook of the Normal Distribution, by J. K. Patel; C. B. Read. Journal of the Royal Statistical Society. Series A (General), 146(1):95, ???? 1983. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981503.

Fryer:1983:BRB

[330] M. J. Fryer. Book review: Simulation and the Monte Carlo Method, by R. Y. Rubinstein. Journal of the Royal Statistical Society. Series A (General), 146 (1):95–96, ???? 1983. CODEN JS-SAEF. ISSN 0035-9238. URL http:/

/www.jstor.org/stable/2981504.

Boreham:1983:JSC

[331] John Boreham. Jack Stafford, CB. Journal of the Royal Statistical Society. Series A (General), 146(1):97–98, ???? 1983. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981505.

Anonymous:1983:BMa

[332] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 146(1):??, ???? 1983. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981506.

Anonymous:1983:FMb

[333] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 146(2):??, ???? 1983. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2982008.

Anonymous:1983:SSE

[334] Anonymous. Sources of statistics on ethnic minorities. Journal of the Royal Statistical Society. Series A (General), 146 (2):99–114, ???? 1983. CODEN JS-SAEF. ISSN 0035-9238. URL http:/

/www.jstor.org/stable/2982009.

Derbyshire:1983:ASM

[335] M. E. Derbyshire. The application of statistical methods in personal social services — a review. Journal of the Royal Statistical Society. Series A (General), 146(2):115–149, ???? 1983. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2982010.

Winkler:1983:CF

[336] Robert L. Winkler and Spyros Makridakis. The combination of forecasts. Journal of the Royal Statistical Society. Series A (General), 146(2):150–157, ???? 1983. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2982011.

Creedy:1983:FEM

[337] J. Creedy. F. Y. Edgeworth’s mathematical training. Journal of the Royal Statistical Society. Series A (General), 146 (2):158–162, ???? 1983. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982012.

Prais:1983:FIT

[338] S. J. Prais. Formal and informal teaching: A further re-consideration of Professor Bennett’s statistics. Journal of the Royal Statistical Society. Series A (General), 146(2):163–169, ???? 1983. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2982013.

Aitkin:1983:CPP

[339] Murray Aitkin. Comment on Professor Prais’s paper. Journal of the Royal Statistical Society. Series A (General), 146 (2):170–171, ???? 1983. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982014.

Leser:1983:SRL

[340] C. E. V. Leser. Short-run and longrun relative price changes. Journal of the Royal Statistical Society. Series A (General), 146(2):172–181, ???? 1983. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2982015.

Hussain:1983:NPR

[341] S. S. Hussain and P. Sprent. Nonparametric regression. Journal of the Royal Statistical Society. Series A (General), 146(2):182–191, ???? 1983. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2982016.

Crowder:1983:BRB

[342] M. J. Crowder. Book review: Asymptotic Efficiency of Statistical Estimators: Concepts and Higher Order Asymptotic Efficiency, by M. Akahira; K. Takeuchi. Journal of the Royal Statistical Society. Series A (General), 146(2):192–193, ???? 1983. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2982017.

Poskitt:1983:BRB

[343] D. S. Poskitt. Book review: Time Series Analysis: Theory and Practice 1, by O. D. Anderson. Journal of the Royal Statistical Society. Series A (General), 146 (2):193–194, ???? 1983. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982018.

Bryant:1983:BRB

[344] J. W. Bryant. Book review: Stochastic Models for Social Processes, by D. J. Bartholomew. Journal of the Royal Statistical Society. Series A (General), 146 (2):194–195, ???? 1983. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982019.

Smith:1983:BRBa

[345] Cedric A. B. Smith. Book review:

The Genetics of Altruism, by Scott A. Boorman; Paul R. Levitt. Journal of the Royal Statistical Society. Series A (General), 146(2):195–196, ???? 1983. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2982020.

Oliver:1983:BRBa

[346] F. R. Oliver. Book review: Incomes Policy, by Robin E. J. Chater; Andrew Dean; Robert F. Elliott. Journal of the Royal Statistical Society. Series A (General), 146(2):196, ???? 1983. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2982021.

Harvey:1983:BRB

[347] A. C. Harvey. Book review: Distributed Lags: Problems of Estimation and Formulation, by P. J. Dhrymes. Journal of the Royal Statistical Society. Series A (General), 146(2):196–197, ???? 1983. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982022.

Alderson:1983:BRB

[348] M. R. Alderson. Book review: The Causes of Cancer: Quantitative Estimates of the Avoidable Risks of Cancer in the United States Today, by R. Doll; R. Peto. Journal of the Royal Statistical Society. Series A (General), 146(2): 197, ???? 1983. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2982023.

Shaw:1983:BRB

[349] K. M. Shaw. Book review: An Introduction to Mathematical Taxonomy, by G. Dunn; B. S. Everitt. Journal of the Royal Statistical Society. Series A (General), 146(2):197–198, ???? 1983. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982024.

Goldie:1983:BRB

[350] C. Goldie. Book review: Contributions to Probability: A Collection of Papers Dedicated to Eugene Lukacs, by J. Gani; V. K. Rohatgi. Journal of the Royal Statistical Society. Series A (General), 146 (2):198–199, ???? 1983. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982025.

Allin:1983:BRB

[351] Paul Allin. Book review: Secondary Analysis in Social Research: A Guide to Data Sources and Methods with Examples, by Catherine Hakim. Journal of the Royal Statistical Society. Series A (General), 146(2):199–200, ???? 1983. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982026.

Ray:1983:BRB

[352] W. D. Ray. Book review: Stochastic Methods of Operations Research, by J. Kohlas. Journal of the Royal Statistical Society. Series A (General), 146(2): 200, ???? 1983. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2982027.

Page:1983:BRBb

[353] Eric Page. Book review: Frontiers in Statistical Quality Control, by H. J. Lenz; G. B. Wetherill; P. Th. Wilrich. Journal of the Royal Statistical Society. Series A (General), 146(2):200–201, ???? 1983. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2982028.

Slater:1983:BRB

[354] M. Slater. Book review: Introduction to Data Analysis and Statistical Inference, by Carl N. Morris; John E. Rolph. Journal of the Royal Statistical Society. Series A (General), 146(2):201, ???? 1983. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2982029.

Curram:1983:BRB

[355] J. Curram. Book review: Acceptance Sampling in Quality Control, by E. G. Schilling. Journal of the Royal Statistical Society. Series A (General), 146 (2):201–202, ???? 1983. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2982030.

Isham:1983:BRB

[356] Valerie Isham. Book review: Non-Negative Matrices and Markov Chains, by E. Seneta. Journal of the Royal Statistical Society. Series A (General), 146(2):202, ???? 1983. CODEN JS-SAEF. ISSN 0035-9238. URL http:/

                                        /www.jstor.org/stable/2982031.

Smith:1983:BRBb

[357] J. Q. Smith. Book review: Prior Information in the Linear Model, by H. Toutenberg. Journal of the Royal Statistical Society. Series A (General), 146 (2):202–203, ???? 1983. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2982032.

Sprent:1983:BRB

[358] P. Sprent. Book review: Recent Advances in Regression Methods, by H. D. Vinod; A. Ullah. Journal of the Royal Statistical Society. Series A (General), 146(2):203, ???? 1983. CODEN JS-SAEF. ISSN 0035-9238. URL http:/

/www.jstor.org/stable/2982033.

Holt:1983:BRB

[359] D. Holt. Book review: Sampling Methods of Censuses and Survey, by F. Yates. Journal of the Royal Statistical Society. Series A (General), 146(2):203–204, ???? 1983. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2982034.

Box:1983:GMJ

[360] G. E. P. Box. G. M. Jenkins, 1933– 1982. Journal of the Royal Statistical Society. Series A (General), 146(2):205– 206, ???? 1983. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2982035.

Finney:1983:MRS

[361] D. J. Finney. Michael Robert Sampford, 1925–1983. Journal of the Royal Statistical Society. Series A (General), 146 (2):206–207, ???? 1983. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982036.

Anonymous:1983:BMb

[362] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 146(2):??, ???? 1983. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2982037.

Anonymous:1983:FMc

[363] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 146(3):??, ???? 1983. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981650.

Bartlett:1983:TK

[364] M. S. Bartlett. A tribute to A. N. Kolmogorov. Journal of the Royal Statistical Society. Series A (General), 146 (3):209–210, ???? 1983. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2981651.

Kendall:1983:THC

[365] David Kendall. A tribute to Harald Cramér. Journal of the Royal Statistical Society. Series A (General), 146 (3):211–212, ???? 1983. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2981652.

Moore:1983:HEN

[366] Peter G. Moore. Higher education: The next decade. Journal of the Royal Statistical Society. Series A (General), 146 (3):213–245, ???? 1983. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2981653.

Felsenstein:1983:SIP

[367] Joseph Felsenstein. Statistical inference of phylogenies. Journal of the Royal Statistical Society. Series A (General), 146 (3):246–272, ???? 1983. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2981654.

Greenfield:1983:EDR

[368] C. C. Greenfield. On estimators for dual record systems. Journal of the Royal Statistical Society. Series A (General), 146(3):273–280, ???? 1983. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2981655.

Walker:1983:SPI

[369] Monica A. Walker. Some problems in interpreting statistics relating to crime.

Journal of the Royal Statistical Society. Series A (General), 146(3):281–293, ???? 1983. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981656.

Haberman:1983:MPB

[370] Steven Haberman, Barry A. Kosmin, and Caren Levy. Mortality patterns of British Jews 1975–79: Insights and applications for the size and structure of British Jewry. Journal of the Royal Statistical Society. Series A (General), 146 (3):294–310, ???? 1983. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981657.

Izenman:1983:JRW

[371] Alan Julian Izenman. J. R. Wolf and H. A. Wolfer: An historical note on the Zurich sunspot relative numbers. Journal of the Royal Statistical Society. Series A (General), 146(3):311–318, ???? 1983. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981658.

Anonymous:1983:ASM

[372] Anonymous. The application of statistical methods in personal social services — a review. Journal of the Royal Statistical Society. Series A (General), 146 (3):319–320, ???? 1983. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981659.

Anonymous:1983:BMc

[373] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 146(3):??, ???? 1983. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981660.

Anonymous:1983:VI

[374] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (General), 146(4):441–444, ???? 1983. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2981449.

Anonymous:1983:FMd

[375] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 146(4):??, ???? 1983. CODEN JSSAEF. ISSN 0035-9238.

                                        URL http://www.jstor.org/stable/
                                        2981450.

Armitage:1983:TEE

[376] P. Armitage. Trials and errors: The emergence of clinical statistics. Journal of the Royal Statistical Society. Series A (General), 146(4):321–334, ???? 1983. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2981451.

Deville:1983:DAO

[377] J.-C. Deville and E. Malinvaud. Data analysis in official socio-economic statistics. Journal of the Royal Statistical Society. Series A (General), 146(4):335– 361, ???? 1983. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2981452.

Lewis:1983:CTS

[378] John A. Lewis. Clinical trials: Statistical developments of practical benefit to the pharmaceutical industry. Journal of the Royal Statistical Society. Series A (General), 146(4):362–393, ???? 1983. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2981453.

Smith:1983:VIN

[379] T. M. F. Smith. On the validity of inferences from non-random sample. Journal of the Royal Statistical Society. Series A (General), 146(4):394–403, ???? 1983. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981454.

Anonymous:1983:PAG

[380] Anonymous. Proceedings of the 149th Annual General Meeting of the Royal Statistical Society. Journal of the Royal Statistical Society. Series A (General), 146(4):404–405, ???? 1983. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981455.

Anonymous:1983:PSGb

[381] Anonymous. Proceedings of a Special General Meeting of the Royal Statistical Society. Journal of the Royal Statistical Society. Series A (General), 146(4): 406, ???? 1983. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981456.

Anonymous:1983:HTR

[382] Anonymous. Honorary Treasurer’s report. Journal of the Royal Statistical Society. Series A (General), 146(4):407, ???? 1983. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981457.

Anonymous:1983:ARC

[383] Anonymous. Annual report of the Council. Journal of the Royal Statistical Society. Series A (General), 146(4):408–429, ???? 1983. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981458.

Michael:1983:BRB

[384] P. Michael. Book review: The Rational Expectations Revolution in Macroeconomics, by D. K. H. Begg; The Philosophy of Economics. by W. Stegmüller; W. Balzer; W. Spohn. Journal of the Royal Statistical Society. Series A (General), 146(4):430–431, ???? 1983. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2981459.

Hill:1983:BRB

[385] I. D. Hill. Book review: Quotes, Damned Quotes, and …, by John Bibby. Journal of the Royal Statistical Society. Series A (General), 146(4):431, ???? 1983. CODEN JSSAEF. ISSN 0035-9238.

                                        URL http://www.jstor.org/stable/
                                        2981460.

Kennedy:1983:BRB

[386] D. P. Kennedy. Book review: Rates of Convergence in the Central Limit Theorem, by P. Hall. Journal of the Royal Statistical Society. Series A (General), 146(4):431–432, ???? 1983. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2981461.

Titterington:1983:BRB

[387] D. M. Titterington. Book review: Theory of Statistical Experiments, by H. Heyer. Journal of the Royal Statistical Society. Series A (General), 146(4):432, ???? 1983. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2981462.

Oliver:1983:BRBb

[388] F. R. Oliver. Book review: Introduction to Economics: Theory and Data, by Dudley Jackson. Journal of the Royal Statistical Society. Series A (General), 146(4):432–433, ???? 1983. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981463.

Maine:1983:BRB

[389] N. P. Maine. Book review: Modern Data Analysis, by R. L. Launer; A. F. Siegel. Journal of the Royal Statistical Society. Series A (General), 146(4):433– 434, ???? 1983. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981464.

Charlton:1983:BRB

[390] John R. H. Charlton. Book review: The Randomised Clinical Trial and Therapeutic Decisions, by N. Tygstrup; J. M. Lechin; E. Juhl. Journal of the Royal Statistical Society. Series A (General), 146(4):434–435, ???? 1983. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981465.

Hartley:1983:BRB

[391] R. Hartley. Book review: Optimisation over Time: Dynamic Programming and Stochastic Control, by P. Whittle. Journal of the Royal Statistical Society. Series A (General), 146(4):435–436, ???? 1983. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981466.

Bather:1983:BRB

[392] J. A. Bather. Book review: The Design and Analysis of Sequential Clinical Trials, by J. Whitehead. Journal of the Royal Statistical Society. Series A (General), 146(4):436–437, ???? 1983. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981467.

Taylor:1983:JB

[393] William F. Taylor. Joseph Berkson, 1899–1982. Journal of the Royal Statistical Society. Series A (General), 146 (4):438–439, ???? 1983. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2981468.

Webb:1983:HD

[394] Norman L. Webb. Henry Durant, 1903– 1982. Journal of the Royal Statistical Society. Series A (General), 146(4):439– 440, ???? 1983. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2981469.

Anonymous:1983:BMd

[395] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 146(4):??, ???? 1983. CODEN JSSAEF. ISSN 0035-9238.

                                        URL http://www.jstor.org/stable/
                                        2981470.

Anonymous:1984:FMa

[396] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 147(1):??, ???? 1984. CODEN JSSAEF. ISSN 0035-9238.

                                        URL http://www.jstor.org/stable/
                                        2981735.

Stern:1984:MFA

[397] R. D. Stern and R. Coe. A model fitting analysis of daily rainfall data. Journal of the Royal Statistical Society. Series A (General), 147(1):1–34, ???? 1984. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2981736.

Spiegelhalter:1984:SKB

[398] David J. Spiegelhalter and Robin P.

Knill-Jones. Statistical and knowledgebased approaches to clinical decisionsupport systems, with an application in gastroenterology. Journal of the Royal Statistical Society. Series A (General), 147(1):35–77, ???? 1984. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981737.

Bawden:1984:CBO

[399] David Bawden. Computer-based online information retrieval for statistical literature. Journal of the Royal Statistical Society. Series A (General), 147(1):78– 86, ???? 1984. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981738.

Woolson:1984:ACI

[400] Robert F. Woolson and William R.

Clarke. Analysis of categorical incomplete longitudinal data. Journal of the Royal Statistical Society. Series A (General), 147(1):87–99, ???? 1984. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2981739.

Draper:1984:BWC

[401] Norman R. Draper. The Box–Wetz criterion versus R^{2}. Journal of the Royal Statistical Society. Series A (General), 147(1):100–103, ???? 1984. CODEN JS-SAEF. ISSN 0035-9238. URL http:/

/www.jstor.org/stable/2981740. See

corrections [615].

Embrechts:1984:BRB

[402] P. A. L. Embrechts. Book review: Combinatorial Integral Geometry, with Applications to Mathematical Stereology, by R. V. Ambartzumian. Journal of the Royal Statistical Society. Series A (General), 147(1):104, ???? 1984. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2981741.

Morton-Williams:1984:BRB

[403] Jean Morton-Williams. Book review: The Design and Understanding of Survey Questions, by William A. Belson. Journal of the Royal Statistical Society. Series A (General), 147(1):105, ???? 1984. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981742.

Fisk:1984:BRB

[404] Peter R. Fisk. Book review: Fix-Point Estimation in Theory and Practice, by R. Bergstrom; H. Wold. Journal of the Royal Statistical Society. Series A (General), 147(1):105–106, ???? 1984. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981743.

Goldstein:1984:BRBa

[405] M. Goldstein. Book review: Bayesian Statistics, by J. M. Bernardo. Journal of the Royal Statistical Society. Series A (General), 147(1):106–107, ???? 1984. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981744.

Stone:1984:BRB

[406] M. Stone. Book review: A Festschrift for Eric Lehmann: In Honour of his Sixty-Fifth Birthday, by P. Bickel; K. Doksum; J. Hodges. Journal of the Royal Statistical Society. Series A (General), 147 (1):107–108, ???? 1984. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981745.

Robinson:1984:BRB

[407] Anthony Robinson. Book review: Residuals and Influence in Regression, by R. Dennis Cook; Sanford Weisberg. Journal of the Royal Statistical Society. Series A (General), 147(1):108, ???? 1984. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981746.

Hand:1984:BRB

[408] D. J. Hand. Book review: Statistics and Related Topics, by M. Csorgo; D. A. Dawson; J. N. K. Rao; A. K. Md. E. Saleh. Journal of the Royal Statistical Society. Series A (General), 147 (1):108–109, ???? 1984. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981747.

Osborn:1984:BRB

[409] John F. Osborn. Book review: Biostatistics: A Foundation of Analysis in the Health Sciences, by Wayne W. Daniel. Journal of the Royal Statistical Society. Series A (General), 147(1):109–110, ???? 1984. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981748.

Davis:1984:BRB

[410] M. H. A. Davis. Book review: Stochastic Calculus and Applications, by R. J. Elliott. Journal of the Royal Statistical Society. Series A (General), 147(1):110, ???? 1984. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981749.

Smith:1984:BRBa

[411] J. Q. Smith. Book review: Specifying Statistical Models: From Parametric to Non-Parametric using Bayesian or Non-Bayesian Approaches, by J. P. Florens. Journal of the Royal Statistical Society. Series A (General), 147(1): 111, ???? 1984. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981750.

Reuter:1984:BRB

[412] G. E. H. Reuter. Book review: Markov Chains, by D. Freedman; Brownian Motion and Diffusion, by D. Freedman; Approximation Countable Chains, by D. Freedman. Journal of the Royal Statistical Society. Series A (General), 147 (1):111–112, ???? 1984. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981751.

Thomas:1984:BRBa

[413] G. E. Thomas. Book review: Matrices with Applications in Statistics, by F. A. Graybill. Journal of the Royal Statistical Society. Series A (General), 147 (1):112–113, ???? 1984. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981752.

Collett:1984:BRB

[414] D. Collett. Book review: Understanding Robust and Exploratory Data Analysis, by David C. Hoaglin; Frederick Mosteller; John W. Tukey. Journal of the Royal Statistical Society. Series A (General), 147(1):113, ???? 1984. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981753.

Hammersley:1984:BRB

[415] J. M. Hammersley. Book review: Percolation Theory for Mathematicians, by Harry Kesten. Journal of the Royal Statistical Society. Series A (General), 147 (1):113–114, ???? 1984. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981754.

Bather:1984:BRB

[416] J. A. Bather. Book review: Probability, Statistics and Analysis, by J. F. C. Kingman; G. E. H. Reuter. Journal of the Royal Statistical Society. Series A (General), 147(1):114–115, ???? 1984. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981755.

Wilson:1984:BRB

[417] D. A. L. Wilson. Book review: Short Term Forecasting: An Introduction to the Box-Jenkins Approach, by Thomas M. O’Donovan. Journal of the Royal Statistical Society. Series A (General), 147(1):115, ???? 1984. CODEN JS-SAEF. ISSN 0035-9238. URL http:/

/www.jstor.org/stable/2981756.

Campbell:1984:BRB

[418] M. J. Campbell. Book review: Environmental Epidemiology: Risk Assessment, by Ross L. Prentice; Alice S. Whittemore. Journal of the Royal Statistical Society. Series A (General), 147(1):115– 116, ???? 1984. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981757.

Yates:1984:BRB

[419] F. Yates. Book review: Neyman–From Life, by Constance Reid. Journal of the Royal Statistical Society. Series A (General), 147(1):116–118, ???? 1984. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981758.

Healy:1984:BRBa

[420] M. J. R. Healy. Book review: Newer Concepts in Nutrition and Their Implications for Policy, by P. V. Sukhatme. Journal of the Royal Statistical Society. Series A (General), 147(1):118–119, ???? 1984. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2981759.

Renshaw:1984:BRB

[421] Eric Renshaw. Book review: Capture-Recapture and Removal Methods for Sampling Closed Populations, by G. C. White; D. R. Anderson; K. P. Burnham; D. L. Otis. Journal of the Royal Statistical Society. Series A (General), 147(1):119, ???? 1984. CODEN JS-SAEF. ISSN 0035-9238. URL http:/

                                        /www.jstor.org/stable/2981760.

Plackett:1984:JAA

[422] R. L. Plackett. John Anthony Anderson, 1939–1983. Journal of the Royal Statistical Society. Series A (General), 147(1):120, ???? 1984. CODEN JS-SAEF. ISSN 0035-9238. URL http:/

                                        /www.jstor.org/stable/2981761.

Stuart:1984:SMK

[423] Alan Stuart. Sir Maurice Kendall, 1907– 1983. Journal of the Royal Statistical Society. Series A (General), 147(1):120– 122, ???? 1984. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2981762.

Anonymous:1984:BMa

[424] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 147(1):??, ???? 1984. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981763.

Anonymous:1984:FMb

[425] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 147(2):??, ???? 1984. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981668.

Anonymous:1984:P

[426] Anonymous. Proceedings. Journal of the Royal Statistical Society. Series A (General), 147(2):123–129, ???? 1984. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981669.

Hill:1984:SSL

[427] I. D. Hill. Statistical Society of London– Royal Statistical Society: The first 100 years: 1834–1934. Journal of the Royal Statistical Society. Series A (General), 147(2):130–139, ???? 1984. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981670.

Plackett:1984:RSS

[428] R. L. Plackett. Royal Statistical Society: The last fifty years; 1934–84. Journal of the Royal Statistical Society. Series A (General), 147(2):140–150, ???? 1984. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981671.

Nelder:1984:PPP

[429] J. A. Nelder. Present position and potential developments: Some personal views: Statistical computing. Journal of the Royal Statistical Society. Series A (General), 147(2):151–160, ????

1984. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981672.

Durbin:1984:PPP

[430] J. Durbin. Present position and potential developments: Some personal views: Time series analysis. Journal of the Royal Statistical Society. Series A (General), 147(2):161–173, ???? 1984. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981673.

Boreham:1984:PPP

[431] John Boreham. Present position and potential developments: Some personal views: Official statistics. Journal of the Royal Statistical Society. Series A (General), 147(2):174–185, ???? 1984. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981674.

Newell:1984:PPP

[432] D. J. Newell. Present position and potential developments: Some personal views: Medical statistics. Journal of the Royal Statistical Society. Series A (General), 147(2):186–197, ???? 1984. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981675.

Sibson:1984:PPP

[433] Robin Sibson. Present position and potential developments: Some personal views: Multivariate analysis. Journal of the Royal Statistical Society. Series A (General), 147(2):198–207, ???? 1984. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981676.

Smith:1984:PPPa

[434] T. M. F. Smith. Present position and potential developments: Some personal views: Sample surveys. Journal of the Royal Statistical Society. Series A (General), 147(2):208–221, ???? 1984. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981677.

Thatcher:1984:TPI

[435] A. R. Thatcher. Topic of public interest:

The 1981 census of population in England and Wales. Journal of the Royal Statistical Society. Series A (General), 147(2):222–232, ???? 1984. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981678.

Kingman:1984:PPP

[436] J. F. C. Kingman. Present position and potential developments: Some personal views: Probability and random processes. Journal of the Royal Statistical Society. Series A (General), 147 (2):233–244, ???? 1984. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981679.

Smith:1984:PPPb

[437] A. F. M. Smith. Present position and potential developments: Some personal views: Bayesian statistics. Journal of the Royal Statistical Society. Series A (General), 147(2):245–259, ???? 1984. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981680.

Goldstein:1984:PPP

[438] Harvey Goldstein. Present position and potential developments: Some personal views: Statistics in the social sciences.

Journal of the Royal Statistical Society. Series A (General), 147(2):260–267, ???? 1984. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981681.

Moore:1984:PPP

[439] Peter G. Moore. Present position and potential developments: Some personal views: Statistics in business and commerce. Journal of the Royal Statistical Society. Series A (General), 147(2):268– 277, ???? 1984. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981682.

Dawid:1984:PPP

[440] A. P. Dawid. Present position and potential developments: Some personal views: Statistical theory: The prequential approach. Journal of the Royal Statistical Society. Series A (General), 147 (2):278–292, ???? 1984. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981683.

Beck:1984:TPI

[441] M. B. Beck. Topic of public interest: Water quality. Journal of the Royal Statistical Society. Series A (General), 147 (2):293–305, ???? 1984. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981684.

Cox:1984:PPP

[442] D. R. Cox. Present position and potential developments: Some personal views: Design of experiments and regression. Journal of the Royal Statistical Society. Series A (General), 147(2):306–315, ???? 1984. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981685.

Baines:1984:PPP

[443] A. Baines. Present position and potential developments: Some personal views: Industrial statistics and operational research. Journal of the Royal Statistical Society. Series A (General), 147(2):316– 326, ???? 1984. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981686.

Hendry:1984:PPP

[444] David F. Hendry. Present position and potential developments: Some personal views: Time-series econometrics. Journal of the Royal Statistical Society. Series A (General), 147(2):327–339, ???? 1984. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981687.

Ripley:1984:PPP

[445] B. D. Ripley. Present position and potential developments: Some personal views: Statistics in the natural sciences. Journal of the Royal Statistical Society. Series A (General), 147(2):340–348, ???? 1984. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981688.

Curnow:1984:PPP

[446] R. N. Curnow. Present position and potential developments: Some personal views: Statistics in biometry and agriculture. Journal of the Royal Statistical Society. Series A (General), 147(2):349– 358, ???? 1984. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981689.

Lindley:1984:PFN

[447] D. V. Lindley. Prospects for the future:

The next 50 years. Journal of the Royal Statistical Society. Series A (General), 147(2):359–367, ???? 1984. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2981690.

Healy:1984:PFW

[448] M. J. R. Healy. Prospects for the future: Where has statistics failed? Journal of the Royal Statistical Society. Series A (General), 147(2):368–374, ???? 1984. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2981691.

Rosenbaum:1984:GRS

[449] S. Rosenbaum. The growth of the Royal Statistical Society. Journal of the Royal Statistical Society. Series A (General), 147(2):375–388, ???? 1984. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2981692.

Anonymous:1984:BMb

[450] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 147(2):??, ???? 1984. CODEN JSSAEF. ISSN 0035-9238.

                                        URL http://www.jstor.org/stable/
                                        2981693.

Anonymous:1984:FMc

[451] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 147(3):??, ???? 1984. CODEN JSSAEF. ISSN 0035-9238.

                                        URL http://www.jstor.org/stable/
                                        2981575.

Miller:1984:SSR

[452] Alan J. Miller. Selection of subsets of regression variables. Journal of the Royal Statistical Society. Series A (General), 147(3):389–425, ???? 1984. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981576.

Yates:1984:TSC

[453] F. Yates. Tests of significance for 2 × 2 contingency tables. Journal of the Royal Statistical Society. Series A (General), 147(3):426–463, ???? 1984. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981577.

Lawrence:1984:LDD

[454] R. J. Lawrence. The lognormal distribution of the duration of strikes. Journal of the Royal Statistical Society. Series A (General), 147(3):464–483, ???? 1984. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981578.

Weiss:1984:ETS

[455] A. A. Weiss and A. P. Andersen. Estimating time series models using the relevant forecast evaluation criterion. Journal of the Royal Statistical Society. Series A (General), 147(3):484–487, ???? 1984. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981579.

Xekalaki:1984:BGW

[456] Evdokia Xekalaki. The bivariate generalized Waring distribution and its application to accident theory. Journal of the Royal Statistical Society. Series A (General), 147(3):488–498, ???? 1984. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981580.

Bezdek:1984:TTH

[457] Roger H. Bezdek. Tests of three hypotheses relating to the Leontief inputoutput model. Journal of the Royal Statistical Society. Series A (General), 147 (3):499–509, ???? 1984. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2981581.

Anonymous:1984:LR

[458] Anonymous. [list of reviews]. Journal of the Royal Statistical Society. Series A (General), 147(3):510, ???? 1984. CODEN JSSAEF. ISSN 0035-9238.

                                        URL http://www.jstor.org/stable/
                                        2981582.

Ray:1984:BRBa

[459] W. D. Ray. Book review: Statistical Methods for Forecasting, by Bovas Abraham; Johannes Ledolter. Journal of the Royal Statistical Society. Series A (General), 147(3):511, ???? 1984. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2981583.

Radcliffe:1984:BRB

[460] J. Radcliffe. Book review: The Biomathematics of Malaria, by Norman T. J. Bailey. Journal of the Royal Statistical Society. Series A (General), 147(3):511– 512, ???? 1984. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2981584.

Daldry:1984:BRBa

[461] S. J. Daldry. Book review: Introductory Applied Probability, by G. P. Beaumont. Journal of the Royal Statistical Society. Series A (General), 147(3):512, ???? 1984. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2981585.

Smith:1984:BRBb

[462] R. P. Smith. Book review: Political Aspects of the Economy, by V. K. Borooah; F. Van der Ploeg. Journal of the Royal Statistical Society. Series A (General), 147(3):512–513, ???? 1984. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981586.

Healy:1984:BRBb

[463] M. J. R. Healy. Book review: Graphical Methods for Data Analysis, by J. M. Chambers; W. S. Cleveland; B. Kleiner; P. A. Tukey. Journal of the Royal Statistical Society. Series A (General), 147(3): 513, ???? 1984. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981587.

Latham:1984:BRB

[464] F. R. Latham. Book review: Statistics for Technology (A Course in Applied Statistics), by C. Chatfield. Journal of the Royal Statistical Society. Series A (General), 147(3):513–514, ???? 1984. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981588.

Bartlett:1984:BRB

[465] M. S. Bartlett. Book review: Pao-Lu Hsu: Collected Papers by K. L. Chung; C.-S. Cheng; T.-P. Chiang. Journal of the Royal Statistical Society. Series A (General), 147(3):514–515, ???? 1984. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981589.

Cox:1984:BRB

[466] T. F. Cox. Book review: Spatial Processes: Models and Applications, by A. D. Cliff; J. K. Ord. Journal of the Royal Statistical Society. Series A (General), 147(3):515, ???? 1984. CODEN JS-SAEF. ISSN 0035-9238. URL http:/

/www.jstor.org/stable/2981590.

Stuart:1984:BRB

[467] Alan Stuart. Book review: Contribution to Statistics, by William G. Cochran. Journal of the Royal Statistical Society. Series A (General), 147(3):515, ???? 1984. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2981591.

Critchley:1984:BRB

[468] Frank Critchley. Book review: (a) The User’s Guide to Multidimensional Scaling, by A. P. M. Coxon; P. M. Davies;

(b) Key texts in Multidimensional Scaling. by P. M. Davies; A. P. M. Coxon. Journal of the Royal Statistical Society. Series A (General), 147(3):516, ???? 1984. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2981592.

Upton:1984:BRB

[469] G. J. G. Upton. Book review: An Introduction to Statistical Modelling, by Annette J. Dobson. Journal of the Royal Statistical Society. Series A (General), 147(3):516, ???? 1984. CODEN JS-SAEF. ISSN 0035-9238. URL http:/

                                        /www.jstor.org/stable/2981593.

Newell:1984:BRB

[470] David Newell. Book review: Evidence, Proof and Probability, by Richard Eggleston. Journal of the Royal Statistical Society. Series A (General), 147(3):517, ???? 1984. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2981594.

Silverman:1984:BRB

[471] B. W. Silverman. Book review: Advanced Methods of Data Exploration and Modelling, by B. S. Everitt; G. Dunn. Journal of the Royal Statistical Society. Series A (General), 147(3):517–518, ???? 1984. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981595.

Barnard:1984:BRB

[472] G. A. Barnard. Book review: The Making of Statisticians, by J. Gani. Journal of the Royal Statistical Society. Series A (General), 147(3):518–519, ???? 1984. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981596.

Oakes:1984:BRB

[473] David Oakes. Book review: Statistics in Practice, by Sheila M. Gore; Douglas G. Altman. Journal of the Royal Statistical Society. Series A (General), 147(3):519, ???? 1984. CODEN JS-SAEF. ISSN 0035-9238. URL http:/

/www.jstor.org/stable/2981597.

Smith:1984:BRBc

[474] J. Q. Smith. Book review: Statistical Decision Theory and Related Topics III, by S. S. Gupta; J. O. Berger. Journal of the Royal Statistical Society. Series A (General), 147(3):520, ???? 1984. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981598.

Loynes:1984:BRB

[475] R. M. Loynes. Book review: Martingale Limit Theory and its Application, by P. Hall; C. C. Heyde. Journal of the Royal Statistical Society. Series A (General), 147(3):520–521, ???? 1984. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981599.

Ray:1984:BRBb

[476] W. D. Ray. Book review: Recent Trends in Statistics, by S. Heiler. Journal of the Royal Statistical Society. Series A (General), 147(3):521, ???? 1984. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2981600.

Higgins:1984:BRBa

[477] J. C. Higgins. Book review: Risk Analysis and its Applications, by D. B. Hertz; H. Thomas. Journal of the Royal Statistical Society. Series A (General), 147 (3):521–522, ???? 1984. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2981601.

Curnow:1984:BRB

[478] R. N. Curnow. Book review: Utilitas Mathematica Volumes 21A and 21B, May 1982, by Frank Yates; A. M. Herzberg. Journal of the Royal Statistical Society. Series A (General), 147(3): 522, ???? 1984. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2981602.

Healy:1984:BRBc

[479] M. J. R. Healy. Book review: New Trends in Data Analysis and Applications, by J. Janssen; J. F. Marcotorchino; J. M. Proth. Journal of the Royal Statistical Society. Series A (General), 147(3):522–523, ???? 1984. CODEN JSSAEF. ISSN 0035-9238.

                                        URL http://www.jstor.org/stable/
                                        2981603.

Ray:1984:BRBc

[480] W. D. Ray. Book review: The Advanced Theory of Statistics, Vol. 3: Design and Analysis and Time Series, by M. G. Kendall; A. Stuart; J. K. Ord. Journal of the Royal Statistical Society. Series A (General), 147(3):523, ???? 1984. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981604.

Barton:1984:BRB

[481] D. E. Barton. Book review: Encyclopedia of Statistical Sciences Vol. I: A-Circular Probable Error, by S. Kotz; N. L. Johnson; Encyclopedia of Statistical Sciences Vol. II: Classification-Eye Estimate, by S. Kotz; N. L. Johnson. Journal of the Royal Statistical Society. Series A (General), 147(3):524, ???? 1984. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981605.

Ray:1984:BRBd

[482] W. D. Ray. Book review: Developments in Statistics, Vol. 4, by Paruchuri R. Krishnaiah. Journal of the Royal Statistical Society. Series A (General), 147(3): 525, ???? 1984. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981606.

Thomas:1984:BRBb

[483] G. E. Thomas. Book review: A Course in Linear Models, by A. M. Kshirsagar. Journal of the Royal Statistical Society. Series A (General), 147(3):525–526, ???? 1984. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981607.

Ray:1984:BRBe

[484] W. D. Ray. Book review: Stationary Processes in Time Series Analysis: The Mathematical Foundations, by P. J. Lambert; D. S. Poskett. Journal of the Royal Statistical Society. Series A (General), 147(3):526, ???? 1984. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2981608.

Plackett:1984:BRB

[485] R. L. Plackett. Book review: Generalized Linear Models, by P. McCullagh; J. A. Nelder. Journal of the Royal Statistical Society. Series A (General), 147 (3):526–527, ???? 1984. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2981609.

Whittle:1984:BRB

[486] P. Whittle. Book review: Prediction Theory and Harmonic Analysis, by V. Mandrekar; H. Salehi. Journal of the Royal Statistical Society. Series A (General), 147(3):527, ???? 1984. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2981610.

Oliver:1984:BRB

[487] F. R. Oliver. Book review: Slower Growth in the Western World, by R. C. O. Matthews. Journal of the Royal Statistical Society. Series A (General), 147 (3):527–528, ???? 1984. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2981611.

Thomas:1984:BRBc

[488] G. E. Thomas. Book review: Statistical Methods in Agriculture and Experimental Biology, by R. Mead; R. N. Curnow. Journal of the Royal Statistical Society. Series A (General), 147 (3):528–529, ???? 1984. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2981612.

Higgins:1984:BRBb

[489] J. C. Higgins. Book review: The Business of Risk, by P. G. Moore. Journal of the Royal Statistical Society. Series A (General), 147(3):529, ???? 1984. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981613.

Pike:1984:BRB

[490] D. J. Pike. Book review: Applied Linear Statistical Methods, by Donald F. Morrison. Journal of the Royal Statistical Society. Series A (General), 147(3):530, ???? 1984. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981614.

Harvey:1984:BRB

[491] Andrew Harvey. Book review: Random Coefficient Autoregressive Models: An Introduction, by D. F. Nicholls; B. G. Quinn. Journal of the Royal Statistical Society. Series A (General), 147 (3):530–531, ???? 1984. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981615.

Daldry:1984:BRBb

[492] S. J. Daldry. Book review: Theoretical Exercises in Probability and Statistics, by N. A. Rahman. Journal of the Royal Statistical Society. Series A (General), 147(3):531, ???? 1984. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2981616.

Healy:1984:BRBd

[493] M. J. R. Healy. Book review: Non-Linear Regression Modelling: A Unified Practical Approach, by D. A. Ratkowsky. Journal of the Royal Statistical Society. Series A (General), 147 (3):531–532, ???? 1984. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2981617.

Martin:1984:BRB

[494] R. J. Martin. Book review: Spatial Statistics, by Brian D. Ripley. Journal of the Royal Statistical Society. Series A (General), 147(3):532–533, ???? 1984. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2981618.

Janacek:1984:BRB

[495] G. J. Janacek. Book review: Physical Applications of Stationary Time-Series, by E. A. Robinson. Journal of the Royal Statistical Society. Series A (General), 147(3):533–534, ???? 1984. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2981619.

Smith:1984:BRBd

[496] Cedric A. B. Smith. Book review: Contributions to Statistics: Essays in Honour of Professor Norman L. Johnson, by P. K. Sen. Journal of the Royal Statistical Society. Series A (General), 147(3): 534, ???? 1984. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2981620.

Ripley:1984:BRB

[497] B. D. Ripley. Book review: Statistics on Spheres, by Geoffrey S. Watson. Journal of the Royal Statistical Society. Series A (General), 147(3):534–535, ???? 1984. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2981621.

Gittins:1984:BRB

[498] J. C. Gittins. Book review: Optimality and Efficiency, by D. J. White. Journal of the Royal Statistical Society. Series A (General), 147(3):535–536, ???? 1984. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981622.

Benjamin:1984:PRE

[499] Bernard Benjamin. Professor Robert Eric Beard, OBE, 1911–1983. Journal of the Royal Statistical Society. Series A (General), 147(3):537, ???? 1984. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981623.

Mentz:1984:CED

[500] Raul Pedro Mentz. Carlos Eugenio Dieulefait, 1901–1982. Journal of the Royal Statistical Society. Series A (General), 147(3):537–538, ???? 1984. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981624.

Corlett:1984:PGG

[501] Tom Corlett. Percy George Gray, 1916– 1983. Journal of the Royal Statistical Society. Series A (General), 147(3):538– 539, ???? 1984. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981625.

Anonymous:1984:BMc

[502] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 147(3):??, ???? 1984. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981626.

Anonymous:1984:FMd

[503] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 147(4):??, ???? 1984.

CODEN JSSAEF. ISSN 0035-9238.

                                        URL http://www.jstor.org/stable/
                                        2981842.

Taffler:1984:CRM

[504] R. J. Taffler and B. Abassi. Country risk: A model for predicting debt servicing problems in developing countries. Journal of the Royal Statistical Society. Series A (General), 147(4):541– 568, ???? 1984. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2981843.

Anonymous:1984:AEP

[505] Anonymous. Assessment of examination performance in different types of schools. Journal of the Royal Statistical Society. Series A (General), 147 (4):569–581, ???? 1984. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2981844.

Nour:1984:MSM

[506] El-Sayed Nour and C. M. Suchindran. Multi-state mortality by cause of death: A life table analysis. Journal of the Royal Statistical Society. Series A (General), 147(4):582–597, ???? 1984. CODEN JSSAEF. ISSN 0035-9238.

                                        URL http://www.jstor.org/stable/
                                        2981845.

Benjamin:1984:SUC

[507] B. Benjamin. The Statistics Users’ Council: A note. Journal of the Royal Statistical Society. Series A (General), 147(4):598–599, ???? 1984. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2981846.

West:1984:BA

[508] Mike West. Bayesian aggregation. Journal of the Royal Statistical Society. Series A (General), 147(4):600–607, ???? 1984. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981847.

Healy:1984:UMG

[509] M. J. R. Healy. The use of R^{2} as a measure of goodness of fit. Journal of the Royal Statistical Society. Series A (General), 147(4):608–609, ???? 1984. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981848.

Taylor:1984:BRB

[510] S. J. Taylor. Book review: [untitled].

Journal of the Royal Statistical Society. Series A (General), 147(4):610, ???? 1984. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981849.

Blight:1984:BRB

[511] P. A. Blight. Book review: [untitled].

Journal of the Royal Statistical Society. Series A (General), 147(4):610–611, ???? 1984. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981850.

Diggle:1984:BRB

[512] Peter J. Diggle. Book review: Computing in Statistical Science through APL, by Francis Anscombe. Journal of the Royal Statistical Society. Series A (General), 147(4):611, ???? 1984. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2981851.

Osmond:1984:BRB

[513] Clive Osmond. Book review: Regression Analysis of Survival Data in Cancer Chemotherapy, by Walter H. Carter; Galen L. Wampler; Donald M. Stablein. Journal of the Royal Statistical Society. Series A (General), 147(4):611–612, ???? 1984. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981852.

Farewell:1984:BRB

[514] V. T. Farewell. Book review: Statistical Methods for Cancer Studies, by R. G. Cornell. Journal of the Royal Statistical Society. Series A (General), 147 (4):612–613, ???? 1984. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981853.

Davidson:1984:BRBa

[515] James Davidson. Book review: Applications of Econometrics, by Julia Hebden. Journal of the Royal Statistical Society. Series A (General), 147(4):613– 614, ???? 1984. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981854.

Bunn:1984:BRB

[516] P. J. Bunn. Book review: Dice, Data and Decisions, by K. W. Kemp. Journal of the Royal Statistical Society. Series A (General), 147(4):614, ???? 1984. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981855.

Goldstein:1984:BRBb

[517] M. Goldstein. Book review: Introduction to Probability Theory and Statistical Inference, by Harold J. Larson. Journal of the Royal Statistical Society. Series A (General), 147(4):614–615, ???? 1984. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981856.

Brown:1984:BRB

[518] Philip J. Brown. Book review: Theory of Point Estimation, by E. L. Lehmann. Journal of the Royal Statistical Society. Series A (General), 147(4):615, ???? 1984. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981857.

Sparrow:1984:BRB

[519] Colin Sparrow. Book review: The Fractal Geometry of Nature, by B. Mandelbrot. Journal of the Royal Statistical Society. Series A (General), 147(4):616– 618, ???? 1984. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981858.

Fearn:1984:BRB

[520] Tom Fearn. Book review: Food Research and Data Analysis, by Harald Martens; Hellmut Russwurm. Journal of the Royal Statistical Society. Series A (General), 147(4):618–619, ???? 1984. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981859.

Curram:1984:BRB

[521] J. B. Curram. Book review: Attribute Sampling Plans, Tables of Tests and Confidence Limits for Proportions, by R. E. Odeh; D. B. Owen. Journal of the Royal Statistical Society. Series A (General), 147(4):619, ???? 1984. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2981860.

Doney:1984:BRB

[522] R. A. Doney. Book review: A First Course in Probability, by Sheldon Ross. Journal of the Royal Statistical Society. Series A (General), 147(4):619–620, ???? 1984. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2981861.

Poskitt:1984:BRB

[523] D. S. Poskitt. Book review: Prediction and Regulation, by Linear Least-Squares Methods. by Peter Whittle. Journal of the Royal Statistical Society. Series A (General), 147(4):620, ???? 1984. CODEN JSSAEF. ISSN 0035-9238.

                                        URL http://www.jstor.org/stable/
                                        2981862.

Anonymous:1984:BMd

[524] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 147(4):??, ???? 1984. CODEN JSSAEF. ISSN 0035-9238.

                                        URL http://www.jstor.org/stable/
                                        2981863.

Anonymous:1984:VI

[525] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (General), 147(5):??, ???? 1984. CODEN JSSAEF. ISSN 0035-9238.

                                        URL http://www.jstor.org/stable/
                                        2981694.

Anonymous:1984:FMe

[526] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 147(5):??, ???? 1984. CODEN JSSAEF. ISSN 0035-9238.

                                        URL http://www.jstor.org/stable/
                                        2981695.

Goodhardt:1984:DCM

[527] G. J. Goodhardt, A. S. C. Ehrenberg, and C. Chatfield. The Dirichlet: A comprehensive model of buying behaviour. Journal of the Royal Statistical Society. Series A (General), 147(5):621–655, ???? 1984. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981696.

Rosenbaum:1984:CAC

[528] Paul R. Rosenbaum. The consequences of adjustment for a concomitant variable that has been affected by the treatment. Journal of the Royal Statistical Society. Series A (General), 147(5):656– 666, ???? 1984. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981697.

Anonymous:1984:PSG

[529] Anonymous. Proceedings of a Special General Meeting of the Royal Statistical Society. Journal of the Royal Statistical Society. Series A (General), 147(5): 667, ???? 1984. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981698.

Anonymous:1984:POH

[530] Anonymous. Proceedings of the one hundred and fiftieth Annual General Meeting of the Royal Statistical Society. Journal of the Royal Statistical Society. Series A (General), 147(5):668, ???? 1984. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981699.

Anonymous:1984:PR

[531] Anonymous. President’s report. Journal of the Royal Statistical Society. Series A (General), 147(5):669–670, ???? 1984. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981700.

Anonymous:1984:HTR

[532] Anonymous. Honorary Treasurer’s report. Journal of the Royal Statistical Society. Series A (General), 147(5):671, ???? 1984. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2981701.

Anonymous:1984:ARC

[533] Anonymous. Annual report of the Council. Journal of the Royal Statistical Society. Series A (General), 147(5):671–696, ???? 1984. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2981702.

Stoodley:1984:BRB

[534] Keith D. C. Stoodley. Book review: Time Series Analysis: Theory and Practice 2, by O. D. Anderson. Journal of the Royal Statistical Society. Series A (General), 147(5):697, ???? 1984. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2981703.

Ray:1984:BRBf

[535] W. D. Ray. Book review: Handbook of Statistics, Vol. 3: Time Series in the Frequency Domain, by D. R. Brillinger; P. R. Krishnaiah. Journal of the Royal Statistical Society. Series A (General), 147(5):697–698, ???? 1984. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2981704.

Davidson:1984:BRBb

[536] James Davidson. Book review: Handbook of Econometrics, Volume 1, by Zvi Griliches; Michael D. Intriligator; Kenneth J. Arrow; Michael D. Intriligator. Journal of the Royal Statistical Society. Series A (General), 147(5):698–699, ???? 1984. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981705.

Anderson:1984:BRB

[537] C. W. Anderson. Book review: Extremes and Related Properties of Random Sequences and Processes, by M. R. Leadbetter; G. Lindgren; H. Rootzen. Journal of the Royal Statistical Society. Series A (General), 147(5):699–700, ???? 1984. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981706.

Daldry:1984:BRBc

[538] S. J. Daldry. Book review: Statistical Inference, by V. K. Rohatgi. Journal of the Royal Statistical Society. Series A (General), 147(5):700–701, ???? 1984. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981707.

Moran:1984:BRB

[539] M. A. Moran. Book review: Multivariate Observations, by G. A. F. Seber. Journal of the Royal Statistical Society. Series A (General), 147(5):701, ???? 1984. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981708.

Bulpitt:1984:BRB

[540] C. J. Bulpitt. Book review: Clinical Trials–Issues and Approaches, by S. H. Shapiro; T. A. Louis. Journal of the Royal Statistical Society. Series A (General), 147(5):702, ???? 1984. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2981709.

Godall:1984:BRB

[541] Edward A. Godall. Book review: The Ideas of Sampling, by A. Stuart. Journal of the Royal Statistical Society. Series A (General), 147(5):702–703, ???? 1984. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/

stable/2981710. See correction [606].

Bunday:1984:BRB

[542] Brian D. Bunday. Book review: Games, Theory and Applications, by L. C. Thomas. Journal of the Royal Statistical Society. Series A (General), 147(5): 703, ???? 1984. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2981711.

Roger:1984:BRB

[543] James Roger. Book review: The Visual Display of Quantitative Information, by Edward R. Tufte. Journal of the Royal Statistical Society. Series A (General), 147(5):704, ???? 1984. CODEN JS-SAEF. ISSN 0035-9238. URL http:/

                                        /www.jstor.org/stable/2981712.

Hartley:1984:BRB

[544] R. Hartley. Book review: Optimisation Over Time: Dynamic Programming and Stochastic Control, by Peter Whittle. Journal of the Royal Statistical Society. Series A (General), 147(5):704–705, ???? 1984. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2981713.

Grebenik:1984:RGD

[545] E. Grebenik. Roy George Douglas Allen 1906–1983. Journal of the Royal Statistical Society. Series A (General), 147 (5):706–707, ???? 1984. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981714.

Kempthorne:1984:GT

[546] O. Kempthorne, K. A. Fox, and W. A.

Fuller. Gerhard Tintner, 1907–1983. Journal of the Royal Statistical Society. Series A (General), 147(5):707–708, ???? 1984. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981715.

Anonymous:1984:BMe

[547] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 147(5):??, ???? 1984. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981716.

Anonymous:1985:FMa

[548] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 148(1):??, ???? 1985. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981507.

Bennett:1985:SSS

[549] R. J. Bennett and R. P. Haining. Spatial structure and spatial interaction: Modelling approaches to the statistical analysis of geographical data. Journal of the Royal Statistical Society. Series A (General), 148(1):1–36, ???? 1985. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2981508.

Oldham:1985:FSR

[550] P. D. Oldham. The fluoridation of the Strathclyde Regional Council’s water supply: Opinion of Lord Jauncey in causa Mrs Catherine McColl against Strathclyde Regional Council: A review. Journal of the Royal Statistical Society. Series A (General), 148(1):37–44, ???? 1985. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981509.

Aitkin:1985:SMU

[551] Murray Aitkin and Robert Healey. Statistical modelling of unemployment rates from the EEC Labour Force Survey. Journal of the Royal Statistical Society. Series A (General), 148(1):45–56, ???? 1985. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981510.

Plackett:1985:TSH

[552] R. L. Plackett. A tribute to Sir Harry Campion. Journal of the Royal Statistical Society. Series A (General), 148 (1):57–58, ???? 1985. CODEN JS-SAEF. ISSN 0035-9238. URL http:/

/www.jstor.org/stable/2981511.

Ray:1985:BRBa

[553] W. D. Ray. Book review: Time Series Analysis: Theory and Practice, by O. D. Anderson. Journal of the Royal Statistical Society. Series A (General), 148(1): 59, ???? 1985. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981512.

Dawid:1985:BRB

[554] A. P. Dawid. Book review: Robustness of Bayesian Analyses, by Joseph B. Kadane. Journal of the Royal Statistical Society. Series A (General), 148(1):59– 60, ???? 1985. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981513.

Duncombe:1985:BRB

[555] P. Duncombe. Book review: Multivariate Descriptive Statistical Analysis: Correspondence Analysis and Related Techniques for Large Matrices, by L. Lebart; A. Morineau; K. M. Warwick. Journal of the Royal Statistical Society. Series A (General), 148(1):60– 61, ???? 1985. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981514.

Stoodley:1985:BRB

[556] K. D. C. Stoodley. Book review: Time Series and Systems Analysis with Applications, by S. M. Pandit; S. M. Wu. Journal of the Royal Statistical Society. Series A (General), 148(1):61–62, ???? 1985. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981515.

Palmer:1985:BRB

[557] M. K. Palmer. Book review: Clinical Trials: A Pratical Approach, by S. J. Pocock. Journal of the Royal Statistical Society. Series A (General), 148(1):62– 63, ???? 1985. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981516.

Barnett:1985:BRB

[558] Vic Barnett. Book review: Statistics Theory and Practice. Selected Papers, by Maurice Kendall (1907–1983). by Alan Stuart. Journal of the Royal Statistical Society. Series A (General), 148(1):63, ???? 1985. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981517.

Ray:1985:BRBb

[559] W. D. Ray. Book review: Statistical Signal Processing, by Edward J. Wegman; James G. Smith. Journal of the Royal Statistical Society. Series A (General), 148(1):63, ???? 1985. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2981518.

Carroll:1985:BRBa

[560] Patrick Carroll. Book review: Practical Statistical Sampling for Auditors, by Arthur J. Wilburn. Journal of the Royal Statistical Society. Series A (General), 148(1):63–64, ???? 1985. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2981519.

Daniels:1985:FD

[561] Henry Daniels. Frank Downton 1925– 84. Journal of the Royal Statistical Society. Series A (General), 148(1):65–66, ???? 1985. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2981520.

Bartlett:1985:HLS

[562] Maurice Bartlett. Hilary Latham seal, 1911–1984. Journal of the Royal Statistical Society. Series A (General), 148(1): 66, ???? 1985. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2981521.

Schmetterer:1985:WW

[563] L. Schmetterer. Wilheim Winkler, 1884– 1984. Journal of the Royal Statistical Society. Series A (General), 148(1):67, ???? 1985. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2981522.

Anonymous:1985:BMa

[564] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 148(1):??, ???? 1985. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981523.

Anonymous:1985:FMb

[565] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 148(2):??, ???? 1985. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981941.

Bodmer:1985:US

[566] W. F. Bodmer. Understanding statistics. Journal of the Royal Statistical Society. Series A (General), 148(2):69–81, ???? 1985. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981942.

Clayton:1985:MGP

[567] David Clayton and Jack Cuzick. Multivariate generalizations of the proportional hazards model. Journal of the Royal Statistical Society. Series A (General), 148(2):82–117, ???? 1985. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981943.

Seneta:1985:SHS

[568] E. Seneta. A sketch of the history of survey sampling in Russia. Journal of the Royal Statistical Society. Series A (General), 148(2):118–125, ???? 1985. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981944.

Brant:1985:UAE

[569] J. D. Brant and S. M. Chalk. The use of automatic editing in the 1981 census. Journal of the Royal Statistical Society. Series A (General), 148(2):126– 146, ???? 1985. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981945.

McDonald:1985:SAD

[570] John McDonald and G. D. Snooks. Statistical analysis of Domesday Book (1086). Journal of the Royal Statistical Society. Series A (General), 148(2):147– 160, ???? 1985. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981946.

Anonymous:1985:PSGa

[571] Anonymous. Proceedings of a Special General Meeting of the Royal Statistical Society. Journal of the Royal Statistical Society. Series A (General), 148(2): 161, ???? 1985. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981947.

Anonymous:1985:RCG

[572] Anonymous. Robert Charles Geary, 1896–1983. Journal of the Royal Statistical Society. Series A (General), 148(2): 162, ???? 1985. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981948.

Whittaker:1985:BRB

[573] J. Whittaker. Book review: Analysis of Ordinal Categorical Data, by Alan Agresti. Journal of the Royal Statistical Society. Series A (General), 148 (2):163–164, ???? 1985. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981949.

Skinner:1985:BRBa

[574] C. J. Skinner. Book review: An Introduction to Multivariate Statistical Analysis, by T. W. Anderson. Journal of the Royal Statistical Society. Series A (General), 148(2):164–165, ???? 1985. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981950.

Littlewood:1985:BRB

[575] Bev Littlewood. Book review: Repairable Systems Reliability: Modelling, Inference, Misconceptions and their Causes, by H. Ascher; H. Feingold. Journal of the Royal Statistical Society. Series A (General), 148(2):165, ???? 1985. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981951.

Kimber:1985:BRBa

[576] Alan Kimber. Book review: Outliers in Statistical Data, by Vic Barnett; Toby Lewis. Journal of the Royal Statistical Society. Series A (General), 148(2):165– 166, ???? 1985. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981952.

Farlie:1985:BRB

[577] D. J. G. Farlie. Book review: Statistical Methods in Research and Production with Special Reference to the Chemical Industry, by O. L. Davies; P. L. Goldsmith. Journal of the Royal Statistical Society. Series A (General), 148 (2):166–167, ???? 1985. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981953.

Skinner:1985:BRBb

[578] C. J. Skinner. Book review: An Introduction to Latent Variable Models, by B. S. Everitt. Journal of the Royal Statistical Society. Series A (General), 148(2): 167, ???? 1985. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2981954.

Poskitt:1985:BRB

[579] D. S. Poskitt. Book review: Alternative Approaches to Time Series Analysis, by J. P. Florens; M. Mouchart; J. P. Raouls; L. Simar. Journal of the Royal Statistical Society. Series A (General), 148(2):167–168, ???? 1985. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2981955.

Farrow:1985:BRB

[580] M. Farrow. Book review: Statistical Modelling, by Warren Gilchrist. Journal of the Royal Statistical Society. Series A (General), 148(2):168, ???? 1985. CODEN JSSAEF. ISSN 0035-9238.

                                        URL http://www.jstor.org/stable/
                                        2981956.

Plackett:1985:BRB

[581] R. L. Plackett. Book review: The Analysis of Cross-Classified Data having Ordered Categories, by Leo A. Goodman; Clifford C. Clogg; Otis Dudley Duncan. Journal of the Royal Statistical Society. Series A (General), 148(2):168– 169, ???? 1985. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2981957.

Shand:1985:BRB

[582] Alexander H. Shand. Book review: Money, Capital and Fluctuations: Early Essays, by F. A. Hayek; Roy Mc-Cloughrey. Journal of the Royal Statistical Society. Series A (General), 148 (2):169–170, ???? 1985. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981958.

Emmett:1985:BRB

[583] B. P. Emmett. Book review: Introduction to Causal Analysis: Exploring Survey Data by Cross-Tabulation, by Ottar Hellevik. Journal of the Royal Statistical Society. Series A (General), 148(2): 170, ???? 1985. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981959.

Gardner:1985:BRB

[584] M. J. Gardner. Book review: A Short Textbook of Medical Statistics, by Austin Bradford Hill. Journal of the Royal Statistical Society. Series A (General), 148 (2):170–171, ???? 1985. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981960.

Hill:1985:BRBa

[585] I. D. Hill. Book review: Charles Babbage — Pioneer of the Computer, by Anthony Hyman. Journal of the Royal Statistical Society. Series A (General), 148(2):171, ???? 1985. CODEN JS-SAEF. ISSN 0035-9238. URL http:/

/www.jstor.org/stable/2981961.

Ray:1985:BRBc

[586] W. D. Ray. Book review: Elements of Simulation, by Byron J. T. Morgan. Journal of the Royal Statistical Society. Series A (General), 148(2):172, ???? 1985. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981962.

Rao:1985:BRB

[587] M. Bhaskara Rao. Book review: General Irreducible Markov Chains and Non-Negative Operators, by E. Nummelin. Journal of the Royal Statistical Society. Series A (General), 148(2):172, ???? 1985. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2981963.

Spooner:1985:BRB

[588] A. Spooner. Book review: W. G. Cochran’s Impact on Statistics, by P. S. R. S. Rao; J. Sedransk. Journal of the Royal Statistical Society. Series A (General), 148(2):172–173, ???? 1985. CODEN JSSAEF. ISSN 0035-9238.

                                        URL http://www.jstor.org/stable/
                                        2981964.

Barlow:1985:BRB

[589] M. T. Barlow. Book review: Probability, by A. N. Shiryayev; R. P. Boas. Journal of the Royal Statistical Society. Series A (General), 148(2):173, ???? 1985. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2981965.

Anonymous:1985:BMb

[590] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 148(2):??, ???? 1985. CODEN JSSAEF. ISSN 0035-9238.

                                        URL http://www.jstor.org/stable/
                                        2981966.

Anonymous:1985:FMc

[591] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 148(3):??, ???? 1985. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981967.

Smith:1985:PSN

[592] T. M. F. Smith, W. B. Wakefield, I. Diamond, and D. E. Barton. Projections of student numbers in higher education. Journal of the Royal Statistical Society. Series A (General), 148(3):175–213, ???? 1985. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981968.

Chatfield:1985:IED

[593] C. Chatfield. The initial examination of data. Journal of the Royal Statistical Society. Series A (General), 148 (3):214–253, ???? 1985. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981969.

Narendranthan:1985:IID

[594] W. Narendranthan, S. Nickell, and D. Metcalf. An investigation into the incidence and dynamic structure of sickness and unemployment in Britain, 1965–75. Journal of the Royal Statistical Society. Series A (General), 148 (3):254–267, ???? 1985. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981970.

Pfeffermann:1985:RCC

[595] D. Pfeffermann and D. J. Holmes. Robustness considerations in the choice of a method of inference for regression analysis of survey data. Journal of the Royal Statistical Society. Series A (General), 148(3):268–278, ???? 1985. CODEN JS-SAEF. ISSN 0035-9238. URL http:/

/www.jstor.org/stable/2981971. See

corrections [616].

Campion:1985:JWN

[596] H. Campion. J. W. Nixon, 1888–1985. Journal of the Royal Statistical Society. Series A (General), 148(3):279, ???? 1985. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2981972.

Priestley:1985:BRB

[597] M. B. Priestley. Book review: The Collected Works of John W. Tukey. Volume 1: Time Series, 1949–1964, by D. R. Brillinger. Journal of the Royal Statistical Society. Series A (General), 148(3): 280, ???? 1985. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2981973.

Ray:1985:BRBd

[598] W. D. Ray. Book review: Compstat Lectures. Volumes 2 and 3. Lectures in Computational Statistics, by J. M. Chambers; J. Gordesch; A. Klas; L. Lebart; P. P. Sint. Journal of the Royal Statistical Society. Series A (General), 148(3):280–281, ???? 1985. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2981974.

vanderPloeg:1985:BRB

[599] F. van der Ploeg. Book review: Operations Research and Economic Theory (Essays in Honor of Martin J. Beckmann), by H. Hauptmann; W. Krelle; K. C. Mosler. Journal of the Royal Statistical Society. Series A (General), 148(3): 281, ???? 1985. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2981975.

Anderson:1985:BRB

[600] C. W. Anderson. Book review: Mathematical Techniques of Applied Probability. Volume 1 — Discrete Time Models: Basic Theory, by J. J. Hunter; Mathematical Techniques of Applied Probability. Volume 2 — Discrete Time Models: Techniques and Applications, by J. J. Hunter. Journal of the Royal Statistical Society. Series A (General), 148 (3):282–283, ???? 1985. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981976.

Kennedy:1985:BRB

[601] D. P. Kennedy. Book review: Introduction to Probability Theory, by Kiyosi Ito. Journal of the Royal Statistical Society. Series A (General), 148(3):283, ???? 1985. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981977.

Sweeting:1985:BRB

[602] T. J. Sweeting. Book review: Asymptotic Statistics II, by P. Mandl; M. Huskova. Journal of the Royal Statistical Society. Series A (General), 148 (3):283–284, ???? 1985. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981978.

Kimber:1985:BRBb

[603] Alan Kimber. Book review: Statistical Theory and Data Analysis, by K. Matusita. Journal of the Royal Statistical Society. Series A (General), 148(3):284, ???? 1985. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981979.

Carroll:1985:BRBb

[604] Patrick Carroll. Book review: Risk Theory, by R. E. Beard; T. Pentikainen; E. Pesonen. Journal of the Royal Statistical Society. Series A (General), 148(3):

285, ???? 1985. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2981980.

Everitt:1985:BRB

[605] B. Everitt. Book review: Cluster Dissection and Analysis: Theory, Fortran Programs and Examples, by H. Spath. Journal of the Royal Statistical Society. Series A (General), 148(3):285–286, ???? 1985. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2981981.

Anonymous:1985:CIS

[606] Anonymous. Correction: The ideas of sampling. Journal of the Royal Statistical Society. Series A (General), 148(3): 286, ???? 1985. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2981982. See [541].

Anonymous:1985:BMc

[607] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 148(3):??, ???? 1985. CODEN JSSAEF. ISSN 0035-9238.

                                        URL http://www.jstor.org/stable/
                                        2981983.

Anonymous:1985:VI

[608] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (General), 148(4):399–400, ???? 1985. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2981889.

Anonymous:1985:FMd

[609] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 148(4):??, ???? 1985. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981890.

Black:1985:ICP

[610] Robert W. Black. Instead of the 1986 census: The potential contribution of enhanced electoral registers. Journal of the Royal Statistical Society. Series A (General), 148(4):287–316, ???? 1985. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981891.

Suissa:1985:EUS

[611] Samy Suissa and Jonathan J. Shuster.

Exact unconditional sample sizes for the 2×2 binomial trial. Journal of the Royal Statistical Society. Series A (General), 148(4):317–327, ???? 1985. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981892.

Holmes:1985:ALS

[612] D. I. Holmes. The analysis of literary style — a review. Journal of the Royal Statistical Society. Series A (General), 148(4):328–341, ???? 1985. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981893.

Henery:1985:APL

[613] Robert J. Henery. On the average probability of losing bets on horses with given starting price odds. Journal of the Royal Statistical Society. Series A (General), 148(4):342–349, ???? 1985. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981894.

Newby:1985:FND

[614] Martin Newby. A further note on the distribution of strike duration. Journal of the Royal Statistical Society. Series A (General), 148(4):350–356, ????

1985. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2981895.

Draper:1985:CBW

[615] Norman R. Draper. Corrections: The Box–Wetz criterion versus R^{2}. Journal of the Royal Statistical Society. Series A (General), 148(4):357, ???? 1985. CODEN JSSAEF. ISSN 0035-9238.

                                        URL http://www.jstor.org/stable/

2981896. See [401].

Pfeffermann:1985:CRC

[616] D. Pfeffermann and D. Holmes. Corrections: Robustness considerations in the choice of a method of inference for regression analysis of survey data. Journal of the Royal Statistical Society. Series A (General), 148(4):357, ???? 1985. CODEN JSSAEF. ISSN 0035-9238.

                                        URL http://www.jstor.org/stable/

2981897. See [595].

Anonymous:1985:PSGb

[617] Anonymous. Proceedings of a Special General Meeting of the Royal Statistical Society. Journal of the Royal Statistical Society. Series A (General), 148(4): 358, ???? 1985. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2981898.

Anonymous:1985:PAG

[618] Anonymous. Proceedings of the 151st Annual General Meeting of the Royal Statistical Society. Journal of the Royal Statistical Society. Series A (General), 148(4):359–360, ???? 1985. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2981899.

Anonymous:1985:PR

[619] Anonymous. President’s report. Journal of the Royal Statistical Society. Series A (General), 148(4):361–364, ???? 1985. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981900.

Anonymous:1985:HTR

[620] Anonymous. Honarary Treasurer’s report. Journal of the Royal Statistical Society. Series A (General), 148(4):365, ???? 1985. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981901.

Anonymous:1985:ARC

[621] Anonymous. Annual report of the Council. Journal of the Royal Statistical Society. Series A (General), 148(4):366–388, ???? 1985. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981902.

Martin:1985:BRB

[622] R. Martin. Book review: Time Series Analysis: Theory and Practice 6, by O. D. Anderson; J. K. Ord; E. A. Robinson. Journal of the Royal Statistical Society. Series A (General), 148(4):389– 390, ???? 1985. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981903.

Jones:1985:BRB

[623] Byron Jones. Book review: Statistical Methods for Engineers and Scientists, by R. M. Bethea; B. S. Duran; T. L. Boullion. Journal of the Royal Statistical Society. Series A (General), 148(4):390, ???? 1985. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981904.

Smith:1985:BRB

[624] A. F. M. Smith. Book review: Bayesian Analysis of Linear Models, by L. D. Broemeling. Journal of the Royal Statistical Society. Series A (General), 148 (4):390–391, ???? 1985. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2981905.

Woodward:1985:BRB

[625] M. Woodward. Book review: Applied Regression Analysis and Experimental Design, by R. J. Brook; G. C. Arnold. Journal of the Royal Statistical Society. Series A (General), 148(4):391, ???? 1985. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2981906.

Lancaster:1985:BRB

[626] T. Lancaster. Book review: Evaluating the Reliability of Econometric Models, by G. C. Chow; P. Corsi. Journal of the Royal Statistical Society. Series A (General), 148(4):391–392, ???? 1985. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2981907.

Taylor:1985:BRB

[627] C. C. Taylor. Book review: Nonparametric Density Estimation: The L_{1} View, by Luc Devroye; Laszlo Gyorfi. Journal of the Royal Statistical Society. Series A (General), 148(4):392–393, ???? 1985. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2981908.

Laxton:1985:BRB

[628] R. R. Laxton. Book review: Interval Orders and Interval Graphs: A Study of Partially Ordered Sets, by P. C. Fishburn. Journal of the Royal Statistical Society. Series A (General), 148(4):393– 394, ???? 1985. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981909.

Maclaren:1985:BRB

[629] N. M. Maclaren. Book review: Applied Statistics Algorithms, by P. Griffiths; A. D. Hill. Journal of the Royal Statistical Society. Series A (General), 148(4): 394, ???? 1985. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981910.

Cox:1985:BRB

[630] David Cox. Book review: Contributions to Probability and Statistics in Honour of Gunnar Blom, by J. Lanke; G. Lindgren. Journal of the Royal Statistical Society. Series A (General), 148(4):394– 395, ???? 1985. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981911.

Hill:1985:BRBb

[631] I. D. Hill. Book review: Human Experimentation: A Guided Step into the Unknown, by W. A. Silverman. Journal of the Royal Statistical Society. Series A (General), 148(4):395, ???? 1985. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981912.

Tremayne:1985:BRB

[632] A. R. Tremayne. Book review: Mathematical Methods in Economics, by F. van der Ploeg. Journal of the Royal Statistical Society. Series A (General), 148 (4):395–396, ???? 1985. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981913.

Bunday:1985:BRB

[633] B. D. Bunday. Book review: Model Building in Mathematical Programming, by H. P. Williams. Journal of the Royal Statistical Society. Series A (General), 148(4):396, ???? 1985. CODEN JS-SAEF. ISSN 0035-9238. URL http:/

                                        /www.jstor.org/stable/2981914.

Boreham:1985:SAG

[634] John Boreham. Simon Abraham Goldberg (1914–1985). Journal of the Royal Statistical Society. Series A (General), 148(4):397, ???? 1985. CODEN JS-SAEF. ISSN 0035-9238. URL http:/

                                        /www.jstor.org/stable/2981915.

Anonymous:1985:BMd

[635] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 148(4):??, ???? 1985. CODEN JSSAEF. ISSN 0035-9238.

                                        URL http://www.jstor.org/stable/
                                        2981916.

Anonymous:1986:FMa

[636] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 149(1):??, ???? 1986. CODEN JSSAEF. ISSN 0035-9238.

                                        URL http://www.jstor.org/stable/
                                        2981881.

Aitkin:1986:SMI

[637] M. Aitkin and N. Longford. Statistical modelling issues in school effectiveness studies. Journal of the Royal Statistical Society. Series A (General), 149(1):1– 43, ???? 1986. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2981882.

Ford:1986:HCT

[638] J. E. Ford. L. H. C. Tippett, 1902– 1985. Journal of the Royal Statistical Society. Series A (General), 149(1):44, ???? 1986. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981883.

Duffy:1986:LCM

[639] J. C. Duffy and R. W. Latcham. Liver cirrhosis mortality in England and Wales compared to Scotland: An age-period-cohort analysis 1941–81. Journal of the Royal Statistical Society. Series A (General), 149(1):45–59, ???? 1986. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2981884.

Conway:1986:SS

[640] Freda Conway. Statistics in schools.

Journal of the Royal Statistical Society. Series A (General), 149(1):60–64, ???? 1986. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981885.

Kalton:1986:SRP

[641] Graham Kalton and Dallas W. Anderson. Sampling rare populations. Journal of the Royal Statistical Society. Series A (General), 149(1):65–82, ???? 1986. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981886.

Sowey:1986:TCB

[642] Eric R. Sowey. A third classified bibliography on random number generation and testing. Journal of the Royal Statistical Society. Series A (General), 149 (1):83–107, ???? 1986. CODEN JS-SAEF. ISSN 0035-9238. URL http:/

/www.jstor.org/stable/2981887.

Anonymous:1986:BMa

[643] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 149(1):??, ???? 1986. CODEN JSSAEF. ISSN 0035-9238.

                                        URL http://www.jstor.org/stable/
                                        2981888.

Anonymous:1986:FMb

[644] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 149(2):??, ???? 1986. CODEN JSSAEF. ISSN 0035-9238.

                                        URL http://www.jstor.org/stable/
                                        2981524.

Nelder:1986:SST

[645] J. A. Nelder. Statistics, science and technology. Journal of the Royal Statistical Society. Series A (General), 149 (2):109–121, ???? 1986. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2981525.

Anonymous:1986:RJW

[646] Anonymous. The report of a Joint Working Party of the Royal Statistical Society and the Institute of Statisticians on statisticians: Supply and demand. Journal of the Royal Statistical Society. Series A (General), 149(2):122–145, ???? 1986. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2981526.

Swerdlow:1986:CRE

[647] A. J. Swerdlow. Cancer registration in England and Wales: Some aspects relevant to interpretation of the data. Journal of the Royal Statistical Society. Series A (General), 149(2):146–160, ???? 1986. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981527.

Chambers:1986:DAP

[648] R. L. Chambers. Design-adjusted parameter estimation. Journal of the Royal Statistical Society. Series A (General), 149(2):161–173, ???? 1986. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981528.

Pruscha:1986:NTS

[649] Helmut Pruscha. A note on time series analysis of yearly temperature data. Journal of the Royal Statistical Society. Series A (General), 149(2):174–185, ???? 1986. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981529.

Kingman:1986:HC

[650] J. F. C. Kingman. Harald Cramér, 1893– 1985. Journal of the Royal Statistical Society. Series A (General), 149(2):186, ???? 1986. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981530.

Anonymous:1986:BMb

[651] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 149(2):??, ???? 1986. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981531.

Anonymous:1986:FMc

[652] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 149(3):??, ???? 1986. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981552.

Harvey:1986:ESB

[653] A. C. Harvey and J. Durbin. The effects of seat belt legislation on British road casualties: A case study in structural time series modelling. Journal of the Royal Statistical Society. Series A (General), 149(3):187–227, ???? 1986. CODEN JSSAEF. ISSN 0035-9238.

                                        URL http://www.jstor.org/stable/
                                        2981553.

Hart:1986:IHI

[654] Nicky Hart. Inequalities in health: The individual versus the environment. Journal of the Royal Statistical Society. Series A (General), 149(3):228–246, ???? 1986. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2981554.

Plackett:1986:OSA

[655] R. L. Plackett. The old statistical account. Journal of the Royal Statistical Society. Series A (General), 149(3):247– 251, ???? 1986. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2981555.

Lindley:1986:BF

[656] D. V. Lindley. Bruno de Finetti, 1906– 1985. Journal of the Royal Statistical Society. Series A (General), 149(3):252, ???? 1986. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2981556.

Jolliffe:1986:CCA

[657] I. T. Jolliffe, B. Jones, and B. J. T. Morgan. Comparison of cluster analyses of the English personal social services authorities. Journal of the Royal Statistical Society. Series A (General), 149 (3):253–270, ???? 1986. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981557.

Bather:1986:BRB

[658] J. A. Bather. Book review: Bandit Problems: Sequential Allocation of Experiments, by D. A. Berry; B. Fristedt. Journal of the Royal Statistical Society. Series A (General), 149(3):271, ???? 1986. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981558.

Mitchell:1986:BRB

[659] J. M. O. Mitchell. Book review: Computer Science and Statistics: Proceedings of the Sixteenth Symposium on the Interface, by L. Billard. Journal of the Royal Statistical Society. Series A (General), 149(3):272, ???? 1986. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2981559.

Chatfield:1986:BRBa

[660] Chris Chatfield. Book review: Comparative Models for Electrical Load Forecasting, by D. W. Bunn; E. D. Farmer. Journal of the Royal Statistical Society. Series A (General), 149(3):272, ???? 1986. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981560.

Currie:1986:BRBa

[661] Iain D. Currie. Book review: Data Distributions. A Statistical Handbook, by R. A. Christensen. Journal of the Royal Statistical Society. Series A (General), 149(3):273, ???? 1986. CODEN JS-SAEF. ISSN 0035-9238. URL http:/

/www.jstor.org/stable/2981561.

Wetherill:1986:BRB

[662] G. B. Wetherill. Book review: The Element of Graphing Data, by W. S. Cleveland. Journal of the Royal Statistical Society. Series A (General), 149(3):273– 274, ???? 1986. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2981562.

Lewis:1986:BRB

[663] S. M. Lewis. Book review: Orthogonal Fractional Factorial Designs, by Aloke Dey. Journal of the Royal Statistical Society. Series A (General), 149(3):274, ???? 1986. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2981563.

Ray:1986:BRBa

[664] W. D. Ray. Book review: Essays in Time Series and Allied Processes. Papers in Honour of E. J. Hannan, by J. Gani; M. B. Priestley. Journal of the Royal Statistical Society. Series A (General), 149(3):274–275, ???? 1986. CODEN JSSAEF. ISSN 0035-9238.

                                        URL http://www.jstor.org/stable/
                                        2981564.

Leslie:1986:BRB

[665] J. Leslie. Book review: Nonparametric Statistical Inference, by J. D. Gibbons. Journal of the Royal Statistical Society. Series A (General), 149(3):275, ???? 1986. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2981565.

Jolliffe:1986:BRB

[666] Ian Jolliffe. Book review: An Introduction to Probability Theory and Statistical Applications, by M. A. Goldberg. Journal of the Royal Statistical Society. Series A (General), 149(3):275–276, ???? 1986. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981566.

Attfield:1986:BRB

[667] C. L. F. Attfield. Book review: Expectations: Theory and Evidence, by K. Holden; D. A. Peel; J. L. Thompson. Journal of the Royal Statistical Society. Series A (General), 149(3):276, ???? 1986. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981567.

Barnard:1986:BRB

[668] G. A. Barnard. Book review: Making Decisions, by D. V. Lindley. Journal of the Royal Statistical Society. Series A (General), 149(3):276–277, ???? 1986. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981568.

Atkinson:1986:BRB

[669] A. C. Atkinson. Book review: Non-Parametric Methods in General Linear Models, by M. L. Puri; P. K. Sen. Journal of the Royal Statistical Society. Series A (General), 149(3):277–278, ???? 1986. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981569.

Currie:1986:BRBb

[670] Iain D. Currie. Book review: Probability Theory and Applications, by E. A. Robinson. Journal of the Royal Statistical Society. Series A (General), 149(3): 278, ???? 1986. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981570.

Cormack:1986:BRB

[671] R. M. Cormack. Book review: Spatial Data Analysis, by Example. Volume 1: Point Pattern and Quantitative Data. by G. Upton; B. Fingleton. Journal of the Royal Statistical Society. Series A (General), 149(3):279, ???? 1986. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2981571.

Chatfield:1986:BRBb

[672] C. Chatfield. Book review: Applied Linear Regression, by S. Weisberg. Journal of the Royal Statistical Society. Series A (General), 149(3):279–280, ???? 1986. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2981572.

Ray:1986:BRBb

[673] W. D. Ray. Book review: Recursive Estimation and Time Series Analysis, by Peter Young. Journal of the Royal Statistical Society. Series A (General), 149(3):280, ???? 1986. CODEN JS-SAEF. ISSN 0035-9238. URL http:/

                                        /www.jstor.org/stable/2981573.

Anonymous:1986:BMc

[674] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 149(3):??, ???? 1986. CODEN JSSAEF. ISSN 0035-9238.

                                        URL http://www.jstor.org/stable/
                                        2981574.

Anonymous:1986:VI

[675] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (General), 149(4):409, ???? 1986. CODEN JSSAEF. ISSN 0035-9238.

                                        URL http://www.jstor.org/stable/
                                        2981717.

Anonymous:1986:FMd

[676] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 149(4):??, ???? 1986. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981718.

Speed:1986:RSN

[677] T. P. Speed and D. Culpin. The role of statistics in nuclear materials accounting: Issues and problems. Journal of the Royal Statistical Society. Series A (General), 149(4):281–313, ???? 1986. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981719.

Brewer:1986:CSO

[678] A. C. Brewer and R. Mead. Continuous second order models of spatial variation with application to the efficiency of field crop experiments. Journal of the Royal Statistical Society. Series A (General), 149(4):314–348, ???? 1986. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981720.

Stuart:1986:LSS

[679] Alan Stuart. Location-shifts in sampling with unequal probabilities. Journal of the Royal Statistical Society. Series A (General), 149(4):349–365, ???? 1986. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981721.

Mackenzie:1986:PHM

[680] Gilbert Mackenzie. A proportional hazard model for accident data. Journal of the Royal Statistical Society. Series A (General), 149(4):366–375, ????

1986. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2981722.

Anonymous:1986:PR

[681] Anonymous. President’s report. Journal of the Royal Statistical Society. Series A (General), 149(4):376, ???? 1986. CODEN JSSAEF. ISSN 0035-9238.

                                        URL http://www.jstor.org/stable/
                                        2981723.

Anonymous:1986:HTR

[682] Anonymous. Honorary Treasurer’s report. Journal of the Royal Statistical Society. Series A (General), 149(4):377, ???? 1986. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2981724.

Anonymous:1986:ARC

[683] Anonymous. Annual report of the Council. Journal of the Royal Statistical Society. Series A (General), 149(4):377–401, ???? 1986. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2981725.

Anonymous:1986:PAG

[684] Anonymous. Proceedings of the 152nd Annual General Meeting of the Royal Statistical Society. Journal of the Royal Statistical Society. Series A (General), 149(4):402, ???? 1986. CODEN JS-SAEF. ISSN 0035-9238. URL http:/

                                        /www.jstor.org/stable/2981726.

Smith:1986:BRB

[685] David K. Smith. Book review: Mathematics in the Social and Life Sciences: Theories, Models and Methods, by M. A. Ball. Journal of the Royal Statistical Society. Series A (General), 149(4):

403, ???? 1986. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981727.

Porteous:1986:BRB

[686] B. T. Porteous. Book review: Patterns Recognition in Practice II, by E. S. Gelsema; L. N. Kanal. Journal of the Royal Statistical Society. Series A (General), 149(4):403–404, ???? 1986. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981728.

Harvey:1986:BRB

[687] Andrew Harvey. Book review: Time Series in the Time Domain, by E. J. Hannan; P. R. Krishnaiah; M. M. Rao. Journal of the Royal Statistical Society. Series A (General), 149(4):404–405, ???? 1986. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981729.

Macdonald:1986:BRB

[688] R. R. Macdonald. Book review: Design and Analysis of Experiments, by R. G. Petersen. Journal of the Royal Statistical Society. Series A (General), 149 (4):405–406, ???? 1986. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981730.

Morris:1986:BRB

[689] Richard W. Morris. Book review: Introduction to Statistical Methods, Vol. II. Applications to the Life Sciences, by J. S. Rustagi. Journal of the Royal Statistical Society. Series A (General), 149 (4):406–407, ???? 1986. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981731.

Ray:1986:BRBc

[690] W. D. Ray. Book review: Stochastic Modelling and Analysis: A Computational Approach, by H. C. Tijms. Journal of the Royal Statistical Society. Series A (General), 149(4):407, ???? 1986. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2981732.

Lloyd:1986:BRB

[691] Rusell Lloyd. Book review: Directory of Statistical Microcomputer Software, by W. A. Woodward; A. C. Elliot; H. L. Gray. Journal of the Royal Statistical Society. Series A (General), 149(4):407– 408, ???? 1986. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2981733.

Anonymous:1986:BMd

[692] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 149(4):??, ???? 1986. CODEN JSSAEF. ISSN 0035-9238.

                                        URL http://www.jstor.org/stable/
                                        2981734.

Anonymous:1987:FMa

[693] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 150(1):??, ???? 1987. CODEN JSSAEF. ISSN 0035-9238.

                                        URL http://www.jstor.org/stable/
                                        2981661.

Jones:1987:WPP

[694] M. C. Jones and Robin Sibson. What is projection pursuit? Journal of the Royal Statistical Society. Series A (General), 150(1):1–37, ???? 1987. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2981662.

Cox:1987:EML

[695] David Cox and Joyce Snell. Evelyn Martin Lansdowne Beale, F.R.S., 1928– 1985. Journal of the Royal Statistical Society. Series A (General), 150(1):38, ???? 1987. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981663.

Walker:1987:IRC

[696] Monica A. Walker. Interpreting race and crime statistics. Journal of the Royal Statistical Society. Series A (General), 150(1):39–56, ???? 1987. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981664.

Creedy:1987:VER

[697] John Creedy. Variations in earnings and responsibility. Journal of the Royal Statistical Society. Series A (General), 150 (1):57–68, ???? 1987. CODEN JS-SAEF. ISSN 0035-9238. URL http:/

/www.jstor.org/stable/2981665.

Goudie:1987:FCF

[698] A. W. Goudie. Forecasting corporate failure: The use of discriminant analysis within a disaggregated model of the corporate sector. Journal of the Royal Statistical Society. Series A (General), 150(1):69–81, ???? 1987. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981666.

Anonymous:1987:BMa

[699] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 150(1):??, ???? 1987. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981667.

Anonymous:1987:FMb

[700] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 150(2):??, ???? 1987. CODEN JSSAEF. ISSN 0035-9238.

                                        URL http://www.jstor.org/stable/
                                        2981627.

Muirhead:1987:MRA

[701] C. R. Muirhead and S. C. Darby. Modelling the relative and absolute risks of radiation-induced cancers. Journal of the Royal Statistical Society. Series A (General), 150(2):83–118, ???? 1987. CODEN JSSAEF. ISSN 0035-9238.

                                        URL http://www.jstor.org/stable/
                                        2981628.

Gordon:1987:RHC

[702] A. D. Gordon. A review of hierarchical classification. Journal of the Royal Statistical Society. Series A (General), 150 (2):119–137, ???? 1987. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2981629.

Maunder:1987:RUK

[703] W. F. Maunder and M. C. Fleming. Reviews of United Kingdom statistical sources: A progress report. Journal of the Royal Statistical Society. Series A (General), 150(2):138–142, ???? 1987. CODEN JSSAEF. ISSN 0035-9238.

                                        URL http://www.jstor.org/stable/
                                        2981630.

Anandalingam:1987:DPE

[704] G. Anandalingam and Nalin Kulatilaka. Decomposing production efficiency into technical, allocative and structural components. Journal of the Royal Statistical Society. Series A (General), 150 (2):143–151, ???? 1987. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981631.

Elis-Williams:1987:ESP

[705] D. G. Elis-Williams. The effect of spatial population distribution on the cost of delivering local services. Journal of the Royal Statistical Society. Series A (General), 150(2):152–166, ???? 1987. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981632.

Maher:1987:BRB

[706] M. J. Maher. Book review: Plots, Transformations and Regression: An Introduction to Graphical Methods of Diagnostic Regression Analysis, by A. C. Atkinson. Journal of the Royal Statistical Society. Series A (General), 150(2): 167, ???? 1987. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981633.

Chatfield:1987:BRBa

[707] Chris Chatfield. Book review: Random Data: Analysis and Measurement Procedures, by J. S. Bendat; A. G. Piersol. Journal of the Royal Statistical Society. Series A (General), 150(2):167–168, ???? 1987. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981634.

Edwards:1987:BRB

[708] A. W. F. Edwards. Book review: Natural Selection, Heredity, and Eugenics, by R. A. Fisher; Leonard Darwin; J. H. Bennett. Journal of the Royal Statistical Society. Series A (General), 150 (2):168–169, ???? 1987. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981635.

Stone:1987:BRB

[709] M. Stone. Book review: Statistical Inference in Linear Models. Statistical Methods of Model Building, Vol 1, by Helga Bunke; Olaf Bunke. Journal of the Royal Statistical Society. Series A (General), 150(2):169, ???? 1987. CODEN JS-SAEF. ISSN 0035-9238. URL http:/

                                        /www.jstor.org/stable/2981636.

Moore:1987:BRB

[710] B. A. Moore. Book review: Multivariate Statistical Modelling, by Ronald Christensen. Journal of the Royal Statistical Society. Series A (General), 150(2):169– 170, ???? 1987. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2981637.

Connolly:1987:BRB

[711] B. W. Connolly. Book review: Fundamentals of Queueing Theory, by Donald Gross; Carl M. Harris. Journal of the Royal Statistical Society. Series A (General), 150(2):170, ???? 1987. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2981638.

Plewis:1987:BRB

[712] Ian Plewis. Book review: Longitudinal Analysis of Labor Market Data, by James Heckman; Burton Singer. Journal of the Royal Statistical Society. Series A (General), 150(2):170–171, ???? 1987. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2981639.

Bailey:1987:BRB

[713] B. J. R. Bailey and S. M. Lewis. Book review: Selected Tables in Mathematical Statistics, by W. J. Kennedy; R. E.

Odeh. Journal of the Royal Statistical Society. Series A (General), 150(2):171– 172, ???? 1987. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981640.

Chatfield:1987:BRBb

[714] Chris Chatfield. Book review: Multivariate Analysis VI, by P. R. Krishnaiah. Journal of the Royal Statistical Society. Series A (General), 150(2):172, ???? 1987. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981641.

Pettitt:1987:BRB

[715] A. N. Pettitt. Book review: Nonparametric Methods, by P. R. Krishnaiah; P. K. Sen. Journal of the Royal Statistical Society. Series A (General), 150 (2):172–173, ???? 1987. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981642.

Powell:1987:BRB

[716] S. Powell. Book review: Advanced Mathematics for Economists: Static and Dynamic Optimization, by P. J. Lambert. Journal of the Royal Statistical Society. Series A (General), 150(2):173, ???? 1987. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981643.

Copas:1987:BRB

[717] J. B. Copas. Book review: Proceedings of the First International Tampere Seminar on Linear Statistical Models and Their Applications, by T. Pukkila; S. Puntanen. Journal of the Royal Statistical Society. Series A (General), 150 (2):173–174, ???? 1987. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2981644.

Janacek:1987:BRB

[718] G. Janacek. Book review: An Introduction to Bispectral Analysis and Bilinear Time Series Models, by T. Subba Rao; M. M. Gabr. Journal of the Royal Statistical Society. Series A (General), 150(2):174, ???? 1987. CODEN JS-SAEF. ISSN 0035-9238. URL http:/

                                        /www.jstor.org/stable/2981645.

Cox:1987:BRB

[719] M. A. A. Cox. Book review: Order Dependence, by B. F. Schriever. Journal of the Royal Statistical Society. Series A (General), 150(2):175, ???? 1987. CODEN JSSAEF. ISSN 0035-9238.

                                        URL http://www.jstor.org/stable/
                                        2981646.

Carroll:1987:BRB

[720] Patrick Carroll. Book review: Claims Reserving in Non-Life Insurance, by G. C. Taylor. Journal of the Royal Statistical Society. Series A (General), 150(2): 175, ???? 1987. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2981647.

Whitehead:1987:BRB

[721] John Whitehead. Book review: Sequential Methods in Statistics, by G. Barrie Wetherill; Kevin D. Glazebrook. Journal of the Royal Statistical Society. Series A (General), 150(2):176, ???? 1987. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2981648.

Anonymous:1987:BMb

[722] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 150(2):??, ???? 1987. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981649.

Anonymous:1987:FMc

[723] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 150(3):??, ???? 1987. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2981471.

Durbin:1987:SSS

[724] J. Durbin. Statistics and statistical science. Journal of the Royal Statistical Society. Series A (General), 150(3):177– 191, ???? 1987. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981472.

Cleveland:1987:GPV

[725] William S. Cleveland and Robert McGill. Graphical perception: The visual decoding of quantitative information on graphical displays of data. Journal of the Royal Statistical Society. Series A (General), 150(3):192–229, ???? 1987. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981473.

Wilson:1987:EPA

[726] P. R. Wilson and D. J. Elliot. An evaluation of the postcode address file as a sampling frame and its use within OPCS. Journal of the Royal Statistical Society. Series A (General), 150(3):230– 240, ???? 1987. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981474.

Collins:1987:PNC

[727] Martin Collins and Wendy Sykes. The problems of non-coverage and unlisted numbers in telephone surveys in Britain. Journal of the Royal Statistical Society. Series A (General), 150(3):241–253, ???? 1987. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2981475.

Anonymous:1987:PR

[728] Anonymous. President’s report. Journal of the Royal Statistical Society. Series A (General), 150(3):254–255, ???? 1987. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2981476.

Anonymous:1987:HTR

[729] Anonymous. Honorary Treasurer’s report. Journal of the Royal Statistical Society. Series A (General), 150(3):256, ???? 1987. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2981477.

Anonymous:1987:ARC

[730] Anonymous. Annual report of the Council. Journal of the Royal Statistical Society. Series A (General), 150(3):256–280, ???? 1987. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2981478.

Glendinning:1987:BRB

[731] Richard H. Glendinning. Book review: Random Polynomials, by A. T. Bharucha-Reid; M. Sambandham. Journal of the Royal Statistical Society. Series A (General), 150(3):281, ???? 1987. CODEN JSSAEF. ISSN 0035-9238.

                                        URL http://www.jstor.org/stable/
                                        2981479.

Jennison:1987:BRB

[732] C. Jennison. Book review: Robust Statistics: The Approach Based on Influence Functions, by F. R. Hampel; E. M. Ronchetti; P. J. Rousseeuw; W. A. Stahel. Journal of the Royal Statistical Society. Series A (General), 150(3):281– 282, ???? 1987. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981480.

Kimber:1987:BRB

[733] Alan Kimber. Book review: Robustness of Statistical Methods and Nonparametric Statistics, by D. Rasch; M. L. Tiku. Journal of the Royal Statistical Society. Series A (General), 150(3):282– 283, ???? 1987. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981481.

Renshaw:1987:BRB

[734] Arthur E. Renshaw. Book review: Statistical Analysis of Finite Mixture Distributions, by D. M. Titterington; A. F. M. Smith; H. E. Makov. Journal of the Royal Statistical Society. Series A (General), 150(3):283, ???? 1987. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2981482.

Joshi:1987:BRB

[735] Heather Joshi. Book review: Women and Equal Pay: The Effects of Legislation on Female Employment and Wages in Britain, by A. Zabalza; Z. Tzannatos. Journal of the Royal Statistical Society. Series A (General), 150(3):283– 285, ???? 1987. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2981483.

Anonymous:1987:BMc

[736] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 150(3):??, ???? 1987. CODEN JSSAEF. ISSN 0035-9238.

                                        URL http://www.jstor.org/stable/
                                        2981484.

Anonymous:1987:VI

[737] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (General), 150(4):407, ???? 1987. CODEN JSSAEF. ISSN 0035-9238.

                                        URL http://www.jstor.org/stable/
                                        2982038.

Anonymous:1987:FMd

[738] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 150(4):??, ???? 1987. CODEN JSSAEF. ISSN 0035-9238.

                                        URL http://www.jstor.org/stable/
                                        2982039.

Weitzman:1987:EMT

[739] Michael P. Weitzman. The evolution of manuscript traditions. Journal of the Royal Statistical Society. Series A (General), 150(4):287–308, ???? 1987. CODEN JSSAEF. ISSN 0035-9238.

                                        URL http://www.jstor.org/stable/
                                        2982040.

Derbyshire:1987:SRG

[740] M. E. Derbyshire. Statistical rationale for Grant-related expenditure assessment (GREA) concerning personal social services. Journal of the Royal Statistical Society. Series A (General), 150 (4):309–333, ???? 1987. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2982041.

Hand:1987:SKE

[741] D. J. Hand. A statistical knowledge enhancement system. Journal of the Royal Statistical Society. Series A (General), 150(4):334–345, ???? 1987. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982042.

Rudd:1987:EQS

[742] Ernest Rudd. The educational qualifications and social class of the parents of undergraduates entering British universities in 1984. Journal of the Royal Statistical Society. Series A (General), 150(4):346–372, ???? 1987. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982043.

Doll:1987:MEC

[743] Richard Doll. Major epidemics of the 20th Century: From coronary thrombosis to AIDS. Journal of the Royal Statistical Society. Series A (General), 150 (4):373–395, ???? 1987. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982044.

Jones:1987:BRB

[744] M. C. Jones. Book review: The Statistical Analysis of Compositional Data, by J. Aitchison. Journal of the Royal Statistical Society. Series A (General), 150(4):396, ???? 1987. CODEN JS-SAEF. ISSN 0035-9238. URL http:/

/www.jstor.org/stable/2982045.

Pemberton:1987:BRB

[745] J. Pemberton. Book review: Computer Science and Statistics: Proceedings of the Seventeenth Symposium on the Interface, by D. M. Allen. Journal of the Royal Statistical Society. Series A (General), 150(4):396–398, ????

1987. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2982046.

Thomasson:1987:BRB

[746] Mary Thomasson. Book review: MINITAB Guide to the Statistical Analysis of Data, by T. W. Anderson; B. P. Eynon. Journal of the Royal Statistical Society. Series A (General), 150(4):398, ???? 1987. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2982047.

Hawkes:1987:BRB

[747] John Hawkes. Book review: Probability and Measure, by P. Billingsley. Journal of the Royal Statistical Society. Series A (General), 150(4):398–399, ???? 1987. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2982048.

Veevers:1987:BRBa

[748] A. Veevers. Book review: Analysis of Experiments with Missing Data, by Yadolah Dodge. Journal of the Royal Statistical Society. Series A (General), 150(4):399–400, ???? 1987. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2982049.

Hills:1987:BRB

[749] Michael Hills. Book review: The Design and Analysis of Clinical Experiments, by Joseph L. Fleiss. Journal of the Royal Statistical Society. Series A (General), 150(4):400, ???? 1987. CODEN JS-SAEF. ISSN 0035-9238. URL http:/

                                        /www.jstor.org/stable/2982050.

Chapman:1987:BRB

[750] Paul G. Chapman. Book review: Handbook of Econometrics, by Zvi Griliches; Michael D. Intrilligator; Kenneth J. Arrow; Michael D. Intrilligator. Journal of the Royal Statistical Society. Series A (General), 150(4):400–401, ???? 1987. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2982051.

Ray:1987:BRB

[751] W. D. Ray. Book review: Time Series and Econometric Modelling, by Ian B. MacNeil; Gary J. Umphrey. Journal of the Royal Statistical Society. Series A (General), 150(4):401, ???? 1987. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2982052.

Cooper:1987:BRB

[752] R. I. Cooper. Book review: Multivariate Statistical Methods: A Primer, by B. F. J. Manly. Journal of the Royal Statistical Society. Series A (General), 150 (4):401–402, ???? 1987. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2982053.

Veevers:1987:BRBb

[753] A. Veevers. Book review: Beyond ANOVA, Basics of Applied Statistics, by R. G. Miller. Journal of the Royal Statistical Society. Series A (General), 150(4): 402, ???? 1987. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2982054.

Prescott:1987:BRB

[754] P. Prescott. Book review: Truncated and Censored Samples from Normal Populations, by Helmet Schneider. Journal of the Royal Statistical Society. Series A (General), 150(4):403, ???? 1987. CODEN JSSAEF. ISSN 0035-9238.

                                        URL http://www.jstor.org/stable/
                                        2982055.

Warren:1987:BRB

[755] David Warren. Book review: Density Estimation for Statistics and Data Analysis, by B. W. Silverman. Journal of the Royal Statistical Society. Series A (General), 150(4):403–404, ???? 1987. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2982056.

Bartlett:1987:BRB

[756] M. S. Bartlett. Book review: The History of Statistics. The Measurement of Uncertainty before 1900, by S. M. Stigler. Journal of the Royal Statistical Society. Series A (General), 150(4):404– 405, ???? 1987. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2982057.

Gower:1987:BRB

[757] J. C. Gower. Book review: Statistical Image Processing and Graphics, by Edward J. Wegman; Douglas J. De-Priest. Journal of the Royal Statistical Society. Series A (General), 150(4):405, ???? 1987. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2982058.

Anonymous:1987:PAG

[758] Anonymous. Proceedings of the 153rd Annual General Meeting of the Royal Statistical Society. Journal of the Royal Statistical Society. Series A (General), 150(4):406, ???? 1987. CODEN JS-SAEF. ISSN 0035-9238. URL http:/

                                        /www.jstor.org/stable/2982059.

Anonymous:1987:BMd

[759] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 150(4):??, ???? 1987. CODEN JSSAEF. ISSN 0035-9238.

URL http://www.jstor.org/stable/
2982060.

Anonymous:1988:FMa

[760] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(1):??, ???? 1988. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2982176.

Anonymous:1988:JRS

[761] Anonymous. Journal of the Royal Statistical Society. Series A (Statistics in Society). Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(1):1, ???? 1988. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982177.

Anonymous:1988:I

[762] Anonymous. [introduction]. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(1):3–4, ???? 1988. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2982178.

Isham:1988:MMT

[763] Valerie Isham. Mathematical modelling of the transmission dynamics of HIV infection and AIDS: A review. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(1):5– 30 + 44–49, ???? 1988. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982179.

Bailey:1988:SMP

[764] Norman T. J. Bailey. Simplified modelling of the population dynamics of HIV/AIDS. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(1):31–34, ???? 1988. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2982180.

Wilkie:1988:AMA

[765] A. D. Wilkie. An actual model for AIDS. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(1):35–39, ???? 1988. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982181.

Jones:1988:NCI

[766] M. E. Jones and P. J. Solomon. A note on the concept of the incubation period in HIV infection. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(1): 40–42, ???? 1988. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982182.

Boldsen:1988:ITD

[767] Jesper Lier Boldsen, Jens Ledet Jensen, Jes Sogaard, and Michael Sorensen. On the incubation time distribution and the Danish AIDS data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(1): 42–43, ???? 1988. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982183.

Healy:1988:STE

[768] M. J. R. Healy and H. E. Tillett. Shortterm extrapolation of the AIDS epidemic. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(1):50–65, ???? 1988. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982184.

Anderson:1988:EHI

[769] Roy M. Anderson. The epidemiology of HIV infection: Variable incubation plus infectious periods and heterogeneity in sexual activity. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(1):66–98, ???? 1988. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2982185.

Johnson:1988:SBA

[770] Anne M. Johnson. Social and behavioural aspects of the HIV epidemic — a review. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(1):99–119, ???? 1988. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2982186.

Isham:1988:ARD

[771] V. Isham, M. J. R. Healy, H. E. Tillett, R. M. Anderson, and A. M. Johnson. Authors’ reply to the discussion. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(1): 120–126, ???? 1988. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982187.

Kingman:1988:SRA

[772] John Kingman, J. Durbin, David Cox, and M. J. R. Healy. Statistical requirements of the AIDS epidemic. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(1): 127–130, ???? 1988. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982188.

Mann:1988:SIB

[773] Jonathan M. Mann. Statement at an informal briefing on AIDS to the 42nd session of the United Nations General Assembly. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(1):131–136, ???? 1988. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2982189.

Bartholomew:1988:PST

[774] D. J. Bartholomew. Probability, statistics and theology. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(1):137–178, ???? 1988. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2982190.

Gower:1988:SA

[775] J. C. Gower. Statistics and agriculture. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(1):179–200, ???? 1988. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982191.

Cormack:1988:SCE

[776] R. M. Cormack. Statistical challenges in the environmental sciences: A personal view. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(1):201–210, ???? 1988. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2982192.

Weale:1988:RVV

[777] Martin Weale. The reconciliation of values, volumes and prices in the national accounts. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(1):211–221, ???? 1988. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2982193.

Salt:1988:BRB

[778] D. W. Salt. Book review: Probability and Random Variables, by G. P. Beaumont. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(1):222–223, ???? 1988. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982194.

Goldstein:1988:BRB

[779] Michael Goldstein. Book review:

Bayesian Statistics 2: Proceedings on the Second Valencia International Meeting, September 6–10th 1983, by J. M. Bernardo; M. H. De Groot; D. V. Lindley; A. F. M. Smith. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(1):223, ???? 1988. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2982195.

Pike:1988:BRB

[780] Derek J. Pike. Book review: Empirical Model-Building and Response Surfaces, by George E. P. Box; Norman R. Draper. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(1):223–224, ???? 1988. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/2982196.

Haigh:1988:BRB

[781] John Haigh. Book review: Elementary Decision Theory, by H. Chernoff; L. E. Moses. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(1):224–225, ???? 1988. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/2982197.

Bowman:1988:BRB

[782] A. W. Bowman. Book review: Goodness-of-Fit Techniques, by Ralph B. D’Agostino; Michael A. Stephens. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151 (1):225–226, ???? 1988. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2982198.

McHale:1988:BRBa

[783] D. McHale. Book review: Methodes Pratiques de la Statistique, by T. Foucart; A. Bensaber; R. Garneri. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(1):226, ???? 1988. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        2982199.

Owen:1988:BRB

[784] Roger J. Owen. Book review: Decision Theory: An Introduction to the Mathematics of Rationality, by Simon French. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(1):227, ???? 1988. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982200.

Wharton:1988:BRB

[785] F. Wharton. Book review: Operational Research Techniques, by S. French; R. Hartley; L. C. Thomas; D. J. White. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151 (1):227–228, ???? 1988. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982201.

Davison:1988:BRB

[786] A. C. Davison. Book review: Exploring Data Tables, Trends and Shapes, by D. C. Hoaglin; F. Mosteller; J. W. Tukey. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151 (1):228–229, ???? 1988. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982202.

Young:1988:BRB

[787] G. A. Young. Book review: Multivariate Statistical Simulation, by Mark E. Johnson. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(1):229–230, ???? 1988. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2982203.

Gilmour:1988:BRB

[788] Harper Gilmour. Book review: Epidemiology of Occupational Health, by M. Karvonen; M. I. Mikheev. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(1):230, ???? 1988. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        2982204.

Clayton:1988:BRB

[789] David Clayton. Book review: Biostatistical Microcomputing in Pascal, by J. H. Klotz; R. D. Meyer. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(1): 230–231, ???? 1988. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2982205.

Biggins:1988:BRB

[790] J. D. Biggins. Book review: Testing Statistical Hypotheses, by E. L. Lehmann. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151 (1):231–232, ???? 1988. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2982206.

Dewey:1988:BRB

[791] Michael Dewey. Book review: Teaching Health Statistics: Twenty Lesson and Seminar Outlines, by S. K. Lwanga; Cho-Yook Tye. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(1):232, ???? 1988. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/2982207.

Dickson:1988:BRB

[792] D. C. M. Dickson. Book review: Advances in Statistical Sciences, Volume 6 — Actuarial Science, by I. B. MacNeill; G. J. Umphrey. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(1):232–233, ???? 1988. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2982208.

Allin:1988:BRB

[793] Paul Allin. Book review: Clinical Trials:

Design, Conduct, and Analysis, by C. L. Meinhert. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(1):233–234, ???? 1988. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2982209.

Osmond:1988:BRB

[794] Clive Osmond. Book review: Modern Statistical Methods in Chronic Disease Epidemiology, by S. H. Moolgavkar; R. L. Prentice. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(1):234–235, ???? 1988. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2982210.

Dixon:1988:BRB

[795] L. Dixon. Book review: Finite Algorithms in Optimization and Data Analysis, by M. R. Osbourne. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(1): 235–236, ???? 1988. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982211.

Skinner:1988:BRB

[796] C. J. Skinner. Book review: Small Area Statistics: An International Symposium, by R. Platek; J. N. K. Rao; C. E. Sarndal; M. P. Singh. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(1):236, ???? 1988. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/2982212.

Bailey:1988:BRBa

[797] R. A. Bailey. Book review: Experimental Design: Sample Size Determination and Block Designs, by D. Rasch; G. Herrendorfer. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(1):237–238, ???? 1988. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/2982213.

Smeeton:1988:BRB

[798] Nigel Smeeton. Book review: Statistics for Toxicologists, by D. S. Salsburg. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(1):238, ???? 1988. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2982214.

Kennedy:1988:BRB

[799] D. P. Kennedy. Book review: Empirical Processes with Applications to Statistics, by G. R. Shorack; J. A. Wellner. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(1):239, ???? 1988. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2982215.

Ripley:1988:BRB

[800] B. D. Ripley. Book review: Stochastic Geometry and its Applications, by D. Stoyan; W. S. Kendall; J. Mecke. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151 (1):239–240, ???? 1988. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982216.

Davies:1988:BRBa

[801] P. Davies. Book review: Kendall’s Advanced Theory of Statistics. Volume 1. Distribution Theory, by Alan Stuart; J. Keith Ord. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(1):240, ???? 1988. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2982217.

Grimmett:1988:BRB

[802] Geoffrey Grimmett. Book review: Paradoxes in Probability Theory and in Mathematical Statistics, by Gabor J. Szekely. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(1):240–241, ???? 1988. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2982218.

Plewis:1988:BRB

[803] Ian Plewis. Book review: Surveying Subjective Phenomena, by Charles F. Turner; Elizabeth Martin. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(1): 241–242, ???? 1988. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982219.

Anonymous:1988:BMa

[804] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(1):??, ???? 1988. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        2982220.

Anonymous:1988:FMb

[805] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(2):??, ???? 1988. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        2982756.

Hill:1988:SAE

[806] I. D. Hill. Some aspects of elections– to fill one seat or many. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(2): 243–275, ???? 1988. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2982757.

Rosenbaum:1988:YHW

[807] S. Rosenbaum. 100 years of heights and weights. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(2):276–309, ???? 1988. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/2982758.

Cox:1988:NSI

[808] D. R. Cox and P. J. Solomon. A note on selective intervention. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(2):

310–315, ???? 1988. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982759.

Plewis:1988:AUE

[809] Ian Plewis. Assessing and understanding the educational progress of children from different ethnic groups. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(2): 316–326, ???? 1988. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982760.

Kempton:1988:DNC

[810] R. A. Kempton and M. Talbot. The development of new crop varieties. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(2): 327–341, ???? 1988. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982761.

MacFarlane:1988:QOH

[811] Alison MacFarlane and C. K. McPherson. The quality of official health statistics. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(2):342–354, ???? 1988. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982762.

Boreham:1988:SS

[812] John Boreham. Statistics in society, 1838, 1888, 1938. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(2):355–358, ???? 1988. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2982763.

Hill:1988:ODJ

[813] Austin Bradford Hill. Obituary: Dr J. A. Fraser Roberts, CBE, MA, DSc, FRCP, FRC Psych, FRS. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(2): 359–360, ???? 1988. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982764.

Wickham:1988:BRBa

[814] Carol A. C. Wickham. Book review: Statistical Methods in Medical Research, by P. Armitage; G. Berry. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(2): 361–362, ???? 1988. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982765.

Stirzaker:1988:BRB

[815] David Stirzaker. Book review: Applied Probability and Queues, by S. Asmussen. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(2):362, ???? 1988. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982766.

Ansell:1988:BRB

[816] M. J. G. Ansell. Book review: Statistics– Problems and Solutions, by E. E. Basset. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(2):363, ???? 1988. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982767.

Campbell:1988:BRB

[817] M. J. Campbell. Book review: An Introduction to Medical Statistics, by M. Bland. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(2):363–364, ???? 1988. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2982768.

Grubel:1988:BRB

[818] Rudolf Grübel. Book review: Über Statistische Methoden bei Punktprozessen, by F. Böker. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(2):364–365, ???? 1988. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2982769.

Mackenzie:1988:BRB

[819] I. G. Mackenzie. Book review: A Handbook of Introductory Statistical Methods, by C. P. Cox. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(2):365, ???? 1988. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2982770.

Jones:1988:BRBa

[820] M. C. Jones. Book review: A Course in Density Estimation, by L. Devroye. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151 (2):365–366, ???? 1988. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982771.

Hand:1988:BRB

[821] D. J. Hand. Book review: Multivariate Analysis of Ecological Communities, by P. G. N. Digby; R. A. Kempton. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151 (2):366–367, ???? 1988. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982772.

Perkins:1988:BRB

[822] Edwin Perkins. Book review: Markov Processes, Characterization and Convergence, by S. N. Ethier; T. G. Kurtz. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151 (2):367–368, ???? 1988. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982773.

OBrien:1988:BRB

[823] Carl M. O’Brien. Book review: A Folio of Distributions: A Collection of Theoretical Quantile-Quantile Plots, by E. B. Fowlkes. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(2):368–369, ???? 1988. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2982774.

Cook:1988:BRBa

[824] Gordon Cook. Book review: Applied Categorical Data Analysis, by D. H. Freeman. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(2):369–370, ???? 1988. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2982775.

Cook:1988:BRBb

[825] Gordon Cook. Book review: The Statistical Consultant in Action, by D. J. Hand; B. S. Everitt. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(2):370–371, ???? 1988. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2982776.

Hansford-Miller:1988:BRB

[826] Frank Hansford-Miller. Book review:

Judgement and Choice: The Psychology of Decision, by R. M. Hogarth. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(2):371, ???? 1988. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982777.

Patterson:1988:BRB

[827] H. D. Patterson. Book review: Cyclic Designs, by J. A. John. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(2):372, ???? 1988. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2982778.

Ball:1988:BRB

[828] Frank Ball. Book review: Point Processes and Their Statistical Inference, by A. F. Karr. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(2):372–373, ???? 1988. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2982779.

Cooper:1988:BRB

[829] R. I. Cooper. Book review: Multivariate Analysis, by Maurice Kendall. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151 (2):373–374, ???? 1988. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2982780.

Tizzard:1988:BRB

[830] Keith Tizzard. Book review: Statistical Tools for Simulation Practitioners, by J. P. C. Kleijnen. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(2):374, ???? 1988. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/2982781.

Scott:1988:BRBa

[831] Marian Scott. Book review: Model Selection, by H. Linhart; W. Zucchini. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(2):375, ???? 1988. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2982782.

Chatfield:1988:BRB

[832] Chris Chatfield. Book review: Statistical Analysis with Missing Data, by R. J. A. Little; D. B. Rubin. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(2): 375–376, ???? 1988. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2982783.

Wickham:1988:BRBb

[833] Carol A. C. Wickham. Book review:

Statistical Tables for the Design of Clinical Trials, by D. Machin; M. J. Campbell. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(2):376, ???? 1988. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982784.

Barnard:1988:BRBa

[834] George A. Barnard. Book review: [untitled]. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(2):377, ???? 1988. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982785.

Bithell:1988:BRB

[835] John F. Bithell. Book review: [untitled]. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(2):377–378, ???? 1988. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982786.

Barnard:1988:BRBb

[836] George A. Barnard. Book review:

Reviews of United Kingdom Statistical Sources: Volume XIX, Intellectual Property Rights, by W. F. Maunder. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(2):378, ???? 1988. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2982787.

McHale:1988:BRBb

[837] David McHale. Book review: Tensor Methods in Statistics, by P. McCullagh. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151 (2):378–379, ???? 1988. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982788.

Cheek:1988:BRB

[838] P. J. Cheek. Book review: Italian Contributions to the Methodology of Statistics, by A. Naddeo. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(2): 379–380, ???? 1988. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982789.

Baston:1988:BRB

[839] V. J. Baston. Book review: Surveys in Game Theory and Related Topics, by H. J. M. Peters; O. J. Vrieze. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(2):380, ???? 1988. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982790.

Moore:1988:BRB

[840] B. A. Moore. Book review: Population Estimation Methods, Canada, by A. Romaniuc. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(2):380–381, ???? 1988. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2982791.

Stuart:1988:BRB

[841] Michael Stuart. Book review: Introduction to Probability and Statistics for Engineers and Scientists, by S. M. Ross. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151 (2):381–382, ???? 1988. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982792.

Weale:1988:BRB

[842] Martin Weale. Book review: National Accounts Analysis, by G. Stuvel. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(2):382, ???? 1988. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982793.

Copas:1988:BRB

[843] J. B. Copas. Book review: Robust Inference, by M. L. Tiku; W. Y. Tan; N. Balakrishnan. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(2):383, ???? 1988. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2982794.

Jones:1988:BRBb

[844] P. W. Jones. Book review: Stochastic Games with Finite State and Action Spaces, by O. J. Vrieze. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(2):384, ???? 1988. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2982795.

Anonymous:1988:BMb

[845] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(2):??, ???? 1988. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        2982796.

Anonymous:1988:VI

[846] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(3): 575–576, ???? 1988. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2982990.

Anonymous:1988:FMc

[847] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(3):??, ???? 1988. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        2982991.

Cole:1988:FSC

[848] T. J. Cole. Fitting smoothed centile curves to reference data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(3): 385–418, ???? 1988. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2982992.

Begg:1988:PBP

[849] Colin B. Begg and Jesse A. Berlin. Publication bias: A problem in interpreting medical data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(3):419–463, ????

1988. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2982993.

Seaman:1988:SDR

[850] John Seaman and Johnn Rivers. Strategies for the distribution of relief food. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151 (3):464–472, ???? 1988. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982994.

Gunluk-Senesen:1988:SEM

[851] G. Günlük-Şenesen and J. M. Bates.

Some experiments with methods of adjusting unbalanced data matrices. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(3): 473–490, ???? 1988. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982995.

Meade:1988:MLM

[852] Nigel Meade. A modified logistic model applied to human populations. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(3): 491–498, ???? 1988. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982996.

Raeside:1988:USM

[853] R. Raeside. The use of sigmoids in modelling and forecasting human populations. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(3):499–513, ???? 1988. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/2982997.

Anonymous:1988:PR

[854] Anonymous. President’s report. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(3):514, ???? 1988. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        2982998.

Anonymous:1988:HTR

[855] Anonymous. Honorary Treasurer’s report. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(3):515, ???? 1988. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2982999.

Anonymous:1988:ARC

[856] Anonymous. Annual report of the Council. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151 (3):515–542, ???? 1988. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983000.

Anonymous:1988:SS

[857] Anonymous. Statistics in society, 1838, 1888, 1938. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(3):543–545, ???? 1988. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/2983001.

Newby:1988:OFG

[858] R. F. Newby. Obituary: Frank Garwood, 1911–88. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(3):546, ???? 1988. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983002.

Conolly:1988:BRB

[859] B. W. Conolly. Book review: Stochastic Analysis of Computer Storage, by O. I. Aven; E. G. Coffman,; Y. A. Kogan. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(3):547–548, ???? 1988. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983003.

Boys:1988:BRB

[860] Richard J. Boys. Book review: Multivariate Statistical Modeling and Data Analysis, by H. Bozdogan; A. K. Gupta. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151 (3):548–549, ???? 1988. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983004.

West:1988:BRB

[861] Mike West. Book review: Econometrics and Structural Change, by L. D. Broemeling; H. Tsurumi. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(3): 549–550, ???? 1988. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983005.

Ray:1988:BRBa

[862] W. D. Ray. Book review: Basic Queueing Theory, by B. D. Bunday. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(3):550, ????

1988. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        2983006.

Mollison:1988:BRB

[863] Denis Mollison. Book review: Spatial Aspects of Influenza Epidemics, by A. D. Cliff; P. Haggett; J. K. Ord. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(3): 550–551, ???? 1988. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983007.

Matatko:1988:BRB

[864] John Matatko. Book review: Uncertainty in Risk Assessment, Risk Management and Decision Making, by V. T. Covello; L. B. Lane; A. Moghissi; V. R. R. Uppuluri. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(3):551, ???? 1988. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/2983008.

Scott:1988:BRBb

[865] R. G. Scott. Book review: Design and Analysis of Experiments, by M. N. Das; N. C. Giri. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(3):552, ???? 1988. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/2983009.

Bailey:1988:BRBb

[866] B. J. R. Bailey. Book review: Probability and Statistics, by M. H. DeGroot. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(3):552–553, ???? 1988. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983010.

Martin:1988:BRB

[867] R. J. Martin. Book review: Spatial Processes and Spatial Time Series Analysis, by F. Droesbeke. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(3): 553–554, ???? 1988. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983011.

Cockcroft:1988:BRB

[868] N. P. Cockcroft. Book review: Applied Statistics: An Analysis of Variance and Regression, by O. J. Dunn; V. A. Clark. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(3):554–555, ???? 1988. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983012.

Bingham:1988:BRB

[869] M. S. Bingham. Book review: Lectures on Topics in Probability Inequalities, by M. L. Easton. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(3):555, ???? 1988. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983013.

Duffy:1988:BRB

[870] Stephen W. Duffy. Book review: Clinimetrics, by A. R. Feinstein. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(3):

555–556, ???? 1988. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983014.

Page:1988:BRB

[871] E. Page. Book review: Contributions to the Theory and Applications of Statistics, a Volume in Honor of Herbert Solomon, by Herbert Solomon; A. E. Gelfand. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(3):556–557, ???? 1988. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/2983015.

Ray:1988:BRBb

[872] W. D. Ray. Book review: Non-Negative Matrices and Applicable Topics in Linear Algebra, by A. Graham. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(3): 557–558, ???? 1988. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983016.

Jolliffe:1988:BRB

[873] I. T. Jolliffe. Book review: Multivariate Data Analysis with Readings, by J. F. Hair,; R. E. Anderson; R. L. Tatham. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151 (3):558–559, ???? 1988. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983017.

Perry:1988:BRB

[874] Joe N. Perry. Book review: Multiple Comparison Procedures, by Y.

Hochberg; A. C. Tamhane. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(3): 559–560, ???? 1988. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983018.

Metcalfe:1988:BRB

[875] A. V. Metcalfe. Book review: Ecology of Photosynthetic Prokaryotes, by T. Lindholm. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(3):560, ???? 1988. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983019.

Glendinning:1988:BRB

[876] R. H. Glendinning. Book review: Advances in the Statistical Sciences: Volume I, Applied Probability, Stochastic Processes, and Sampling Theory, by I. B. MacNeill; G. J. Umphrey. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(3): 560–561, ???? 1988. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983020.

McKenzie:1988:BRB

[877] Ed. McKenzie. Book review: Advances in the Statistical Sciences: Volume IV, Stochastic Hydrology, by I. B. MacNeill; G. J. Umphrey. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(3):561, ???? 1988. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983021.

Aylen:1988:BRB

[878] Johnathan Aylen. Book review: Reviews of United Kingdom Statistical Sources: Volume XVI, Iron and Steel and Shipbuilding, by W. F. Maunder. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(3): 561–562, ???? 1988. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983022.

Curwen:1988:BRB

[879] P. Curwen. Book review: Reviews of United Kingdom Statistical: Volume XXII, Printing and Publishing, by W. F. Maunder. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(3):562–563, ???? 1988. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/2983023.

Critchley:1988:BRB

[880] Frank Critchley. Book review: Multivariate Data Analysis, by F. Murtagh; A. Heck. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(3):563–564, ???? 1988. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/2983024.

Crowder:1988:BRB

[881] Martin Crowder. Book review: Asymptotic Theory of Statistical Inference, by B. L. S. Prakasa Rao. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(3): 564–565, ???? 1988. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983025.

Kemp:1988:BRB

[882] C. D. Kemp. Book review: Stochastic Simulation, by B. D. Ripley. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(3): 565–567, ???? 1988. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983026.

Sugden:1988:BRB

[883] Roger A. Sugden. Book review: Multiple Imputation for Nonresponse in Surveys, by D. B. Rubin. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(3):567, ???? 1988. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983027.

Bertrand:1988:BRB

[884] Philip V. Bertrand. Book review:

Akaike Information Criterion Statistics, by Y. Sakamoto; M. Ishiguro; G. Kitagawa. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(3):567–568, ???? 1988. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983028.

Janacek:1988:BRB

[885] G. J. Janacek. Book review: Maximum Entropy and Bayesian Spectral Analysis and Estimation Problems, by C. R. Smith; G. Erickson. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(3): 568–569, ???? 1988. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983029.

Abakuks:1988:BRB

[886] Andris Abakuks. Book review: Salt Lab: Microcomputer-Based Practical Classes and Demonstrations for Teaching Basic Statistical Concepts, by W. D. Stirling. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(3):569, ???? 1988. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983030.

Easton:1988:BRB

[887] D. Easton. Book review: Cancer Modeling, by J. R. Thompson; B. W. Brown. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(3):570, ???? 1988. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983031.

Paterson:1988:BRB

[888] Lindsay Paterson. Book review: Optimal Paired Comparison Designs for Factorial Experimets, by E. E. M. Van Berkum. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(3):570–571, ???? 1988. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983032.

Davies:1988:BRBb

[889] P. Davies. Book review: Cluster Analysis in Clinical Chemicstry: A Model, by W. Vogt; D. Nagel; H. Sator. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(3): 571–572, ???? 1988. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983033.

Elton:1988:BRB

[890] R. A. Elton. Book review: Statistical Methods for the Analysis of Biomedical Data, by R. F. Woolson. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(3): 572–573, ???? 1988. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983034.

Priestly:1988:BRB

[891] M. B. Priestly. Book review: The Spectral Analysis of Time Series, by J. G. Zurbenko. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(3):573–574, ???? 1988. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983035.

Anonymous:1988:BMc

[892] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 151(3):??, ???? 1988. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983036.

Anonymous:1989:FMa

[893] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(1):??, ???? 1989. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2982818.

Redfern:1989:PRS

[894] Philip Redfern. Population registers:

Some administrative and statistical pros and cons. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(1):1–41, ???? 1989. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982819.

Crafts:1989:TCB

[895] N. F. R. Crafts, S. J. Leybourne, and T. C. Mills. Trends and cycles in British industrial production, 1700–1913. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(1): 43–60, ???? 1989. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982820.

Wakeford:1989:CLN

[896] Richard Wakeford, Keith Binks, and David Wilkie. Childhood leukaemia and nuclear installations. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(1): 61–86, ???? 1989. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982821.

Silver:1989:MNT

[897] Mick Silver and Khashayar Mahdavy. The measurement of a nation’s terms of trade effect and real national disposable income within a national accounting framework. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(1):87–107, ???? 1989. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2982822.

Anonymous:1989:SSa

[898] Anonymous. Statistics in society, 1839, 1889, 1939. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(1):109–111, ???? 1989. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2982823.

Palmer:1989:BRB

[899] M. K. Palmer. Book review: Multivariate Interpretation of Clinical Laboratory Data, by A. Albert; E. K. Harris. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152 (1):113–114, ???? 1989. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982824.

Lawrance:1989:BRB

[900] A. J. Lawrance. Book review: The Teaching of Practical Statistics, by C. W. Anderson; R. M. Loynes. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(1): 114–115, ???? 1989. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982825.

Farebrother:1989:BRB

[901] R. W. Farebrother. Book review: Plots, Transformations and Regression: An Introduction to Graphical Methods of Diagnostic Regression Analysis, by A. C. Atkinson. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(1):115–116, ???? 1989. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2982826.

Jolliffe:1989:BRBa

[902] F. R. Jolliffe. Book review: Applied Statistics: A First Course, by M. L. Berenson; D. M. Levine; D. Rindskopf. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(1):116, ???? 1989. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2982827.

Baxter:1989:BRB

[903] L. A. Baxter. Book review: Means and their Inequalities, by P. S. Bullen; D. S. Mitrinović; P. M. Vasić. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(1): 116–117, ???? 1989. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2982828.

Diamond:1989:BRB

[904] I. Diamond. Book review: The World Fertility Survey: An Assessment, by J. Cleland; C. Scott. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(1):117–118, ???? 1989. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        2982829.

Schlittgen:1989:BRB

[905] R. Schlittgen. Book review: Randomization Tests, by E. S. Edgington. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(1): 118–119, ???? 1989. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2982830.

Jones:1989:BRBa

[906] M. C. Jones. Book review: Spline Smoothing and Nonparametric Regression, by R. L. Eubank. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(1): 119–120, ???? 1989. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982831.

Grubel:1989:BRBa

[907] Rudolf Grübel. Book review: Regression Analysis and Empirical Processes, by S. A. van de Geer. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(1):120, ???? 1989. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2982832.

Jones:1989:BRBb

[908] P. W. Jones. Book review: Decision Analysis, by G. Gregory. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(1):121, ???? 1989. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2982833.

OBrien:1989:BRB

[909] Carl M. O’Brien. Book review:

Probabilistic Methods in Mathematical Physics, by K. Itô; N. Ikeda. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(1): 121–122, ???? 1989. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982834.

Covey-Crump:1989:BRB

[910] P. A. K. Covey-Crump. Book review: Reason and Being, by B. G. Kuznetsov. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(1):122, ???? 1989. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982835.

Burman:1989:BRB

[911] J. P. Burman. Book review: Don and Mandarin–Memoirs of an Economist, by Donald MacDougall. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(1):123, ???? 1989. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2982836.

Burrows:1989:BRB

[912] B. L. Burrows. Book review: Linear Structures, by J. R. Magnus. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(1): 123–124, ???? 1989. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982837.

Smeeton:1989:BRB

[913] Nigel Smeeton. Book review: Reviews of United Kingdom Statistical Sources: Vol. XX, Religion, by W. F. Maunder; M. C. Fleming. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(1):124–125, ???? 1989. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2982838.

Bull:1989:BRB

[914] P. A. Bull. Book review: Reviews of United Kingdom Statistical Sources: Vol. XXI, Finance, by W. F. Maunder; M. C. Fleming. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(1):125–126, ???? 1989. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2982839.

Geary:1989:BRB

[915] D. N. Geary. Book review: Mixture Models: Inference and Applications to Clustering, by G. J. McLachlan; K. E. Basford. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(1):126–127, ???? 1989. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2982840.

Silver:1989:BRBa

[916] Mick Silver. Book review: The European Economic Recovery: A Need for New Policies?, by F. Modigliani. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(1): 127–128, ???? 1989. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982841.

Lovie:1989:BRB

[917] Pat Lovie. Book review: British Imports of Consumer Goods: A Study of Import Penetration 1974–1985, by A. D. Morgan. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(1):128–129, ???? 1989. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/2982842.

Hansford-Miller:1989:BRB

[918] Frank Hansford-Miller. Book review: The Process of Science: Contemporary Philosophical Approaches to Understanding Scientific Practice, by N. J. Nersessian. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(1):129–130, ???? 1989. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/2982843.

Schimek:1989:BRB

[919] M. G. Schimek. Book review: Automation in Survey Processing, by Netherlands Central Bureau of Statistics. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(1): 130–131, ???? 1989. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2982844.

Dewey:1989:BRB

[920] Michael Dewey. Book review: Lecture Notes on Medical Statistics, by A. Petrie. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(1):131, ???? 1989. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2982845.

Krotki:1989:BRB

[921] Karol J. Krotki. Book review: A Study on the Future of the Census of Population: Alternative Approaches, by P. Redfern. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(1):132–133, ???? 1989. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2982846.

Seheult:1989:BRB

[922] A. H. Seheult and P. J. Green. Book review: Robust Regression and Outlier Detection, by P. J. Rousseeuw; A. M. Leroy. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(1):133–134, ???? 1989. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982847.

Parramore:1989:BRB

[923] K. Parramore. Book review: Linear Models for Unbalanced Data, by S. R. Searle. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(1):134–135, ???? 1989. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2982848.

Jolliffe:1989:BRBb

[924] F. R. Jolliffe. Book review: Understanding Data, by P. Sprent. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(1):135, ???? 1989. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2982849.

Grimmett:1989:BRB

[925] Geoffrey Grimmett. Book review: Counterexamples in Probability, by J. Stoyanov. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(1):135–136, ???? 1989. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982850.

Goldstein:1989:BRBa

[926] Michael Goldstein. Book review: Probability and Bayesian Statistics, by R. Viertl. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(1):136–137, ???? 1989. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2982851.

Gilmour:1989:BRB

[927] Harper Gilmour. Book review: Statistics for Health Management and Research, by M. Woodward; L. M. A. Francis. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(1):137, ???? 1989. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2982852.

Hinde:1989:BRBa

[928] John Hinde. Book review: Directory of Statistical Microcomputer Software, 1988 Edition, by W. A. Woodward; A. C. Elliott; H. L. Gray; D. C. Matlock. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152 (1):137–138, ???? 1989. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2982853.

Anonymous:1989:BMa

[929] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(1):??, ???? 1989. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        2982854.

Anonymous:1989:FMb

[930] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(2):??, ???? 1989. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2982911.

Hadjipavlou:1989:RSB

[931] S. Hadjipavlou and G. Carr-Hill. A revised set of blast casualty rates for civil defence use: An overview. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(2): 139–156, ???? 1989. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982912.

Steadman:1989:CME

[932] J. P. Steadman. Computer models of the effects of nuclear attack on Britain: The Home Office and its critics. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(2): 157–181, ???? 1989. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982913.

Smith:1989:SPA

[933] R. I. Smith, D. Fowler, and J. N. Cape.

The statistics of phytotoxic air pollutants. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(2):183–198, ???? 1989. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982914.

Curnow:1989:SAD

[934] R. N. Curnow and T. B. L. Kirkwood. Statistical analysis of deoxyribonucleic acid sequence data — a review. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(2): 199–220, ???? 1989. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2982915.

McPherson:1989:SVS

[935] Glen McPherson. The scientists’ view of statistics — a neglected area. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(2): 221–240, ???? 1989. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2982916.

Bound:1989:SS

[936] J. A. Bound and A. S. C. Ehrenberg. Significant sameness. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(2): 241–247, ???? 1989. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2982917.

Kemsley:1989:LE

[937] W. F. F. Kemsley and S. Rosenbaum. Letters to the editors. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(2): 249–250, ???? 1989. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2982918.

Anonymous:1989:SSb

[938] Anonymous. Statistics in society, 1839, 1889, 1939. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(2):251–252, ???? 1989. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2982919.

Shanbhag:1989:BRBa

[939] D. N. Shanbhag. Book review: Pareto Distributions, by B. C. Arnold. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(2): 253–254, ???? 1989. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982920.

Morgan:1989:BRB

[940] Byron J. T. Morgan. Book review: Introduction to Optimization, by E. M. L. Beale. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(2):254–255, ???? 1989. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982921.

Shanbhag:1989:BRBb

[941] D. N. Shanbhag. Book review: Data Distributions, by R. Christensen. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152 (2):255, ???? 1989. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982922.

Grubel:1989:BRBb

[942] R. Grübel. Book review: Seminar on Stochastic Processes, 1987, by E.

Çinlar; K. L. Chung; R. K. Getoor. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(2):255, ???? 1989. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982923.

Holgate:1989:BRB

[943] P. Holgate. Book review: Lognormal Distributions: Theory and Applications, by E. L. Crow; K. Shimizu. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(2):256, ???? 1989. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982924.

Goldstein:1989:BRBb

[944] M. Goldstein. Book review: Bayesian Analysis and Uncertainty in Economic Theory, by R. M. Cyert; M. H. DeGroot. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(2):256–257, ???? 1989. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982925.

Jolliffe:1989:BRBc

[945] F. R. Jolliffe. Book review: The Method of Paired Comparisons, by H. A. David. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(2):258, ???? 1989. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982926.

Cressie:1989:BRB

[946] Noel Cressie. Book review: Asymptotics for Generalized Chi-Square Goodnessof-Fit Tests, by F. C. Drost. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(2): 258–259, ???? 1989. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982927.

Bibby:1989:BRBa

[947] John Bibby. Book review: Pascal’s Arithmetical Triangle, by A. W. F. Edwards. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(2):259–260, ???? 1989. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982928.

Watkins:1989:BRB

[948] Alan Watkins. Book review: Maximum-Entropy and Bayesian Methods in Science and Engineering, by G. J. Erickson; C. R. Smith. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(2):260, ???? 1989. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2982929.

Hinde:1989:BRBb

[949] John Hinde. Book review: Introduction to Optimization Methods and their Application in Statistics, by B. S. Everitt. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152 (2):260–261, ???? 1989. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982930.

Bell:1989:BRB

[950] J. F. Bell. Book review: Linear Least Squares Computations, by R. W. Farebrother. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(2):261–262, ???? 1989. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/2982931.

Gilchrist:1989:BRB

[951] Robert Gilchrist. Book review: A Guide to SPSS/PC+, by N. Frude. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(2): 262–263, ???? 1989. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2982932.

Nicholl:1989:BRB

[952] J. P. Nicholl. Book review: Roads Accident Statistics, by T. P. Hutchinson. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(2):263, ???? 1989. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2982933.

Sharples:1989:BRB

[953] L. D. Sharples. Book review: Essentials of Medical Statistics, by B. R. Kirkwood. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(2):263–264, ???? 1989. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2982934.

Everitt:1989:BRB

[954] B. S. Everitt. Book review: Principles of Multivariate Analysis, by W. Krzanowski. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(2):264–265, ???? 1989. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2982935.

Ahmad:1989:BRBa

[955] R. Ahmad. Book review: Information Theory, Statistical Decision Functions, Random Processes, by S. Kubik. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152 (2):265–266, ???? 1989. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982936.

Hinde:1989:BRBc

[956] John Hinde. Book review: Genstat 5:

An Introduction, by P. Lane; N. Galwey; N. Alvey. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(2):266, ???? 1989. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982937.

Darby:1989:BRB

[957] Sarah C. Darby. Book review: Misclassification of Smoking Habits and Passive Smoking: A Review of the Evidence, by P. N. Lee. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(2):266–267, ???? 1989. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2982938.

Quinney:1989:BRB

[958] D. A. Quinney. Book review: Linear Structures, by J. R. Magnus. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(2):267, ???? 1989. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        2982939.

Crowder:1989:BRB

[959] Martin Crowder. Book review: Matrix Differential Calculus with Applications in Statistics and Econometrics, by J. R. Magnus; H. Neudecker. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(2): 268–269, ???? 1989. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2982940.

Baczkowski:1989:BRB

[960] A. J. Baczkowski. Book review: Pattern-Analytic Clustering: Theory, Method, Research and Configural Findings, by L. L. McQuitty. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(2):269, ???? 1989. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/2982941.

Currie:1989:BRB

[961] Iain D. Currie. Book review: Distribution-Free Tests, by H. R. Neave; P. L. Worthington. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(2):269–270, ???? 1989. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        2982942.

Stephens:1989:BRB

[962] H. J. Stephens. Book review: Selected Tables in Mathematical Statistics, by R. E. Odeh; J. M. Davenport. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(2):271, ???? 1989. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982943.

Freeman:1989:BRB

[963] G. H. Freeman. Book review: A Manual of Crop Experimentation, by S. C. Pearce; G. M. Clarke; G. V. Dyke; R. E. Kempson. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(2):271–272, ???? 1989. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2982944.

Magnus:1989:BRB

[964] Jan R. Magnus. Book review: Estimation of Variance Components and Applications, by C. R. Rao; J. Kleffe. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152 (2):272–273, ???? 1989. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982945.

Bibby:1989:BRBb

[965] John Bibby. Book review: Econometrics of Planning and Efficiency, by J. K. Sengupta; G. K. Kadekodi. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(2):273, ???? 1989. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2982946.

Ray:1989:BRB

[966] W. D. Ray. Book review: Elementary Applications of Probability Theory, by H. C. Tuckwell. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(2):273–274, ????

1989. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        2982947.

Ahmad:1989:BRBb

[967] R. Ahmad. Book review: Tauberian Theorems for Generalised Functions, by V. S. Vlidimirov; Yu. N. Drozzinov; B. I. Zavialov. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(2):274–275, ???? 1989. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/2982948.

Silver:1989:BRBb

[968] Mick Silver. Book review: Store Choice, Store Location and Market Analysis, by Neil Wrigley. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(2):275, ???? 1989. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/2982949.

Conway:1989:BRB

[969] D. A. Conway. Book review: Two Models of an Open Economy, by B.-C. Ysander. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(2):276, ???? 1989. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2982950.

Anonymous:1989:BMb

[970] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(2):??, ???? 1989. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        2982951.

Anonymous:1989:VI

[971] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152 (3):428, ???? 1989. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983124.

Anonymous:1989:FMc

[972] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(3):??, ???? 1989. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983125.

Kingman:1989:SR

[973] John Kingman. Statistical responsibility. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152 (3):277–285, ???? 1989. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983126.

Simpson:1989:SRF

[974] Stephen Simpson. School roll forecasts:

Their uses, their accuracy and educational reform. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(3):287–304, ???? 1989. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983127.

Darby:1989:RSS

[975] Sarah Darby. Royal Statistical Society meeting on cancer near nuclear installations. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(3):305, ???? 1989. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983128.

Gardner:1989:RRI

[976] Martin J. Gardner. Review of reported increases of childhood cancer rates in the vicinity of nuclear installations in the UK. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(3):307–325, ???? 1989. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983129.

Wheldon:1989:ARR

[977] T. E. Wheldon. The assessment of risk of radiation-induced childhood leukaemia in the vicinity of nuclear installations. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152 (3):327–339, ???? 1989. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983130.

Doll:1989:ECL

[978] Richard Doll. The epidemiology of childhood leukaemia. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(3):341–351, ???? 1989. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983131.

Hills:1989:SMU

[979] Michael Hills and Freda Alexander. Statistical methods used in assessing the risk of disease near a source of possible environmental pollution: A review. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152 (3):353–384, ???? 1989. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983132.

Anonymous:1989:RCS

[980] Anonymous. Report of the Council for the session 1988–89. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(3): 385–414, ???? 1989. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983133.

Anonymous:1989:SSc

[981] Anonymous. Statistics in society, 1839, 1889, 1939. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(3):415–416, ???? 1989. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983134.

Lindley:1989:OHJ

[982] D. V. Lindley. Obituary: Harold Jeffreys, 1891–1989. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(3):417–418, ???? 1989. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983135.

Smith:1989:OPM

[983] Cedric A. B. Smith. Obituary: Pat Moran, 1917–1988. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(3):419, ???? 1989. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983136.

Silver:1989:BRBc

[984] Mick Silver. Book review: Repeat-Buying: Facts, Theory and Applications, by A. S. C. Ehrenberg. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(3): 420–421, ???? 1989. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983137.

Stern:1989:BRB

[985] George Stern. Book review: British Social Trends since 1900: A Guide to the Changing Social Structure of Britain, by A. H. Halsey. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(3):421, ???? 1989. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983138.

Tong:1989:BRB

[986] H. Tong. Book review: The Statistical Theory of Linear Systems, by E. J. Hannan; M. Deistler. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(3):422, ???? 1989. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983139.

Silver:1989:BRBd

[987] Mick Silver. Book review: National Income and Economic Progress: Essays in Honour of Colin Clark, by D. Ironmonger; J. O. N. Perkins; Tran Van Hoa. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(3):422–423, ???? 1989. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983140.

Veevers:1989:BRB

[988] A. Veevers. Book review: Quality Control and Reliability, by P. R. Krishnaiah; C. R. Rao. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(3):423–424, ???? 1989. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983141.

Bibby:1989:BRBc

[989] John Bibby. Book review: Linear Structures, by J. R. Magnus. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(3): 424–425, ???? 1989. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983142.

Holden:1989:BRB

[990] Kenneth Holden. Book review: Applied Time Series Analysis for Business and Economic Forecasting, by S. M. Nazem. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152 (3):425–426, ???? 1989. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983143.

Gates:1989:BRB

[991] John Gates. Book review: Probability, Method and Measurement, by A. O’Hagan. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(3):426–427, ???? 1989. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983144.

Plewis:1989:BRB

[992] Ian Plewis. Book review: Introduction to Probability, by D. R. Robinson; A. W. Bowman; Introduction to Statistics by A. W. Bowman; D. R. Robinson. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(3):427, ???? 1989. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983145.

Anonymous:1989:BMc

[993] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 152(3):??, ???? 1989. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        2983146.